### v1

# VisionTS: Visual Masked Autoencoders Are Free-Lunch Zero-Shot Time Series Forecasters

Abstract
Foundation models have emerged as a promising approach in time series forecasting (TSF). Existing approaches either fine-tune large language models (LLMs) or build large-scale time-series datasets to develop TSF foundation models. However, these methods face challenges due to the severe cross-domain gap or in-domain heterogeneity. In this paper, we explore a new road to building a TSF foundation model from rich and high-quality natural images, based on the intrinsic similarities between images and time series. To bridge the gap between the two domains, we reformulate the TSF task as an image reconstruction task, which is further processed by a visual masked autoencoder (MAE) self-supervised pre-trained on the ImageNet dataset. Surprisingly, without further adaptation in the time-series domain, the proposed VisionTS could achieve superior zero-shot forecasting performance compared to existing TSF foundation models. With minimal fine-tuning, VisionTS could further improve the forecasting and achieve state-of-the-art performance in most cases. These findings suggest that visual models could be a free lunch for TSF and highlight the potential for future cross-domain research between computer vision and TSF. Our code is publicly available at https://github.com/Keytoyze/VisionTS.
Figure 1: Long-term forecasting performance comparison. Our VisionTS, without any training on time series data, outperforms the largest foundation model MoiraiLarge in the zero-shot setting.
1 Introduction
Foundation models (Bommasani et al., 2021) have revolutionized natural language processing (NLP) and computer vision (CV) in recent years (Brown et al., 2020; He et al., 2022). By pretraining on large-scale data, they have shown remarkable few-shot and even zero-shot performance across various downstream tasks. This has motivated an emergent paradigm shift in time series forecasting (TSF), moving from a traditional one-model-per-dataset framework to universal forecasting with a single pre-trained model (Woo et al., 2024; Goswami et al., 2024). A TSF foundation model can greatly reduce the need for downstream data and demonstrate strong forecasting performance on diverse domains, such as energy consumption planning, weather forecasting, and traffic flow.
We have recently witnessed two roads to building a TSF foundation model. The first tries to fine-tune the weights of large language models (LLMs) which have been pre-trained on text data for TSF tasks (i.e., text-based) (Zhou et al., 2023; Jin et al., 2024), based on the observation that LLMs and TSF models share a similar left-to-right forecasting paradigm. However, due to the significant gap between these two modalities, the effectiveness of such transferability between language and time series has recently been questioned by Tan et al. (2024).
The second road focuses on constructing large-scale time-series datasets collected from diverse domains to train a TSF foundation model from scratch (i.e., TS-based) (Woo et al., 2024; Das et al., 2024). Nevertheless, unlike images or language with unified formats, time series data is highly heterogeneous in length, frequency, dimensionality, domains, and semantics, limiting the transferability between pre-training and downstream domains. Until recently, constructing a high-quality dataset remains challenging and is still in the early exploration stage.
In this paper, we investigate a third road that is less explored yet promising: building TSF foundation models with pre-trained visual models. Our starting point is the intrinsic similarities between natural images and TS: 1 Similar modalities: Unlike discrete texts, both images and TS are continuous. 2 Similar origin: Both TS and images are observations of real-world physical systems, whereas languages are products of human cognitive processes. 3 Similar information density: Languages are human-generated signals with high semantic density, while images and TS data are natural signals with heavy redundancy (He et al., 2022). 4 Similar features: As shown in Fig. 2, images often display many features of real-world time series, which are rarely found in language data. These findings suggest that visual models could be easier to adapt to TSF tasks than language models. Based on these insights, we are motivated to answer the question: “Can a visual model pre-trained on images be a free-lunch foundation model for zero-shot time series forecasting?
Figure 2: An image of the ImageNet dataset (Deng et al., 2009), in which the pixel arrays can display many well-known features of real-world time series, such as trend, seasonality, and stationarity (Qiu et al., 2024). By self-supervised pre-training on ImageNet, it is reasonable that a visual model could understand these features and exhibit a level of time-series forecasting ability.
We focus on visual masked autoencoder (MAE) [^1] We use fonts to distinguish MAE (Masked Autoencoder) and MAE (Mean Absolute Error) in this paper., a popular CV foundation model (He et al., 2022) by self-supervised pre-training on ImageNet (Deng et al., 2009). Inspired by the well-known prompt technique in NLP (Schick & Schütze, 2021), we propose a simple method to reformulate TSF as a patch-level image reconstruction task to bridge the gap between pre-training and downstream tasks. Specifically, we transform 1D TS data into 2D matrices via segmentation. Then, we render the matrices into images and align the forecasting window with masked image patches. This method allows us to make zero-shot forecasting without further adaptation.
We evaluate our proposed VisionTS on multiple TSF benchmarks, including long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021) and Monash (Godahewa et al., 2021). As demonstrated in Fig. 1, without any further adaptation in the time-series domain, a vanilla MAE can surprisingly achieve a comparable performance or even outperform the state-of-the-art (SOTA) zero-shot TSF foundation models, including text-based and TS-based methods. By fine-tuning MAE on each downstream dataset for only one epoch, VisionTS can lead to a SOTA performance on most long-term TSF benchmarks. Our findings suggest that time series and natural images may be two sides of a coin, and visual models can be a free lunch for time series forecasting. We hope our findings can inspire future cross-domain research on CV and TSF.
Our contributions are summarized as follows:
  • We explore a road to building a TSF foundation model from natural images, which is conceptually different from the existing text-based and TS-based methods.
  • We introduce VisionTS, a novel TSF foundation model based on a visual MAE. To bridge the gap between the two modalities, we reformulate the TSF task into an image reconstruction task.
  • Comprehensive evaluations of VisionTS demonstrate its significant zero-shot performance, which even surpasses few-shot text-based TSF foundation models and achieves comparable or superior results to zero-shot TS-based models.
2 Preliminaries
Time Series Forecasting (TSF)
For a multivariate time series with $M$ variables, let ${\bm{x}}_{t}\in\mathbb{R}^{M}$ represent the value at $t$-th time step. Given a historical sequence (i.e., look-back window) ${\bm{X}}_{t-L:t}=[{\bm{x}}_{t-L},\cdots,{\bm{x}}_{t-1}]\in\mathbb{R}^{L\times M}$ with context length $L$, the TSF task is to predict future values (i.e., forecast horizon) with prediction length $H$: $\hat{{\bm{X}}}_{t:t+H}=[{\bm{x}}_{t},\cdots,{\bm{x}}_{t+H-1}]\in\mathbb{R}^{H\times M}$.
Patch-Level Image Reconstruction
To obtain high-quality visual representation for downstream CV tasks, He et al. (2022) proposed masked autoencoder (MAE) to pre-train a Vision Transformer (ViT) (Dosovitskiy et al., 2021) using a patch-level image reconstruction task on ImageNet. Specifically, for an image of size $W\times W$ (where $W$ represents both the width and height, as ImageNet images are square), the image is evenly divided into $N\times N$ patches, each with a width and height of $S=\nicefrac{{W}}{{N}}$. During pre-training, some random patches are masked, while the remaining visible patches are fed into the ViT with their position encodings. MAE are trained to reconstruct the masked pixel values from these visible patches.
3 Methodology
As noted in the Introduction, TS and images share intrinsic similarities, suggesting the transfer potential of visual pre-trained models (particularly MAE in this paper) for TSF tasks. This section explains how to reformulate TSF tasks into MAE’s pre-training task, i.e., patch-level image reconstruction.
Our idea is straightforward: map the look-back/forecasting windows to visible/masked patches, respectively. This idea is supported by the recent success of prompt tuning (Schick & Schütze, 2021) in NLP, where the predictions for [mask] token in pre-trained language models (e.g., BERT (Devlin et al., 2019)) are directly used for downstream tasks. By unifying the forms of the two tasks, we bridge the gap between the two domains, enabling a MAE for zero-shot TSF directly without adapting the pre-trained parameters.
Notably, this idea is limited to univariate forecasting since multivariates are intractable to be encoded in a single image. Fortunately, recent work shows that channel independence — predicting each variable separately for multivariate forecasting — can still be effective (Nie et al., 2022; Han et al., 2024). Therefore, we leave the exploration of multivariate interactions for future work.
However, implementing this idea poses a challenge: the dimension of time-series data (1D) is different from images (2D). Moreover, the size of images in the pre-training dataset is fixed at 224 $\times$ 224, while the lengths of time series data can vary dynamically. In the following, we describe the details of VisionTS to address this challenge. Our architecture is depicted in Fig. 3.
Figure 3: VisionTS architecture. The input is first segmented by period, rendered into a grayscale image, and then aligned with the visible patches on the left through resampling. MAE is used to predict the masked patches on the right, and the reconstructed image is then reversed to forecasting.
Segmentation
[…]
Normalization
MAE standardizes each image based on the mean and standard deviation computed on ImageNet. Therefore, we apply instance normalization to ${\bm{I}}_{\text{raw}}$, which is also a standard practice in current TSF (Kim et al., 2022). Notably, we observed that normalizing ${\bm{I}}_{\text{raw}}$ to a standard deviation of $r$, where $r$ is a hyperparameter less than 1, yields superior performance. One explanation is that the magnitude of inputs/outputs during MAE pretraining is constrained by the limited range of color values. Therefore, reducing the magnitude of ${\bm{I}}_{\text{raw}}$ prevents exceeding these limits. However, an excessively low $r$ can result in values that are difficult to distinguish. We found that a moderate value (0.4) of $r$ performs well across most scenarios (Section B.9). Let ${\bm{I}}_{\text{norm}}$ denote the normalized matrix, which is computed as follows:
${\bm{I}}_{\text{norm}}=r\cdot\frac{{\bm{I}}_{\text{raw}}-\text{Mean}({\bm{I}}_{\text{raw}})}{\text{Standard-Deviation}({\bm{I}}_{\text{raw}})}.$
Rendering
It is well-known that each image has three channels. We simply render ${\bm{I}}_{\text{norm}}$ as a grayscale image ${\bm{I}}_{\text{grey}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor\times 3}$, where all three channels are identical to ${\bm{I}}_{\text{norm}}$. This choice is purely result-driven: In our early experiments, we added a convolutional layer with three output channels to convert the grayscale image into a color image and then fine-tuned it to find the optimal color transformation, which, however, did not significantly influence the performance.
Alignment
Our goal is to predict the columns on the right of ${\bm{I}}_{\text{grey}}$ to forecast the future sequence. A straightforward approach is to treat ${\bm{I}}_{\text{grey}}$ as the visible left portion and the predicted columns as the masked right portion. However, since the image size during pre-training may not match the size of ${\bm{I}}_{\text{grey}}$, we propose to resize ${\bm{I}}_{\text{grey}}$ to align with the pre-training data. Formally, let the total number of 2D patches used in pre-training be $N\times N$ and the size of each patch be $S\times S$. We set the number of visible patches to $N\times n$ and the masked patches to $N\times(N-n)$, where $n=\lfloor N\cdot\nicefrac{{L}}{{(L+H)}}\rfloor$ is determined by the ratio of context length $L$ to prediction length $H$. We resample the image ${\bm{I}}_{\text{grey}}$ to adjust the size from the original dimensions $(P,\lfloor\nicefrac{{L}}{{P}}\rfloor)$ to $(N\cdot S,n\cdot S)$, making it more compatible with MAE. We select bilinear interpolation for the resampling process.
Moreover, we found that reducing the width of the visible portion can further improve performance, likely because the number of input periods ($\lfloor\nicefrac{{L}}{{P}}\rfloor$) is generally smaller than the period length $P$. Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$ to reduce the width, i.e.,
$n=\left\lfloor c\cdot N\cdot\frac{L}{L+H}\right\rfloor.$
Reconstruction and Forecasting
After obtaining the MAE-reconstructed image, we simply reverse the previous steps for forecasting. Specifically, we resize the entire image back to the original time series segmentations through the same bilinear interpolation, and average the three channels to obtain a single-channel image. After de-normalizing and flattening, the forecasting window can be extracted.
4 Experiments
We use MAE (Base) with 112M parameters as our backbone. We conduct experiments on popular TSF benchmarks, including long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021) and Monash benchmark (Godahewa et al., 2021). We select representative baselines for comparison, including TS-based foundation models: Moirai (Woo et al., 2024) and TimesFM (Das et al., 2024); Text-based foundation models: TimeLLM (Jin et al., 2024), GPT4TS (Zhou et al., 2023), and LLMTime (Gruver et al., 2023); and other popular TSF baselines covering both Transformer-based, MLP-based and CNN-based architectures. Details are elaborated in Appendix A.
4.1 Zero-Shot Time Series Forecasting
Setups
We first evaluate VisionTS’s zero-shot TSF performance without any fine-tuning on time-series modalities. To prevent data leakage and assess the out-of-distribution capabilities, we selected six widely-used datasets from the long-term TSF benchmark that are not included in Moirai’s pre-training set for evaluation. Since most baselines cannot perform zero-shot forecasting, we report their few-shot results by fine-tuning on the 10% of the individual target datasets. We also evaluate the Monash benchmark including 29 test datasets, which is more challenging for VisionTS since they were used in Moirai’s pre-training but not for VisionTS. We set the hyperparameters to $r=c=0.4$. Following common practice (Nie et al., 2022; Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section B.1.
Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Section B.2. Bold: the best result. Comparisons with TimesFM and LLMTime are detailed in Section B.3.
[Uncaptioned image] Zero-Shot [Uncaptioned image] Few-Shot (10% In-distribution Downstream Dataset)
Pretrain [Uncaptioned image] Images [Uncaptioned image] Time-series [Uncaptioned image] Text [Uncaptioned image] No Pretrain
Method VisionTS MoiraiS MoiraiB MoiraiL TimeLLM GPT4TS DLinear PatchTST TimesNet Autoformer Informer
ETTh1 MSE 0.390 0.400 0.434 0.510 0.556 0.590 0.691 0.633 0.869 0.702 1.199
MAE 0.414 0.424 0.439 0.469 0.522 0.525 0.600 0.542 0.628 0.596 0.809
ETTh2 MSE 0.333 0.341 0.346 0.354 0.370 0.397 0.605 0.415 0.479 0.488 3.872
MAE 0.375 0.379 0.382 0.377 0.394 0.421 0.538 0.431 0.465 0.499 1.513
ETTm1 MSE 0.374 0.448 0.382 0.390 0.404 0.464 0.411 0.501 0.677 0.802 1.192
MAE 0.372 0.410 0.388 0.389 0.427 0.441 0.429 0.466 0.537 0.628 0.821
ETTm2 MSE 0.282 0.300 0.272 0.276 0.277 0.293 0.316 0.296 0.320 1.342 3.370
MAE 0.321 0.341 0.321 0.320 0.323 0.335 0.368 0.343 0.353 0.930 1.440
Electricity MSE 0.207 0.233 0.188 0.188 0.175 0.176 0.180 0.180 0.323 0.431 1.195
MAE 0.294 0.320 0.274 0.273 0.270 0.269 0.280 0.273 0.392 0.478 0.891
Weather MSE 0.269 0.242 0.238 0.260 0.234 0.238 0.241 0.242 0.279 0.300 0.597
MAE 0.292 0.267 0.261 0.275 0.273 0.275 0.283 0.279 0.301 0.342 0.495
Average MSE 0.309 0.327 0.310 0.329 0.336 0.360 0.407 0.378 0.491 0.678 1.904
MAE 0.345 0.357 0.344 0.350 0.368 0.378 0.416 0.389 0.446 0.579 0.995
$\bf 1^{st}$ count 7 0 3 1 2 1 0 0 0 0 0
Results on Long-Term TSF Benchmark
Table 1 shows that VisionTS surprisingly achieves the best forecasting performance in most cases (7 out of 14). Specifically, VisionTS demonstrates a relative average MSE reduction of approximately 6% compared to MoiraiSmall and MoiraiLarge, and performs comparably to MoiraiBase. When compared to the various few-shot baselines, VisionTS shows a relative average MSE reduction ranging from 8% to 84%. Given that all baselines except for VisionTS are trained on the time-series domain, this result is particularly encouraging. It suggests that the transferability from images to time-series is stronger than from text to time-series, and even comparable to the in-domain transferability between time-series. We also include a comparison with traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section B.4, where VisionTS still outperforms all of these traditional methods.
Results on Monash Benchmark
Fig. 4 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset (e.g., FFNN, WaveNet, and TBATS) and significantly outperforms the other cross-domain baseline (i.e., LLMTime). It achieves second place among all baselines, just behind Moirai that pre-trained on all the training datasets. This highlights VisionTS’s strong zero-shot forecasting ability and the effective cross-modality transferability between image and time series.
Qualitative Analysis: When does VisionTS perform well, and when does it not?
In Appendix D, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails). When the input exhibits strong regularity (Fig. 8), VisionTS effectively forecasts both the periodicity (via segmentation) and trends (via MAE’s capabilities). In contrast, Moirai, akin to seasonal naïve methods, struggles to capture inter-period trends. For less-structured input (Figs. 9, 10 and 11), Moirai adopts a conservative approach with lower volatility to minimize errors, while VisionTS takes a more aggressive stance. This strategy occasionally yields more accurate trend predictions (Figs. 9 and 10) but may also result in greater MAE (Fig. 11).
Figure 4: Aggregated results on the Monash TSF Benchmark, with full results in Section B.5.
Base Large Huge
112M 330M 657M
ETTh1 0.390 0.378 0.391
ETTh2 0.333 0.340 0.339
ETTm1 0.374 0.379 0.383
ETTm2 0.282 0.286 0.284
Avg. 0.345 0.346 0.349
Table 2: MSE of different MAE variants, averaged on four prediction lengths.
Analysis: Impact of MAE’s parameter size
In Table 2 (full results in Section B.6), we observe that the overall performance of three MAE variants (112M, 330M, and 657M) outperforms MoiraiSmall and MoiraiLarge. Particularly, larger models show a slight decrease in performance. This may be due to larger models overfitting image-specific features, reducing their transferability. A similar phenomenon was reported in Moirai, where larger models were found to degrade downstream task performance. We leave the exploration of scaling laws in image-based TSF foundation models for the future.
Analysis: Impact of interpolation and image orientations
We conduct experiments to validate our choices in the Alignment step, detailed in Section B.7. First, we test three different interpolation strategies, which shows that Bilinear interpolation performs best. Second, we apply horizontal and vertical flips on the image to examine whether the assumed left-to-right, top-to-bottom order is efficient. Results show that these changes do not significantly affect performance, suggesting that image reconstruction is isotropic and not influenced by certain orientation.
Computational Cost and Hyperparameter Analysis
Section B.8 compares the computation cost between various baselines, showing that VisionTS are faster than GPT4TS and three Moirai variants. Additionally, Section B.9 illustrates the impact of three hyperparameters: normalization constant $r$, alignment constant $c$, and context length $L$. When $r$ and $c$ are around 0.4, performance is generally well across most benchmarks. For context length, performance generally improves with increasing $L$, but may also show a decrease in some datasets like ETTh2.
4.2 Full-Shot Long-Term Time Series Forecasting
Table 3: Full-short forecasting performance on the Informer long-term TSF benchmark.
Pretrain [Uncaptioned image] Images [Uncaptioned image] Text [Uncaptioned image] No Pretrain
Method VisionTS Time-LLM GPT4TS DLinear PatchTST TimesNet FEDformer Autoformer Stationary ETSformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$ETTh1$ 96 0.347 0.376 0.376 0.402 0.370 0.389 0.375 0.399 0.370 0.399 0.384 0.402 0.376 0.419 0.449 0.459 0.513 0.491 0.494 0.479 0.865 0.713
192 0.385 0.400 0.407 0.421 0.412 0.413 0.405 0.416 0.413 0.421 0.436 0.429 0.420 0.448 0.500 0.482 0.534 0.504 0.538 0.504 1.008 0.792
336 0.407 0.415 0.430 0.438 0.448 0.431 0.439 0.443 0.422 0.436 0.491 0.469 0.459 0.465 0.521 0.496 0.588 0.535 0.574 0.521 1.107 0.809
720 0.439 0.443 0.457 0.468 0.441 0.449 0.472 0.490 0.447 0.466 0.521 0.500 0.506 0.507 0.514 0.512 0.643 0.616 0.562 0.535 1.181 0.865
avg 0.395 0.409 0.418 0.432 0.418 0.421 0.423 0.437 0.413 0.431 0.458 0.450 0.440 0.460 0.496 0.487 0.570 0.537 0.542 0.510 1.040 0.795
$ETTh2$ 96 0.269 0.328 0.286 0.346 0.280 0.335 0.289 0.353 0.274 0.336 0.340 0.374 0.358 0.397 0.346 0.388 0.476 0.458 0.340 0.391 3.755 1.525
192 0.332 0.374 0.361 0.391 0.348 0.380 0.383 0.418 0.339 0.379 0.402 0.414 0.429 0.439 0.456 0.452 0.512 0.493 0.430 0.439 5.602 1.931
336 0.351 0.395 0.390 0.414 0.380 0.405 0.448 0.465 0.329 0.380 0.452 0.452 0.496 0.487 0.482 0.486 0.552 0.551 0.485 0.479 4.721 1.835
720 0.390 0.430 0.405 0.434 0.406 0.436 0.605 0.551 0.379 0.422 0.462 0.468 0.463 0.474 0.515 0.511 0.562 0.560 0.500 0.497 3.647 1.625
avg 0.336 0.382 0.361 0.396 0.354 0.389 0.431 0.447 0.330 0.379 0.414 0.427 0.437 0.449 0.450 0.459 0.526 0.516 0.439 0.452 4.431 1.729
$ETTm1$ 96 0.281 0.322 0.291 0.341 0.300 0.340 0.299 0.343 0.290 0.342 0.338 0.375 0.379 0.419 0.505 0.475 0.386 0.398 0.375 0.398 0.672 0.571
192 0.322 0.353 0.341 0.369 0.343 0.368 0.335 0.365 0.332 0.369 0.374 0.387 0.426 0.441 0.553 0.496 0.459 0.444 0.408 0.410 0.795 0.669
336 0.356 0.379 0.359 0.379 0.376 0.386 0.369 0.386 0.366 0.392 0.410 0.411 0.445 0.459 0.621 0.537 0.495 0.464 0.435 0.428 1.212 0.871
720 0.391 0.413 0.433 0.419 0.431 0.416 0.425 0.421 0.416 0.420 0.478 0.450 0.543 0.490 0.671 0.561 0.585 0.516 0.499 0.462 1.166 0.823
avg 0.338 0.367 0.356 0.377 0.363 0.378 0.357 0.379 0.351 0.381 0.400 0.406 0.448 0.452 0.588 0.517 0.481 0.456 0.429 0.425 0.961 0.734
$ETTm2$ 96 0.169 0.256 0.162 0.248 0.163 0.249 0.167 0.269 0.165 0.255 0.187 0.267 0.203 0.287 0.255 0.339 0.192 0.274 0.189 0.280 0.365 0.453
192 0.225 0.294 0.235 0.304 0.222 0.291 0.224 0.303 0.220 0.292 0.249 0.309 0.269 0.328 0.281 0.340 0.280 0.339 0.253 0.319 0.533 0.563
336 0.278 0.334 0.280 0.329 0.273 0.327 0.281 0.342 0.274 0.329 0.321 0.351 0.325 0.366 0.339 0.372 0.334 0.361 0.314 0.357 1.363 0.887
720 0.372 0.392 0.366 0.382 0.357 0.376 0.397 0.421 0.362 0.385 0.408 0.403 0.421 0.415 0.433 0.432 0.417 0.413 0.414 0.413 3.379 1.338
avg 0.261 0.319 0.261 0.316 0.254 0.311 0.267 0.334 0.255 0.315 0.291 0.333 0.305 0.349 0.327 0.371 0.306 0.347 0.293 0.342 1.410 0.810
$Illness$ 24 2.034 0.937 1.792 0.807 1.869 0.823 2.215 1.081 1.319 0.754 2.317 0.934 3.228 1.260 3.483 1.287 2.294 0.945 2.527 1.020 5.764 1.677
36 1.866 0.888 1.833 0.833 1.853 0.854 1.963 0.963 1.430 0.834 1.972 0.920 2.679 1.080 3.103 1.148 1.825 0.848 2.615 1.007 4.755 1.467
48 1.784 0.870 2.269 1.012 1.886 0.855 2.130 1.024 1.553 0.815 2.238 0.940 2.622 1.078 2.669 1.085 2.010 0.900 2.359 0.972 4.763 1.469
60 1.910 0.912 2.177 0.925 1.877 0.877 2.368 1.096 1.470 0.788 2.027 0.928 2.857 1.157 2.770 1.125 2.178 0.963 2.487 1.016 5.264 1.564
avg 1.899 0.902 2.018 0.894 1.871 0.852 2.169 1.041 1.443 0.798 2.139 0.931 2.847 1.144 3.006 1.161 2.077 0.914 2.497 1.004 5.137 1.544
$Weather$ 96 0.142 0.192 0.155 0.199 0.148 0.188 0.176 0.237 0.149 0.198 0.172 0.220 0.217 0.296 0.266 0.336 0.173 0.223 0.197 0.281 0.300 0.384
192 0.191 0.238 0.223 0.261 0.192 0.230 0.220 0.282 0.194 0.241 0.219 0.261 0.276 0.336 0.307 0.367 0.245 0.285 0.237 0.312 0.598 0.544
336 0.246 0.282 0.251 0.279 0.246 0.273 0.265 0.319 0.245 0.282 0.280 0.306 0.339 0.380 0.359 0.395 0.321 0.338 0.298 0.353 0.578 0.523
720 0.328 0.337 0.345 0.342 0.320 0.328 0.333 0.362 0.314 0.334 0.365 0.359 0.403 0.428 0.419 0.428 0.414 0.410 0.352 0.388 1.059 0.741
avg 0.227 0.262 0.244 0.270 0.227 0.255 0.249 0.300 0.226 0.264 0.259 0.287 0.309 0.360 0.338 0.382 0.288 0.314 0.271 0.334 0.634 0.548
$Traffic$ 96 0.344 0.236 0.392 0.267 0.396 0.264 0.410 0.282 0.360 0.249 0.593 0.321 0.587 0.366 0.613 0.388 0.612 0.338 0.607 0.392 0.719 0.391
192 0.372 0.249 0.409 0.271 0.412 0.268 0.423 0.287 0.379 0.256 0.617 0.336 0.604 0.373 0.616 0.382 0.613 0.340 0.621 0.399 0.696 0.379
336 0.383 0.257 0.434 0.296 0.421 0.273 0.436 0.296 0.392 0.264 0.629 0.336 0.621 0.383 0.622 0.337 0.618 0.328 0.622 0.396 0.777 0.420
720 0.422 0.280 0.451 0.291 0.455 0.291 0.466 0.315 0.432 0.286 0.640 0.350 0.626 0.382 0.660 0.408 0.653 0.355 0.632 0.396 0.864 0.472
avg 0.380 0.256 0.422 0.281 0.421 0.274 0.434 0.295 0.391 0.264 0.620 0.336 0.610 0.376 0.628 0.379 0.624 0.340 0.621 0.396 0.764 0.416
$Electricity$ 96 0.126 0.218 0.137 0.233 0.141 0.239 0.140 0.237 0.129 0.222 0.168 0.272 0.193 0.308 0.201 0.317 0.169 0.273 0.187 0.304 0.274 0.368
192 0.144 0.237 0.152 0.247 0.158 0.253 0.153 0.249 0.157 0.240 0.184 0.289 0.201 0.315 0.222 0.334 0.182 0.286 0.199 0.315 0.296 0.386
336 0.162 0.256 0.169 0.267 0.172 0.266 0.169 0.267 0.163 0.259 0.198 0.300 0.214 0.329 0.231 0.338 0.200 0.304 0.212 0.329 0.300 0.394
720 0.192 0.286 0.200 0.290 0.207 0.293 0.203 0.301 0.197 0.290 0.220 0.320 0.246 0.355 0.254 0.361 0.222 0.321 0.233 0.345 0.373 0.439
avg 0.156 0.249 0.165 0.259 0.170 0.263 0.166 0.264 0.162 0.253 0.193 0.295 0.214 0.327 0.227 0.338 0.193 0.296 0.208 0.323 0.311 0.397
$\bf 1^{st}$ count 46 4 12 0 19 0 0 0 0 0 0
Setups
We evaluate the full-shot capability of each baseline trained on individual long-term TSF benchmarks. In addition to the six datasets used for zero-shot forecasting, we also include the popular Traffic and Illness datasets. As self-attention and feed-forward layers contain rich knowledge that can be transferred to TSF, we choose to fine-tune only the layer normalization (LN) layers while freezing the other parameters, which is also adopted by Zhou et al. (2023). Training details are elaborated in Section C.1.
Main Results
Table 3 summarizes the full-shot results, with standard deviations detailed in Section C.2. It shows that VisionTS outperforms other baselines in most cases (46 out of 80), surpassing the non-pretrained PatchTST and the language-pretrained GPT4TS. Remarkably, except for Illness with the least data, VisionTS demands only a single epoch of fine-tuning. This suggests that even minimal fine-tuning enables VisionTS to effectively adapt to time series. Compared with Table 1, fine-tuning provides limited benefits for ETTh1 and ETTh2 but leads to significant improvements for other datasets. We attribute this to the smaller data scale of ETTh1 and ETTh2.
- w/o VM VM2Attn VM2Trsf Rand-VM
ETTh1 0.395 0.785 0.448 0.459 0.534
ETTh2 0.336 0.420 0.418 0.448 0.411
ETTm1 0.338 0.676 0.397 0.398 0.433
ETTm2 0.261 0.379 0.274 0.292 0.288
Avg. 0.333 0.565 0.384 0.399 0.417
Table 4: MSE results for ablation studies, averaged on four prediction lengths.
All LN Bias MLP Attn Freeze
ETTh1 0.534 0.395 0.401 0.534 0.554 0.419
ETTh2 0.411 0.336 0.347 0.401 0.392 0.340
ETTm1 0.433 0.338 0.343 0.441 0.444 0.374
ETTm2 0.288 0.261 0.256 0.292 0.289 0.305
Avg. 0.417 0.333 0.337 0.417 0.420 0.360
Table 5: MSE for different fine-tuning strategies, averaged on four prediction lengths.
Ablation Study
Tan et al. (2024) proposed several ablation variants for text-based foundation models, including w/o LLM (removing the LLM), LLM2Attn/LLM2Trsf (replacing the LLM with a single self-attention/Transformer layer), and RandLLM (randomly initializing the LLM). They found no significant performance differences and concluded that textual knowledge is not necessary for TSF. We conducted similar ablations to assess the role of the vision model (VM), including w/o VM, VM2Attn, VM2Trsf, and RandVM. Table 5 with full results in Section C.3 shows that these variants lead to worse performance, indicating that visual knowledge is beneficial for TSF.
Analysis: Fine-tuning strategies
As stated before, we fine-tune only the layer normalization (LN). We also tested fine-tuning the bias, MLP, or attention layers, in addition to full fine-tuning and freezing. All hyperparameters were kept constant. Note that freezing differs from the previous zero-shot experiment, where a longer context length was used. Table 5 with full results in Section C.3 show that fine-tuning LN is the best. Modifying MLP or attention layers results in significant performance drops, suggesting that valuable knowledge resides in these components.
5 Related Work
Depending on the pre-training data, TSF foundation models can be categorized into Text-based and TS-based. We firstly review these related work, then introduce recent research for image-based time series analysis.
Text-based TSF Foundation Models
Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks. For example, Zhou et al. (2023) fine-tuned a pre-trained GPT (Radford et al., 2019) on each time-series downstream task, such as forecasting, classification, imputation, and anomaly detection. Based on Llama (Touvron et al., 2023), Jin et al. (2024) froze the pre-trained LLM and reprogrammed the time series to align with the language modality. Bian et al. (2024) adopted a two-stage approach by continually pre-training GPT (Radford et al., 2019) on the time-series domain. Nevertheless, the TSF performance of LLMs has recently been questioned by Tan et al. (2024), which designed several ablation studies to show that textual knowledge is unnecessary for forecasting. In this paper, we attribute it to the large modality gap. Some recent approaches focus on directly transforming the time series into natural texts for LLMs, allowing for zero-shot forecasting. For example, PromptCast (Xue & Salim, 2023) uses pre-defined templates to describe numerical time series data, while LLMTime (Gruver et al., 2023) directly separates time steps using commas and separates digits using spaces to construct the text input. However, due to the efficiency issue caused by the autoregressive decoding strategy, their practical use is limited.
Time Series-Based TSF Foundation Models
Self-supervised pre-training a TSF model on the same dataset used for downstream TSF tasks is a well-explored topic (Ma et al., 2023; Zhang et al., 2024), such as denoising autoencoders (Zerveas et al., 2021) or contrastive learning (Woo et al., 2022a; Yue et al., 2022). They follow a similar paradigm to the masked autoencoder (MAE) in computer vision. However, these methods rarely examine the cross-dataset generalization capabilities. Recently, research has shifted towards training universal foundation models, by collecting large-scale time series datasets from diverse domains (Goswami et al., 2024; Liu et al., 2024; Das et al., 2024; Dong et al., 2024; Feng et al., 2024) or generating numerous synthetic time series data (Yang et al., 2024). As a representative method, Woo et al. (2024) collected 27 billion observations across nine domains and trained TSF foundation models of various scales, achieving strong zero-shot performance. However, given the severe heterogeneity, constructing high-quality large datasets poses significant challenges for building these foundation models.
Image-Based Time-Series Analysis
Previous research has investigated encoding time series data into images and used convolutional neural networks (CNNs) trained from scratch for classification (Wang & Oates, 2015a; b; Hatami et al., 2018) or forecasting (Li et al., 2020; Sood et al., 2021; Semenoglou et al., 2023). Recent researchers explored using pre-trained models for these imaging time series. Li et al. (2024) used a pre-trained vision transformer (ViT) for classification. Wimmer & Rekabsaz (2023) and Zhang et al. (2023) employed vision-language multimodal pre-trained models to extract predictive features and generate text descriptions. However, these studies did not deeply examine the transferability from natural images to TSF. Despite early efforts by Zhou et al. (2023) to fine-tune a BEiT (Bao et al., 2022) trained on images for time series forecasting, it still falls short of the leading text-based and TS-based TSF foundation models. To the best of our knowledge, we are the first to show that an image-based foundation model, without further time-series adaptation, can match or even surpass other types of TSF foundation models.
6 Conclusion
In this paper, we explore a novel approach to building a time series forecasting (TSF) foundation model using natural images, offering a new perspective distinct from the traditional text-based and TS-based methods. By leveraging the intrinsic similarities between images and time series, we introduced VisionTS, an MAE-based TSF foundation model that reformulates the TSF task as an image reconstruction problem. Our extensive evaluations demonstrate that VisionTS achieves outstanding forecasting performance in zero-shot and full-shot settings, being a free lunch for a TSF foundation model. We hope our findings could open new avenues for further cross-domain research.
Limitations and Future Directions.
As a preliminary study, we employed a basic MAE model and reformulated TSF as a patch-based image reconstruction task. Reformulating TSF as a more refined image inpainting task and utilizing more advanced models like diffusion models (Rombach et al., 2022; Peebles & Xie, 2023) presents a promising research direction.
Appendix A Details of Experiments
Table 6: The statistics of 8 long-term TSF datasets.
Dataset ETTh1 ETTh2 ETTm1 ETTm2 Electricity Traffic Weather Illness
Variates 7 7 7 7 321 862 21 7
Timesteps 17,420 17,420 69,680 69,680 26,304 17,544 52,695 966
Frequency 1hour 1hour 15min 15min 1hour 1hour 10min 1week
Periodicity 24 24 96 96 24 24 144 52
Long-term TSF Benchmark
We evaluate our model on 8 widely used long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021), summarized in Table 6. We also report the periodicity of each dataset for our segmentation step.
  1. 1.
    ETTh1/ETTh2/ETTm1/ETTm2 record 7 indicators from electricity transformers between July 2016 and July 2018. Data is collected hourly for ETTh1/ETTh2 and every 15 minutes for ETTm1/ETTm2.
  2. 2.
    Electricity contains hourly electricity consumption (in KWh) data from 321 customers between 2012 and 2014.
  3. 3.
    Traffic provides hourly road occupancy rates measured by various sensors on San Francisco Bay area freeways over 2 years, sourced from the California Department of Transportation.
  4. 4.
    Illness tracks 7 weekly indicators related to influenza-like illness from the CDC in the United States, spanning 2002 to 2021.
  5. 5.
    Weather includes 21 meteorological indicators, such as temperature and humidity, recorded every 10 minutes throughout 2020.
Monash Benchmark
Following Woo et al. (2024), we tested 29 Monash datasets (Godahewa et al., 2021) using GluonTS (Alexandrov et al., 2020), including M1 Monthly, M3 Monthly, M3 Other, M4 Monthly, M4 Weekly, M4 Daily, M4 Hourly, Tourism Quarterly, Tourism Monthly, CIF 2016, Australian Electricity Demand, Bitcoin, Pedestrian Counts, Vehicle Trips, KDD Cup, Weather, NN5 Daily, NN5 Weekly, Carparts, FRED-MD, Traffic Hourly, Traffic Weekly, Rideshare, Hospital, COVID Deaths, Temperature Rain, Sunspot, Saugeen River Flow, and US Births.
Baselines
The baseline models selected for comparison are briefly described below:
  1. 1.
    Moirai (Woo et al., 2024) is a TSF foundation model trained on the Large-scale Open Time Series Archive (LOTSA), with over 27B observations across nine domains. It has three variants: small, base, and large.
  2. 2.
    TimesFM (Das et al., 2024) is a decoder-style TSF foundation model, using a large time-series corpus comprising both real-world and synthetic datasets.
  3. 3.
    Time-LLM (Jin et al., 2024) is a text-based TSF foundation model built on Llama, which reprograms time series data to align with the language modality, keeping the LLM frozen.
  4. 4.
    GPT4TS (Zhou et al., 2023) (OneFitsAll) is another text-based model based on GPT, fine-tuned for forecasting tasks.
  5. 5.
    LLMTime (Gruver et al., 2023) encodes time series data to a text sequence, supporting zero-shot forecasting.
  6. 6.
    DLinear (Zeng et al., 2023) proposes a linear forecasting model, enhanced by seasonal-trend decomposition or normalization.
  7. 7.
    PatchTST (Nie et al., 2022) uses Transformer encoders with patching and channel independence techniques for improved predictions.
  8. 8.
    TimesNet (Wu et al., 2023) applies convolution kernels along the time dimension, using temporal decomposition and periodical segmentation to capture temporal patterns.
  9. 9.
    FEDformer (Zhou et al., 2022) employs a sparse frequency domain representation, using frequency-enhanced blocks for cross-time dependency.
  10. 10.
    Autoformer (Wu et al., 2021) uses series decomposition blocks and Auto-Correlation to capture cross-time dependency.
  11. 11.
    Stationary (Liu et al., 2022) introduces stationarization and de-stationary attention mechanisms.
  12. 12.
    ETSFormer (Woo et al., 2022b) leverages exponential smoothing principles, including exponential smoothing and frequency attention mechanisms.
  13. 13.
    Informer (Zhou et al., 2021) proposes ProbSparse self-attention and distillation operations.
For the long-term TSF benchmark, we include TS-based foundation model results from their original papers, Text-based model results from Tan et al. (2024), and other baseline results from Zhou et al. (2023). For the Monash benchmark, we include results from Woo et al. (2024).
Metrics and Environment
Performance is assessed using Mean Squared Error (MSE) and Mean Absolute Error (MAE), with lower values indicating better forecasting accuracy. All experiments are conducted using Time-Series-Library (https://github.com/thuml/Time-Series-Library) on an NVIDIA A800 GPU.
Appendix B Zero-Shot Forecasting
B.1 Hyperparameters
Table 7: Hyperparameters for VisionTS used in our zero-shot forecasting (Informer benchmark).
ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
Normalization constant $r$ 0.4 0.4 0.4 0.4 0.4 0.4
Alignment constant $c$ 0.4 0.4 0.4 0.4 0.4 0.4
Context length $L$ 2880 1728 2304 4032 4032 2880
We conduct hyperparameter tuning on validation sets to determine the optimal context length $L$. Final used hyperparameters are summarized in Table 7.
B.2 Full forecasting results of the long-term TSF benchmark
Table 8: Full results of Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Bold: the best result.
[Uncaptioned image] Zero-Shot [Uncaptioned image] Few-Shot (10% Downstream Dataset)
Pretrain [Uncaptioned image] Images [Uncaptioned image] Time-series [Uncaptioned image] Text [Uncaptioned image] No Pretrain
Method VisionTime MoiraiS MoiraiB MoiraiL TimeLLM GPT4TS DLinear PatchTST TimesNet Autoformer Informer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$ETTh1$ 96 0.353 0.383 0.375 0.402 0.384 0.402 0.380 0.398 0.448 0.460 0.458 0.456 0.492 0.495 0.516 0.485 0.861 0.628 0.613 0.552 1.179 0.792
192 0.392 0.410 0.399 0.419 0.425 0.429 0.440 0.434 0.484 0.483 0.570 0.516 0.565 0.538 0.598 0.524 0.797 0.593 0.722 0.598 1.199 0.806
336 0.407 0.423 0.412 0.429 0.456 0.450 0.514 0.474 0.589 0.540 0.608 0.535 0.721 0.622 0.657 0.550 0.941 0.648 0.750 0.619 1.202 0.811
720 0.406 0.441 0.413 0.444 0.470 0.473 0.705 0.568 0.700 0.604 0.725 0.591 0.986 0.743 0.762 0.610 0.877 0.641 0.721 0.616 1.217 0.825
avg 0.390 0.414 0.400 0.424 0.434 0.439 0.510 0.469 0.556 0.522 0.590 0.525 0.691 0.600 0.633 0.542 0.869 0.628 0.702 0.596 1.199 0.809
$ETTh2$ 96 0.271 0.328 0.281 0.334 0.277 0.327 0.287 0.325 0.275 0.326 0.331 0.374 0.357 0.411 0.353 0.389 0.378 0.409 0.413 0.451 3.837 1.508
192 0.328 0.367 0.340 0.373 0.340 0.374 0.347 0.367 0.374 0.373 0.402 0.411 0.569 0.519 0.403 0.414 0.490 0.467 0.474 0.477 3.856 1.513
336 0.345 0.381 0.362 0.393 0.371 0.401 0.377 0.393 0.406 0.429 0.406 0.433 0.671 0.572 0.426 0.441 0.537 0.494 0.547 0.543 3.952 1.526
720 0.388 0.422 0.380 0.416 0.394 0.426 0.404 0.421 0.427 0.449 0.449 0.464 0.824 0.648 0.477 0.480 0.510 0.491 0.516 0.523 3.842 1.503
avg 0.333 0.375 0.341 0.379 0.346 0.382 0.354 0.377 0.370 0.394 0.397 0.421 0.605 0.538 0.415 0.431 0.479 0.465 0.488 0.499 3.872 1.513
$ETTm1$ 96 0.341 0.347 0.404 0.383 0.335 0.360 0.353 0.363 0.346 0.388 0.390 0.404 0.352 0.392 0.410 0.419 0.583 0.501 0.774 0.614 1.162 0.785
192 0.360 0.360 0.435 0.402 0.366 0.379 0.376 0.380 0.373 0.416 0.429 0.423 0.382 0.412 0.437 0.434 0.630 0.528 0.754 0.592 1.172 0.793
336 0.377 0.374 0.462 0.416 0.391 0.394 0.399 0.395 0.413 0.426 0.469 0.439 0.419 0.434 0.476 0.454 0.725 0.568 0.869 0.677 1.227 0.908
720 0.416 0.405 0.490 0.437 0.434 0.419 0.432 0.417 0.485 0.476 0.569 0.498 0.490 0.477 0.681 0.556 0.769 0.549 0.810 0.630 1.207 0.797
avg 0.374 0.372 0.448 0.410 0.382 0.388 0.390 0.389 0.404 0.427 0.464 0.441 0.411 0.429 0.501 0.466 0.677 0.537 0.802 0.628 1.192 0.821
$ETTm2$ 96 0.228 0.282 0.205 0.282 0.195 0.269 0.189 0.260 0.177 0.261 0.188 0.269 0.213 0.303 0.191 0.274 0.212 0.285 0.352 0.454 3.203 1.407
192 0.262 0.305 0.261 0.318 0.247 0.303 0.247 0.300 0.241 0.314 0.251 0.309 0.278 0.345 0.252 0.317 0.270 0.323 0.694 0.691 3.112 1.387
336 0.293 0.328 0.319 0.355 0.291 0.333 0.295 0.334 0.274 0.327 0.307 0.346 0.338 0.385 0.306 0.353 0.323 0.353 2.408 1.407 3.255 1.421
720 0.343 0.370 0.415 0.410 0.355 0.377 0.372 0.386 0.417 0.390 0.426 0.417 0.436 0.440 0.433 0.427 0.474 0.449 1.913 1.166 3.909 1.543
avg 0.282 0.321 0.300 0.341 0.272 0.321 0.276 0.320 0.277 0.323 0.293 0.335 0.316 0.368 0.296 0.343 0.320 0.353 1.342 0.930 3.370 1.440
$Electricity$ 96 0.177 0.266 0.205 0.299 0.158 0.248 0.152 0.242 0.139 0.241 0.139 0.237 0.150 0.253 0.140 0.238 0.299 0.373 0.261 0.348 1.259 0.919
192 0.188 0.277 0.220 0.310 0.174 0.263 0.171 0.259 0.151 0.248 0.156 0.252 0.164 0.264 0.160 0.255 0.305 0.379 0.338 0.406 1.160 0.873
336 0.207 0.296 0.236 0.323 0.191 0.278 0.192 0.278 0.169 0.270 0.175 0.270 0.181 0.282 0.180 0.276 0.319 0.391 0.410 0.474 1.157 0.872
720 0.256 0.337 0.270 0.347 0.229 0.307 0.236 0.313 0.240 0.322 0.233 0.317 0.223 0.321 0.241 0.323 0.369 0.426 0.715 0.685 1.203 0.898
avg 0.207 0.294 0.233 0.320 0.188 0.274 0.188 0.273 0.175 0.270 0.176 0.269 0.180 0.280 0.180 0.273 0.323 0.392 0.431 0.478 1.195 0.891
$Weather$ 96 0.220 0.257 0.173 0.212 0.167 0.203 0.177 0.208 0.161 0.210 0.163 0.215 0.171 0.224 0.165 0.215 0.184 0.230 0.221 0.297 0.374 0.401
192 0.244 0.275 0.216 0.250 0.209 0.241 0.219 0.249 0.204 0.248 0.210 0.254 0.215 0.263 0.210 0.257 0.245 0.283 0.270 0.322 0.552 0.478
336 0.280 0.299 0.260 0.282 0.256 0.276 0.277 0.292 0.261 0.302 0.256 0.292 0.258 0.299 0.259 0.297 0.305 0.321 0.320 0.351 0.724 0.541
720 0.330 0.337 0.320 0.322 0.321 0.323 0.365 0.350 0.309 0.332 0.321 0.339 0.320 0.346 0.332 0.346 0.381 0.371 0.390 0.396 0.739 0.558
avg 0.269 0.292 0.242 0.267 0.238 0.261 0.260 0.275 0.234 0.273 0.238 0.275 0.241 0.283 0.242 0.279 0.279 0.301 0.300 0.342 0.597 0.495
Average 0.309 0.345 0.327 0.357 0.310 0.344 0.329 0.350 0.336 0.368 0.360 0.378 0.407 0.416 0.378 0.389 0.491 0.446 0.678 0.579 1.904 0.995
$\bf 1^{st}$ count 32 0 10 8 10 6 0 0 0 0 0
Table 8 shows the full results of zero-shot/few-shot long-term forecasting performance. VisionTS achieves the best results in most cases (32 out of 62), outperforming MoiraiBase (10 out of 62) and MoiraiLarge (8 out of 62).
B.3 Comparison of TimesFM and LLMTime
Table 9: Comparison of TimesFM and LLMTime for zero-shot forecasting in terms of MAE. Owing to expensive evaluations for LLMTime, the results are reported on the last test window of the original test split, as done in Das et al. (2024); Gruver et al. (2023).
Method VisionTS TimesFM LLMTime
ETTh1 96 0.35 0.45 0.42
192 0.45 0.53 0.50
ETTh2 96 0.24 0.35 0.33
192 0.60 0.62 0.70
ETTm1 96 0.12 0.19 0.37
192 0.23 0.26 0.71
ETTm2 96 0.19 0.24 0.29
192 0.24 0.27 0.31
Average 0.30 0.36 0.45
Table 10: Comparison of traditional zero-shot forecasting baselines.
Method VisionTS ETS ARIMA Seasonal Naïve Seasonal Avg
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$ETTh1$ 96 0.353 0.383 1.289 0.710 0.900 0.719 0.512 0.433 0.589 0.585
192 0.392 0.410 1.319 0.730 0.906 0.724 0.581 0.469 0.598 0.590
336 0.407 0.423 1.324 0.742 0.908 0.731 0.650 0.501 0.610 0.597
720 0.406 0.441 1.329 0.751 0.932 0.753 0.655 0.514 0.656 0.624
avg 0.390 0.414 1.315 0.733 0.912 0.732 0.600 0.479 0.613 0.599
$ETTh2$ 96 0.271 0.328 0.399 0.408 0.488 0.508 0.391 0.380 0.457 0.494
192 0.328 0.367 0.500 0.459 0.497 0.514 0.482 0.429 0.466 0.500
336 0.345 0.381 0.562 0.498 0.507 0.522 0.532 0.466 0.476 0.509
720 0.388 0.422 0.558 0.506 0.572 0.557 0.525 0.474 0.542 0.548
avg 0.333 0.375 0.505 0.468 0.516 0.525 0.483 0.437 0.485 0.513
$ETTm1$ 96 0.341 0.347 1.204 0.659 0.702 0.568 0.423 0.387 0.369 0.399
192 0.360 0.360 1.251 0.685 0.704 0.570 0.463 0.406 0.374 0.402
336 0.377 0.374 1.276 0.702 0.709 0.574 0.496 0.426 0.382 0.407
720 0.416 0.405 1.311 0.724 0.713 0.580 0.574 0.464 0.394 0.416
avg 0.374 0.372 1.261 0.693 0.707 0.573 0.489 0.421 0.380 0.406
$ETTm2$ 96 0.228 0.282 0.257 0.324 0.397 0.434 0.263 0.301 0.365 0.411
192 0.262 0.305 0.331 0.366 0.402 0.436 0.321 0.337 0.369 0.414
336 0.293 0.328 0.402 0.406 0.407 0.439 0.376 0.370 0.375 0.418
720 0.343 0.370 0.512 0.462 0.413 0.443 0.471 0.422 0.380 0.423
avg 0.282 0.321 0.376 0.390 0.405 0.438 0.358 0.357 0.372 0.417
Average 0.344 0.370 0.864 0.571 0.635 0.567 0.482 0.424 0.463 0.484
$\bf 1^{st}$ count 41 0 0 0 1
Due to the step-by-step output of the decoder architecture, the efficiency of TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023) are relatively slower. Thus, Das et al. (2024) only reported results for the last test window of the original split. We compared VisionTS with their results under the same setting, as shown in Table 9. VisionTS outperforms TimesFM and LLMTime in terms of MAE, indicating that image-based TSF models are on par with or even better than TS-based and text-based models.
B.4 Comparison of traditional methods
In addition to deep learning models, we also compare traditional methods, including ARIMA, ETS, and two methods that require periodicity as our VisionTS: Seasonal Naïve (repeating the last period) and Seasonal Avg (similar to Seasonal Naïve but repeating the average of all periods in the look-back window). Due to the high computational cost of ARIMA and ETS, we only compare them on the small-scale benchmarks, i.e., four ETT datasets. Table 10 shows that VisionTS also achieves the best performance.
B.5 Full forecasting results of the Monash TSF benchmark
Table 11: The statistics of 29 Monash TSF datasets. $P=1$ indicates no significant periodicity.
Datasets Frequency Periodiciy ($P$)
Rideshare H 1
M4 Daily, Bitcoin, COVID Deaths D 1
M4 Weekly, NN5 Weekly, Traffic Weekly W 1
FRED-MD M 1
M3 Other Q 2
CIF 2016 (horizon=6) M 3
Tourism Quarterly Q 4
M3 Monthly, M4 Monthly, CIF 2016 (horizon=12), Car Parts M 6
Vehicle Trips, NN5 Daily D 7
M1 Monthly, Tourism Monthly, Hospital M 12
M4 Hourly, Traffic Hourly H 24
Sunspot, Saugeen River Flow D 30
Pedestrian Counts H 48
KDD cup H 96
Aus. Elec. Demand 30T 336
Weather, Temperature Rain, US Births D 365
Table 12: Full results of Fig. 4: Forecasting results (MAE) on the Monash TSF benchmark. We reported the reproduction results of LLMTime based on the GPT3.5 API from Woo et al. (2024).
VisionTS LLMTime MoiraiSmall Naive SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer
M1 Monthly 1987.69 2562.84 2082.26 2707.75 2259.04 2166.18 2237.5 1905.28 2080.13 2088.25 2052.32 2162.58 1860.81 1820.37 2184.42 2723.88
M3 Monthly 741 877.97 713.41 837.14 743.41 623.71 630.59 626.46 654.8 692.97 732 692.48 728.81 648.6 699.3 798.38
M3 Other 283.76 300.3 263.54 278.43 277.83 215.35 189.42 194.98 193.02 234.43 318.13 240.17 247.56 221.85 245.29 239.24
M4 Monthly 654.27 728.27 597.6 671.27 625.24 563.58 589.52 582.6 575.36 596.19 611.69 612.52 615.22 578.48 655.51 780.47
M4 Weekly 314.83 518.44 339.76 347.99 336.82 333.32 296.15 335.66 321.61 293.21 364.65 338.37 351.78 277.73 359.46 378.89
M4 Daily 215.63 266.52 189.1 180.83 178.27 178.86 176.6 193.26 179.67 181.92 231.36 177.91 299.79 190.44 189.47 201.08
M4 Hourly 288.37 576.06 268.04 1218.06 1218.06 1220.97 386.27 3358.1 1310.85 257.39 285.35 385.49 886.02 425.75 393.63 320.54
Tourism Quarterly 12931.88 16918.9 18352.44 15845.1 15014.19 7656.49 9972.42 8925.52 10475.5 9092.58 10268 8981.04 9511.37 8640.56 9137.12 9521.67
Tourism Monthly 2560.19 5608.61 3569.85 5636.83 5302.1 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98
CIF 2016 570907.24 599314 655888.58 578597 581875.97 714819 855578 642421 469059 563205.6 603551 1495923 3200418 679035 5998225 4057973
Aus. Elec. Demand 237.44 760.81 266.57 659.6 659.6 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.5 231.45
Bitcoin 2.34E+18 1.7E+18 1.76E+18 7.8E+17 5.33E+18 5.3E+18 9.9E+17 1.1E+18 3.6E+18 6.66E+17 1.9E+18 1.5E+18 2E+18 1.1E+18 2.5E+18 2.6E+18
Pedestrian Counts 46.37 97.77 54.88 170.88 170.87 170.94 222.38 216.5 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29
Vehicle Trips 22.08 31.48 24.46 31.42 29.98 30.76 21.21 30.95 30.07 27.24 22.61 22.93 22 28.16 24.15 28.01
KDD cup 35.13 42.72 39.81 42.13 42.04 42.06 39.2 44.88 52.2 36.85 34.82 37.16 48.98 49.1 37.08 44.46
Weather 2.07 2.17 1.96 2.36 2.24 2.51 2.3 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03
NN5 Daily 3.52 7.1 5.37 8.26 6.63 3.8 3.7 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16
NN5 Weekly 14.67 15.76 15.07 16.71 15.66 15.3 14.98 15.7 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34
Carparts 0.58 0.44 0.53 0.65 0.55 0.53 0.58 0.56 0.56 0.41 0.53 0.39 0.39 0.98 0.4 0.39
FRED-MD 1893.67 2804.64 2568.48 2825.67 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.8 2508.4 4666.04
Traffic Hourly 0.01 0.03 0.02 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01
Traffic Weekly 1.2 1.15 1.17 1.19 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.2 1.42
Rideshare 4.6 6.28 1.35 6.29 6.29 7.62 6.45 6.29 3.37 6.3 6.07 6.59 6.28 5.55 2.75 6.29
Hospital 19.36 25.68 23 24.07 21.76 18.54 17.43 17.97 19.6 19.24 19.17 22.86 18.25 20.18 19.35 36.19
COVID Deaths 137.51 653.31 124.32 353.71 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66
Temperature Rain 5.88 6.37 5.3 9.39 8.18 8.22 7.14 8.21 7.19 6.13 6.76 5.56 5.37 7.28 5.81 5.24
Sunspot 2.81 5.07 0.11 3.93 4.93 4.93 2.57 4.93 2.57 3.83 2.27 7.97 0.77 14.47 0.17 0.13
Saugeen River Flow 24.47 34.84 24.07 21.5 21.5 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06
US Births 820.21 1374.99 872.51 1152.67 1192.2 586.93 399 419.73 526.33 574.93 441.7 557.87 424.93 422 504.4 452.87
Normalized MAE 0.714 1.041 0.657 1.000 1.028 0.927 0.758 0.872 0.898 0.785 0.760 0.741 0.759 0.783 0.749 0.770
Rank 2 16 1 14 15 13 5 11 12 10 7 3 6 9 4 8
Setup
Table 11 lists the sampling frequency and the selected period $P$ for each dataset. Datasets with $P=1$ indicate no significant periodicity, where we use a context length of $L=300$. For other datasets with $P>1$, we select a longer context length of $L=1000$. All datasets were tested with the hyperparameters $r=c=0.4$ as we had done for the long-term TSF benchmark.
Results
Table 12 presents VisionTS ’s MAE test results, with the normalized MAE calculated by dividing each dataset’s MAE by the naive forecast’s MAE and aggregated using the geometric mean across datasets. We include the result of each baseline from Woo et al. (2024). Particularly, we find that VisionTS outperforms Moirai on some datasets with $P=1$ (e.g., FRED-MD and M4 Weekly), showing that VisionTS can still work effectively without significant periodicity.
B.6 Impact of MAE’s parameter size
Table 13: Full results of Table 2: zero-shot forecasting results of different MAE variants. Bold: best results among three variants. We also include the results from Moirai for reference.
Method MAE (Base) MAE (Large) MAE (Huge) Moirai (Small) Moirai (Base) Moirai (Huge)
112M 330M 657M 14M 91M 311M
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
$ETTh1$ 96 0.353 0.383 0.346 0.382 0.362 0.384 0.375 0.402 0.384 0.402 0.380 0.398
192 0.392 0.410 0.379 0.406 0.407 0.414 0.399 0.419 0.425 0.429 0.440 0.434
336 0.407 0.423 0.391 0.416 0.399 0.419 0.412 0.429 0.456 0.450 0.514 0.474
720 0.406 0.441 0.397 0.433 0.395 0.433 0.413 0.444 0.470 0.473 0.705 0.568
avg 0.390 0.414 0.378 0.409 0.391 0.412 0.400 0.424 0.434 0.439 0.510 0.469
$ETTh2$ 96 0.271 0.328 0.286 0.334 0.285 0.333 0.281 0.334 0.277 0.327 0.287 0.325
192 0.328 0.367 0.346 0.375 0.337 0.369 0.340 0.373 0.340 0.374 0.347 0.367
336 0.345 0.381 0.356 0.387 0.357 0.388 0.362 0.393 0.371 0.401 0.377 0.393
720 0.388 0.422 0.371 0.409 0.379 0.412 0.380 0.416 0.394 0.426 0.404 0.421
avg 0.333 0.375 0.340 0.377 0.339 0.375 0.341 0.379 0.346 0.382 0.354 0.377
$ETTm1$ 96 0.341 0.347 0.344 0.349 0.352 0.351 0.404 0.383 0.335 0.360 0.353 0.363
192 0.360 0.360 0.365 0.363 0.360 0.367 0.435 0.402 0.366 0.379 0.376 0.380
336 0.377 0.374 0.381 0.376 0.381 0.383 0.462 0.416 0.391 0.394 0.399 0.395
720 0.416 0.405 0.429 0.411 0.440 0.412 0.490 0.437 0.434 0.419 0.432 0.417
avg 0.374 0.372 0.379 0.375 0.383 0.378 0.448 0.410 0.382 0.388 0.390 0.389
$ETTm2$ 96 0.228 0.282 0.225 0.282 0.229 0.282 0.205 0.282 0.195 0.269 0.189 0.260
192 0.262 0.305 0.262 0.305 0.265 0.306 0.261 0.318 0.247 0.303 0.247 0.300
336 0.293 0.328 0.299 0.331 0.286 0.324 0.319 0.355 0.291 0.333 0.295 0.334
720 0.343 0.370 0.358 0.377 0.355 0.374 0.415 0.410 0.355 0.377 0.372 0.386
avg 0.282 0.321 0.286 0.324 0.284 0.322 0.300 0.341 0.272 0.321 0.276 0.320
$Electricity$ 96 0.177 0.266 0.177 0.268 0.170 0.259 0.205 0.299 0.158 0.248 0.152 0.242
192 0.188 0.277 0.192 0.283 0.182 0.273 0.220 0.310 0.174 0.263 0.171 0.259
336 0.207 0.296 0.213 0.303 0.207 0.295 0.236 0.323 0.191 0.278 0.192 0.278
720 0.256 0.337 0.256 0.337 0.250 0.333 0.270 0.347 0.229 0.307 0.236 0.313
avg 0.207 0.294 0.209 0.298 0.202 0.290 0.233 0.320 0.188 0.274 0.188 0.273
$Weather$ 96 0.220 0.257 0.222 0.257 0.235 0.265 0.173 0.212 0.167 0.203 0.177 0.208
192 0.244 0.275 0.246 0.275 0.276 0.288 0.216 0.250 0.209 0.241 0.219 0.249
336 0.280 0.299 0.283 0.301 0.304 0.309 0.260 0.282 0.256 0.276 0.277 0.292
720 0.330 0.337 0.338 0.343 0.351 0.350 0.320 0.322 0.321 0.323 0.365 0.350
avg 0.269 0.292 0.272 0.294 0.292 0.303 0.242 0.267 0.238 0.261 0.260 0.275
Average 0.309 0.345 0.311 0.346 0.315 0.347 0.327 0.357 0.310 0.344 0.329 0.350
$\bf 1^{st}$ count 38 17 17 - - -
Table 13 compares zero-shot forecasting performance of three MAE variants (112M, 330M, and 657M), showing that the three variants are similar, but larger models show a slight decrease. Particularly, the smallest model excels in ETTh2, ETTm1, ETTm2, and Weather, while the largest model excels in Electricity.
B.7 Impact of the different image encoding strategies
Table 14: Impact of resampling filters and image orientations.
Interpolation strategies in resampling Image orientation
Method Bilinear Bicubic Nearest Neighbor Method - Horizontal flip Vertical flip
Metric MSE MAE MSE MAE MSE MAE Metric MSE MAE MSE MAE MSE MAE
$ETTh1$ 96 0.353 0.383 0.351 0.383 0.426 0.424 $ETTh1$ 96 0.353 0.383 0.348 0.379 0.355 0.385
192 0.392 0.410 0.392 0.409 0.450 0.443 192 0.392 0.410 0.386 0.404 0.394 0.411
336 0.407 0.423 0.407 0.422 0.451 0.450 336 0.407 0.423 0.401 0.416 0.408 0.423
720 0.406 0.441 0.405 0.440 0.454 0.470 720 0.406 0.441 0.399 0.430 0.406 0.442
avg 0.390 0.414 0.389 0.414 0.445 0.446 avg 0.390 0.414 0.384 0.407 0.391 0.415
$ETTh2$ 96 0.271 0.328 0.274 0.329 0.298 0.349 $ETTh2$ 96 0.271 0.328 0.274 0.329 0.274 0.330
192 0.328 0.367 0.330 0.367 0.343 0.380 192 0.328 0.367 0.331 0.370 0.330 0.367
336 0.345 0.381 0.345 0.380 0.373 0.401 336 0.345 0.381 0.347 0.386 0.345 0.381
720 0.388 0.422 0.386 0.419 0.404 0.431 720 0.388 0.422 0.376 0.416 0.388 0.422
avg 0.333 0.375 0.334 0.374 0.354 0.390 avg 0.333 0.375 0.332 0.375 0.334 0.375
$ETTm1$ 96 0.341 0.347 0.366 0.354 0.399 0.374 $ETTm1$ 96 0.341 0.347 0.345 0.348 0.342 0.347
192 0.360 0.360 0.383 0.367 0.397 0.376 192 0.360 0.360 0.364 0.362 0.360 0.360
336 0.377 0.374 0.396 0.381 0.386 0.380 336 0.377 0.374 0.378 0.375 0.377 0.374
720 0.416 0.405 0.429 0.409 0.417 0.409 720 0.416 0.405 0.419 0.408 0.417 0.405
avg 0.374 0.372 0.393 0.378 0.400 0.384 avg 0.374 0.372 0.376 0.373 0.374 0.372
$ETTm2$ 96 0.228 0.282 0.246 0.296 0.264 0.326 $ETTm2$ 96 0.228 0.282 0.230 0.286 0.228 0.283
192 0.262 0.305 0.273 0.313 0.273 0.328 192 0.262 0.305 0.264 0.308 0.262 0.305
336 0.293 0.328 0.303 0.334 0.297 0.343 336 0.293 0.328 0.298 0.332 0.293 0.328
720 0.343 0.370 0.343 0.370 0.334 0.369 720 0.343 0.370 0.350 0.373 0.343 0.369
avg 0.282 0.321 0.291 0.328 0.292 0.341 avg 0.282 0.321 0.285 0.325 0.282 0.321
Average 0.344 0.370 0.352 0.373 0.373 0.391 Average 0.344 0.370 0.344 0.370 0.345 0.371
$\bf 1^{st}$ count 30 18 2 $\bf 1^{st}$ count 28 16 21
Table 15: Computational cost in terms of seconds for forecasting a batch of 32 time series data.
Context Length 1k 2k 3k 4k 1k
Prediction Length 1k 1k 2k 3k 4k
PatchTST 0.01 0.01 0.01 0.02 0.01 0.01 0.01 0.02
GPT4TS 0.03 0.05 0.09 0.12 0.03 0.03 0.03 0.03
MoiraiSmall 0.04 0.05 0.06 0.07 0.04 0.05 0.06 0.07
MoiraiBase 0.06 0.08 0.11 0.13 0.06 0.08 0.11 0.13
MoiraiLarge 0.14 0.19 0.25 0.30 0.14 0.19 0.25 0.30
VisionTS ($c=1.0$) 0.03 0.03 0.03 0.03 0.02 0.02 0.02 0.02
VisionTS ($c=0.4$) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02
Table 14 summarizes the impact of interporation strategies and image orientations in the Alignment step. It shows that the smoother Bilinear and Bicubic interporation perform similarly, both significantly better than the rougher Nearest Neighbor. This suggests that smooth resizing effectively handles time series interpolation. Moreover, image orientation has little impact on performance.
B.8 Computation cost
We evaluate the computation cost of different baselines on an NVIDIA A800 GPU. For VisionTS, we test two variants, $c=0.4$ and $c=1.0$, since the alignment constant $c$ influences the number of visible patches and determines the computation cost of MAE encoder. Results are averaged on 20 runs.
Table 15 compares the computation cost between various baselines, showing that VisionTS are faster than GPT4TS and three Moirai variants. While computation time increases with context length for all the other Transformer-based baselines, VisionTS remains nearly constant. This is because VisionTS encodes input sequences into an image with constant size, ensuring $O(1)$ efficiency. In contrast, Transformer-based methods operate at $O(L^{2})$ relative to context length $L$.
B.9 Hyperparameter analysis
Figs. 5, 6 and 7 show the influence of three hyperparameters, $r$, $c$, and $L$. We report the MSE averaged on four prediction lengths {96, 192, 336, 720}.
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) ETTm2
(e) Electricity
(f) Weather
Figure 5: MSE (Y-axis) performance of different normalization constants $r$ (X-axis).
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) ETTm2
(e) Electricity
(f) Weather
Figure 6: MSE (Y-axis) performance of different alignment constants $c$ (X-axis).
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) ETTm2
(e) Electricity
(f) Weather
Figure 7: MSE (Y-axis) performance of different context lengths $L$ (X-axis).
Appendix C Full-Shot Forecasting
C.1 Training details
Table 16: Final hyperparameters for VisionTS used in our full-shot forecasting.
ETTh1 ETTh2 ETTm1 ETTm2 Illness Weather Traffic Electricity
Normalization constant $r$ 0.4 0.4 0.4 0.4 1.0 1.0 0.4 0.4
Alignment constant $c$ 0.4 0.4 0.4 0.4 0.4 0.7 0.4 0.4
Context length $L$ 1152 1152 2304 1152 104 576 1152 1152
Based on the principle of channel independence (Nie et al., 2022; Han et al., 2024), we treat the variables of each time series as individual data samples. We use an Adam optimizer with a learning rate 0.0001 and a batch size 256 to fine-tune MAE. All experiments are repeated three times. The training epoch is one for all the datasets except Illness, for which we train MAE for 100 epochs with an early stop due to the limited training dataset scale. We conduct tuning on validation sets for the three hyperparameters, $r$, $c$, and $L$. The final hyperparameters used are summarized in Table 16.
C.2 Standard deviations
Table 17: Standard deviations of full-shot experiments.
Method VisionTS Time-LLM GPT4TS
Metric MSE MAE MSE MAE MSE MAE
ETTh1 96 0.347 $\pm$ 0.002 0.376 $\pm$ 0.000 0.376 $\pm$ 0.003 0.402 $\pm$ 0.002 0.370 $\pm$ 0.003 0.389 $\pm$ 0.001
192 0.385 $\pm$ 0.001 0.400 $\pm$ 0.000 0.407 $\pm$ 0.003 0.421 $\pm$ 0.002 0.412 $\pm$ 0.003 0.413 $\pm$ 0.001
336 0.407 $\pm$ 0.001 0.415 $\pm$ 0.001 0.430 $\pm$ 0.004 0.438 $\pm$ 0.001 0.448 $\pm$ 0.003 0.431 $\pm$ 0.001
720 0.439 $\pm$ 0.001 0.443 $\pm$ 0.000 0.457 $\pm$ 0.003 0.468 $\pm$ 0.001 0.441 $\pm$ 0.003 0.449 $\pm$ 0.001
ETTh2 96 0.269 $\pm$ 0.003 0.328 $\pm$ 0.002 0.286 $\pm$ 0.003 0.346 $\pm$ 0.002 0.280 $\pm$ 0.001 0.335 $\pm$ 0.001
192 0.332 $\pm$ 0.001 0.374 $\pm$ 0.001 0.361 $\pm$ 0.003 0.391 $\pm$ 0.002 0.348 $\pm$ 0.002 0.380 $\pm$ 0.001
336 0.351 $\pm$ 0.002 0.395 $\pm$ 0.002 0.390 $\pm$ 0.003 0.414 $\pm$ 0.002 0.380 $\pm$ 0.002 0.405 $\pm$ 0.001
720 0.390 $\pm$ 0.003 0.430 $\pm$ 0.002 0.405 $\pm$ 0.003 0.434 $\pm$ 0.002 0.406 $\pm$ 0.002 0.436 $\pm$ 0.001
ETTm1 96 0.281 $\pm$ 0.001 0.322 $\pm$ 0.001 0.291 $\pm$ 0.001 0.341 $\pm$ 0.001 0.300 $\pm$ 0.001 0.340 $\pm$ 0.000
192 0.322 $\pm$ 0.006 0.353 $\pm$ 0.002 0.341 $\pm$ 0.001 0.369 $\pm$ 0.001 0.343 $\pm$ 0.001 0.368 $\pm$ 0.000
336 0.356 $\pm$ 0.003 0.379 $\pm$ 0.002 0.359 $\pm$ 0.002 0.379 $\pm$ 0.001 0.376 $\pm$ 0.001 0.386 $\pm$ 0.000
720 0.391 $\pm$ 0.001 0.413 $\pm$ 0.001 0.433 $\pm$ 0.001 0.419 $\pm$ 0.001 0.431 $\pm$ 0.001 0.416 $\pm$ 0.000
ETTm2 96 0.169 $\pm$ 0.003 0.256 $\pm$ 0.002 0.162 $\pm$ 0.001 0.248 $\pm$ 0.001 0.163 $\pm$ 0.001 0.249 $\pm$ 0.001
192 0.225 $\pm$ 0.003 0.294 $\pm$ 0.003 0.235 $\pm$ 0.002 0.304 $\pm$ 0.001 0.222 $\pm$ 0.001 0.291 $\pm$ 0.000
336 0.278 $\pm$ 0.002 0.334 $\pm$ 0.001 0.280 $\pm$ 0.002 0.329 $\pm$ 0.001 0.273 $\pm$ 0.001 0.327 $\pm$ 0.001
720 0.372 $\pm$ 0.002 0.392 $\pm$ 0.002 0.366 $\pm$ 0.002 0.382 $\pm$ 0.001 0.357 $\pm$ 0.001 0.376 $\pm$ 0.001
Weather 96 0.142 $\pm$ 0.000 0.192 $\pm$ 0.001 0.155 $\pm$ 0.001 0.199 $\pm$ 0.001 0.148 $\pm$ 0.001 0.188 $\pm$ 0.000
192 0.191 $\pm$ 0.000 0.238 $\pm$ 0.000 0.223 $\pm$ 0.001 0.261 $\pm$ 0.001 0.192 $\pm$ 0.001 0.230 $\pm$ 0.000
336 0.246 $\pm$ 0.003 0.282 $\pm$ 0.001 0.251 $\pm$ 0.001 0.279 $\pm$ 0.001 0.246 $\pm$ 0.001 0.273 $\pm$ 0.000
720 0.328 $\pm$ 0.004 0.337 $\pm$ 0.001 0.345 $\pm$ 0.001 0.342 $\pm$ 0.001 0.320 $\pm$ 0.001 0.328 $\pm$ 0.000
Traffic 96 0.344 $\pm$ 0.001 0.236 $\pm$ 0.000 0.392 $\pm$ 0.001 0.267 $\pm$ 0.000 0.396 $\pm$ 0.001 0.264 $\pm$ 0.000
192 0.372 $\pm$ 0.001 0.249 $\pm$ 0.001 0.409 $\pm$ 0.001 0.271 $\pm$ 0.000 0.412 $\pm$ 0.001 0.268 $\pm$ 0.000
336 0.383 $\pm$ 0.001 0.257 $\pm$ 0.001 0.434 $\pm$ 0.001 0.296 $\pm$ 0.000 0.421 $\pm$ 0.001 0.273 $\pm$ 0.000
720 0.422 $\pm$ 0.001 0.280 $\pm$ 0.000 0.451 $\pm$ 0.001 0.291 $\pm$ 0.000 0.455 $\pm$ 0.001 0.291 $\pm$ 0.000
Electricity 96 0.126 $\pm$ 0.000 0.218 $\pm$ 0.000 0.137 $\pm$ 0.000 0.233 $\pm$ 0.000 0.141 $\pm$ 0.000 0.239 $\pm$ 0.000
192 0.146 $\pm$ 0.001 0.239 $\pm$ 0.001 0.152 $\pm$ 0.000 0.247 $\pm$ 0.000 0.158 $\pm$ 0.000 0.253 $\pm$ 0.000
336 0.161 $\pm$ 0.001 0.255 $\pm$ 0.001 0.169 $\pm$ 0.000 0.267 $\pm$ 0.000 0.172 $\pm$ 0.000 0.266 $\pm$ 0.000
720 0.193 $\pm$ 0.000 0.286 $\pm$ 0.000 0.200 $\pm$ 0.000 0.290 $\pm$ 0.000 0.207 $\pm$ 0.000 0.293 $\pm$ 0.000
$\bf 1^{st}$ count 42 2 12
We report the standard deviations of our full-shot experiments computed on three runs in Table 17, including the results of Time-LLM and GPT4TS from Tan et al. (2024) for reference.
C.3 Ablation study and fine-tuning strategy comparison
Table 18: Full results of Tables 5 and 5: Ablation studies (left) and fine-tuning strategies (right). Results are averaged on four prediction lengths: {96, 192, 336, 720}.
Ablation on Visual MAE (VM) Ablation on trained parameters
- w/o VM VM2Attn VM2Trsf Rand-VM All LN Bias MLP Attn Freeze
ETTh1 MSE 0.395 0.785 0.448 0.459 0.534 ETTh1 MSE 0.534 0.395 0.401 0.534 0.554 0.419
MAE 0.409 0.649 0.458 0.462 0.470 MAE 0.470 0.409 0.414 0.471 0.479 0.418
ETTh2 MSE 0.336 0.420 0.418 0.448 0.411 ETTh2 MSE 0.411 0.336 0.347 0.401 0.392 0.340
MAE 0.382 0.453 0.445 0.457 0.432 MAE 0.432 0.382 0.392 0.419 0.414 0.376
ETTm1 MSE 0.338 0.676 0.397 0.398 0.433 ETTm1 MSE 0.433 0.338 0.343 0.441 0.444 0.374
MAE 0.367 0.562 0.415 0.410 0.413 MAE 0.413 0.367 0.368 0.415 0.415 0.372
ETTm2 MSE 0.261 0.379 0.274 0.292 0.288 ETTm2 MSE 0.288 0.261 0.256 0.292 0.289 0.305
MAE 0.319 0.415 0.334 0.344 0.341 MAE 0.341 0.319 0.318 0.342 0.339 0.334
Average MSE 0.333 0.565 0.384 0.399 0.417 Average MSE 0.417 0.333 0.337 0.417 0.420 0.360
MAE 0.369 0.520 0.413 0.418 0.414 MAE 0.414 0.369 0.373 0.412 0.412 0.375
$\bf 1^{st}$ count 10 0 0 0 0 $\bf 1^{st}$ count 0 7 2 0 0 1
We compare the following ablation variants to verify the role of the visual model (VM), similar to Tan et al. (2024).
  • w/o VM removes all the transformer blocks in encoders and decoders.
  • VM2Attn replaces both the encoder and decoder with a self-attention layer, matching MAE structure but with random initialization.
  • VM2Trsf is similar to VM2Attn but replaces them with a Transformer block (i.e., a self-attention layer plus an MLP layer).
  • Rand-VM keeps the same architecture as the vanilla MAE, but all the weights are randomly initialized.
We also compare fine-tuning different components in MAE as follows:
  • All fine-tunes all the trainable weights in MAE.
  • LN fine-tunes only the layer normalization, which is the default setting used in our experiments.
  • Bias fine-tunes only the bias term of all the linear layers, proposed by Zaken et al. (2022).
  • MLP and Attn fine-tune only the feed-forward layer and the self-attention layer, respectively.
  • Freeze does not fine-tune any weight. Note that it differs from the previous zero-shot experiment, where a longer context length was used (see Table 7 and Table 16).
The results are shown in Table 18, suggesting that visual knowledge is crucial for VisionTS and fine-tuning the layer normalization is the best.
Appendix D Visualization
We visualized the predictions of VisionTS in the zero-shot setting, including its input and reconstructed images. We also visualized the predictions of MoiraiLarge and Seasonal Naïve, wich their MAE metrics for comparison. Figs. 8, 9 and 10 show examples where VisionTS performed well, with Fig. 8 depicting a more regular pattern, while Figs. 9 and 10 display less obvious patterns. Fig. 11 illustrates a case where VisionTS underperformed, as it aggressively predicted the trend despite the lack of clear patterns in the input sequence, whereas MoiraiLarge made more conservative predictions.
(a) Input Image
(b) Reconstructed Image
(c) VisionTS (MAE = 0.312)
(d) MoiraiLarge (MAE = 0.503)
(e) Seasonal Naïve (MAE = 0.774)
Figure 8: Forecasting visualization on a sample from ETTh1. (a-b) Input/output images of VisionTS. (c-e) Forecasting visualization.
(a) Input Image
(b) Reconstructed Image
(c) VisionTS (MAE = 0.157)
(d) MoiraiLarge (MAE = 0.251)
(e) Seasonal Naïve (MAE = 0.235)
Figure 9: Forecasting visualization on a sample from ETTh2. (a-b) Input/output images of VisionTS. (c-e) Forecasting visualization.
(a) Input Image
(b) Reconstructed Image
(c) VisionTS (MAE = 0.821)
(d) MoiraiLarge (MAE = 1.285)
(e) Seasonal Naïve (MAE = 1.523)
Figure 10: Forecasting visualization on a sample from ETTh2. (a-b) Input/output images of VisionTS. (c-e) Forecasting visualization.
(a) Input Image
(b) Reconstructed Image
(c) VisionTS (MAE = 0.327)
(d) MoiraiLarge (MAE = 0.172)
(e) Seasonal Naïve (MAE = 0.364)
Figure 11: Forecasting visualization on a sample from ETTh1, where Moirai outperforms VisionTS in terms of MAE. (a-b) Input/output images of VisionTS. (c-e) Forecasting visualization.

### v3

@@
 Foundation models have emerged as a promising approach in time series forecasting (TSF).
-Existing approaches either fine-tune large language models (LLMs) or build large-scale time-series datasets to develop TSF foundation models.
+Existing approaches either repurpose large language models (LLMs) or build large-scale time series datasets to develop TSF foundation models for universal forecasting.
 However, these methods face challenges due to the severe cross-domain gap or in-domain heterogeneity.
-In this paper, we explore a new road to building a TSF foundation model from rich and high-quality natural images, based on the intrinsic similarities between images and time series.
-To bridge the gap between the two domains, we reformulate the TSF task as an image reconstruction task, which is further processed by a visual masked autoencoder (MAE) self-supervised pre-trained on the ImageNet dataset.
-Surprisingly, without further adaptation in the time-series domain, the proposed VisionTS could achieve superior zero-shot forecasting performance compared to existing TSF foundation models.
-With minimal fine-tuning, VisionTS could further improve the forecasting and achieve state-of-the-art performance in most cases.
-These findings suggest that visual models could be a free lunch for TSF and highlight the potential for future cross-domain research between computer vision and TSF.
-Our code is publicly available at https://github.com/Keytoyze/VisionTS.
-Figure 1: Long-term forecasting performance comparison. Our VisionTS, without any training on time series data, outperforms the largest foundation model MoiraiLarge in the zero-shot setting.
+This paper explores a new road to building a TSF foundation model from rich, high-quality natural images.
+Our key insight is that a visual masked autoencoder, pre-trained on the ImageNet dataset, can naturally be a numeric series forecaster.
+By reformulating TSF as an image reconstruction task, we bridge the gap between image pre-training and TSF downstream tasks.
+Surprisingly, without further adaptation in the time series domain, the proposed VisionTS could achieve better zero-shot forecast performance than existing TSF foundation models.
+With fine-tuning for one epoch, VisionTS could further improve the forecasting and achieve state-of-the-art performance in most cases.
+Extensive experiments reveal intrinsic similarities between images and real-world time series, suggesting that visual models may offer a “free lunch” for TSF and highlight the potential for future cross-modality research.
+Our code is available in the Supplementary Material.
+Time Series Forecasting, Foundation Models, Transfer Learning
+Figure 1: Long-term forecasting performance comparison. Our VisionTS, without any training on time series data, outperforms the pure time series foundation model MoiraiLarge in the zero-shot setting.
 1 Introduction
@@
 A TSF foundation model can greatly reduce the need for downstream data and demonstrate strong forecasting performance on diverse domains, such as energy consumption planning, weather forecasting, and traffic flow.
-We have recently witnessed two roads to building a TSF foundation model.
-The first tries to fine-tune the weights of large language models (LLMs) which have been pre-trained on text data for TSF tasks (i.e., text-based) (Zhou et al., 2023; Jin et al., 2024), based on the observation that LLMs and TSF models share a similar left-to-right forecasting paradigm.
+We have recently witnessed two roads to building a TSF foundation model for universal forecasting.
+The first tries to repurpose large language models (LLMs) that have been pre-trained on text data for TSF tasks (i.e., text-based) (Zhou et al., 2023; Jin et al., 2024), based on the observation that LLMs and TSF models share a similar left-to-right forecasting paradigm.
 However, due to the significant gap between these two modalities, the effectiveness of such transferability between language and time series has recently been questioned by Tan et al.
 (2024).
-The second road focuses on constructing large-scale time-series datasets collected from diverse domains to train a TSF foundation model from scratch (i.e., TS-based) (Woo et al., 2024; Das et al., 2024).
-Nevertheless, unlike images or language with unified formats, time series data is highly heterogeneous in length, frequency, dimensionality, domains, and semantics, limiting the transferability between pre-training and downstream domains.
+The second road focuses on constructing large-scale time-series datasets collected from diverse domains to train a TSF foundation model from scratch (i.e., time series-based or TS-based) (Woo et al., 2024; Das et al., 2024).
+Nevertheless, unlike images or language with unified formats, time series data is highly heterogeneous in length, frequency, number of variates, domains, and semantics, limiting the transferability between pre-training and downstream domains.
 Until recently, constructing a high-quality dataset remains challenging and is still in the early exploration stage.
 In this paper, we investigate a third road that is less explored yet promising: building TSF foundation models with pre-trained visual models.
-Our starting point is the intrinsic similarities between natural images and TS: 1 Similar modalities: Unlike discrete texts, both images and TS are continuous.
-2 Similar origin: Both TS and images are observations of real-world physical systems, whereas languages are products of human cognitive processes.
-3 Similar information density: Languages are human-generated signals with high semantic density, while images and TS data are natural signals with heavy redundancy (He et al., 2022).
-4 Similar features: As shown in Fig.
+Our key idea is that pixel variations in a natural image can be interpreted as temporal sequences, which share many intrinsic similarities with time series: 1 Similar modalities: Unlike discrete texts, both images and time series are continuous; 2 Similar origin: Both time series and images are observations of real-world physical systems, whereas languages are products of human cognitive processes; 3 Similar information density: Languages are human-generated signals with high semantic density, while images and time series are natural signals with heavy redundancy (He et al., 2022); and 4 Similar features: As shown in Fig.
 2, images often display many features of real-world time series, which are rarely found in language data.
-These findings suggest that visual models could be easier to adapt to TSF tasks than language models.
-Based on these insights, we are motivated to answer the question: “Can a visual model pre-trained on images be a free-lunch foundation model for zero-shot time series forecasting?
+Based on these findings, images could be a promising modality for transferring to TSF.
+We are motivated to answer the question: Can a visual model pre-trained on images be a free-lunch foundation model for time series forecasting?
 Figure 2: An image of the ImageNet dataset (Deng et al., 2009), in which the pixel arrays can display many well-known features of real-world time series, such as trend, seasonality, and stationarity (Qiu et al., 2024).
-By self-supervised pre-training on ImageNet, it is reasonable that a visual model could understand these features and exhibit a level of time-series forecasting ability.
+By self-supervised pre-training on ImageNet, it is reasonable that a visual model could understand these features and exhibit a level of time series forecasting ability.
 We focus on visual masked autoencoder (MAE) [^1] We use fonts to distinguish MAE (Masked Autoencoder) and MAE (Mean Absolute Error) in this paper., a popular CV foundation model (He et al., 2022) by self-supervised pre-training on ImageNet (Deng et al., 2009).
+As an image reconstruction and completion model, MAE can naturally be a numeric series forecaster.
 Inspired by the well-known prompt technique in NLP (Schick & Schütze, 2021), we propose a simple method to reformulate TSF as a patch-level image reconstruction task to bridge the gap between pre-training and downstream tasks.
-Specifically, we transform 1D TS data into 2D matrices via segmentation.
+Specifically, we transform 1D time-series data into 2D matrices via segmentation.
 Then, we render the matrices into images and align the forecasting window with masked image patches.
-This method allows us to make zero-shot forecasting without further adaptation.
-We evaluate our proposed VisionTS on multiple TSF benchmarks, including long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021) and Monash (Godahewa et al., 2021).
+This allows us to make a zero-shot forecast without further adaptation.
+We evaluate our proposed VisionTS on large-scale benchmarks, including 8 long-term TSF (Zhou et al., 2021), 29 Monash (Godahewa et al., 2021), and 23 GIFT-Eval (Aksu et al., 2024) datasets, spanning diverse domains, frequencies, and multivariates.
+To the best of our knowledge, the scale of our evaluation benchmark is the largest among existing TSF foundation models.
 As demonstrated in Fig.
-1, without any further adaptation in the time-series domain, a vanilla MAE can surprisingly achieve a comparable performance or even outperform the state-of-the-art (SOTA) zero-shot TSF foundation models, including text-based and TS-based methods.
-By fine-tuning MAE on each downstream dataset for only one epoch, VisionTS can lead to a SOTA performance on most long-term TSF benchmarks.
+1, without further adaptation on time series, a vanilla MAE can surprisingly achieve a comparable performance or even outperform the strong zero-shot TSF foundation models.
+By fine-tuning MAE in each downstream dataset for a single epoch, VisionTS can lead to SOTA performance in most long-term TSF benchmarks.
+To further understand and explain the transferability, we use an MAE encoder to visualize both modalities, showing a level of similarity between time series and natural image representations.
+Additionally, we observe considerable heterogeneity within time-series data across domains, and images can serve as a bridge to connect these isolated time-series representations.
+This could further explain why VisionTS performs better than some cross-domain TSF models.
 Our findings suggest that time series and natural images may be two sides of a coin, and visual models can be a free lunch for time series forecasting.
-We hope our findings can inspire future cross-domain research on CV and TSF.
+We hope our findings inspire future cross-modal research on CV and TSF.
 Our contributions are summarized as follows:
-  • We explore a road to building a TSF foundation model from natural images, which is conceptually different from the existing text-based and TS-based methods.
+  • We explore a road to building a TSF foundation model from natural images, conceptually different from the existing text-based and TS-based pre-training methods.
   • We introduce VisionTS, a novel TSF foundation model based on a visual MAE. To bridge the gap between the two modalities, we reformulate the TSF task into an image reconstruction task.
-  • Comprehensive evaluations of VisionTS demonstrate its significant zero-shot performance, which even surpasses few-shot text-based TSF foundation models and achieves comparable or superior results to zero-shot TS-based models.
+  • Comprehensive evaluations of VisionTS on large-scale benchmarks across multiple domains demonstrate its significant forecasting performance, surpassing few-shot text-based TSF foundation models and achieving comparable or superior results to zero-shot TS-based models.
 2 Preliminaries
@@
 3 Methodology
-As noted in the Introduction, TS and images share intrinsic similarities, suggesting the transfer potential of visual pre-trained models (particularly MAE in this paper) for TSF tasks.
-This section explains how to reformulate TSF tasks into MAE’s pre-training task, i.e., patch-level image reconstruction.
-Our idea is straightforward: map the look-back/forecasting windows to visible/masked patches, respectively.
-This idea is supported by the recent success of prompt tuning (Schick & Schütze, 2021) in NLP, where the predictions for [mask] token in pre-trained language models (e.g., BERT (Devlin et al., 2019)) are directly used for downstream tasks.
-By unifying the forms of the two tasks, we bridge the gap between the two domains, enabling a MAE for zero-shot TSF directly without adapting the pre-trained parameters.
-Notably, this idea is limited to univariate forecasting since multivariates are intractable to be encoded in a single image.
-Fortunately, recent work shows that channel independence — predicting each variable separately for multivariate forecasting — can still be effective (Nie et al., 2022; Han et al., 2024).
-Therefore, we leave the exploration of multivariate interactions for future work.
+As noted in the Introduction, time series and images share intrinsic similarities, suggesting the transfer potential of pre-trained visual models (particularly MAE in this paper) for TSF.
+To reformulate TSF tasks into MAE’s pre-training task, our high-level idea is straightforward: map the look-back/forecasting windows to visible/masked patches, respectively.
+This idea is supported by the prompt tuning (Schick & Schütze, 2021) in NLP, where the predictions for [mask] token in pre-trained language models, e.g., BERT (Devlin et al., 2019), are directly used for downstream tasks.
+By unifying the forms of the two tasks, we bridge the gap between the two modalities without further training.
 However, implementing this idea poses a challenge: the dimension of time-series data (1D) is different from images (2D).
@@
 Segmentation
-[…]
+Given a univariate input $X\in\mathbb{R}^{L}$, the first goal is to transform it into a 2D matrix.
+We propose to segment it into $\lfloor\nicefrac{{L}}{{P}}\rfloor$ subsequences of length $P$, where $P$ is the periodicity.
+Notably, when the time series lacks clear periodicity, we can set $P=1$ directly, which is also effective in our experiments (Section C.5).
+In practice, $P$ can be determined using statistical methods like Fast Fourier Transform (Wu et al., 2023; Chen et al., 2024) or domain knowledge like sampling frequency (Godahewa et al., 2021; Alexandrov et al., 2020).
+In this paper, we select $P$ based on the sampling frequency, elaborated in Section B.2.
+After that, these subsequences are then stacked into a 2D matrix, denoted by ${\bm{I}}_{\text{raw}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor}$.
+This encoding strategy is proven to be efficient by recent work like TimesNet (Wu et al., 2023) and SparseTSF (Lin et al., 2024), as it allows for the simultaneous capture of both variations within the same period (i.e., intra-period) and across periods with the same phase (i.e., inter-period).
+Moreover, it ensures that each element in ${\bm{I}}_{\text{raw}}$ and its neighbors align with the spatial locality property of images (Krizhevsky et al., 2012), where nearby pixels tend to be similar due to the inherent cohesiveness of objects in the real world.
+Therefore, this further narrows the gap between time series and images.
 Normalization
@@
 However, an excessively low $r$ can result in values that are difficult to distinguish.
-We found that a moderate value (0.4) of $r$ performs well across most scenarios (Section B.9).
+We found that a moderate value (0.4) of $r$ performs well across most scenarios (See Section C.8 for more details).
 Let ${\bm{I}}_{\text{norm}}$ denote the normalized matrix, which is computed as follows:
@@
 Rendering
-It is well-known that each image has three channels.
-We simply render ${\bm{I}}_{\text{norm}}$ as a grayscale image ${\bm{I}}_{\text{grey}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor\times 3}$, where all three channels are identical to ${\bm{I}}_{\text{norm}}$.
-This choice is purely result-driven: In our early experiments, we added a convolutional layer with three output channels to convert the grayscale image into a color image and then fine-tuned it to find the optimal color transformation, which, however, did not significantly influence the performance.
+Since each image has three channels, we simply render ${\bm{I}}_{\text{norm}}$ as a grayscale image ${\bm{I}}_{\text{grey}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor\times 3}$, where all three channels are identical to ${\bm{I}}_{\text{norm}}$.
+This choice is purely result-driven: In our early experiments, we added a convolutional layer with three output channels to convert the grayscale image into a color image and then fine-tuned it to find the optimal color transformation, which, however, did not significantly improve the performance.
 Alignment
@@
 We select bilinear interpolation for the resampling process.
-Moreover, we found that reducing the width of the visible portion can further improve performance, likely because the number of input periods ($\lfloor\nicefrac{{L}}{{P}}\rfloor$) is generally smaller than the period length $P$.
-Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$ to reduce the width, i.e.,
-$n=\left\lfloor c\cdot N\cdot\frac{L}{L+H}\right\rfloor.$
+Moreover, we found that reducing the width of the visible portion can further improve performance.
+One possible explanation is that MAE uses a large masked ratio during pre-training, with only 25% of patches visible.
+Reducing the image width may align the masked ratio more closely with pre-training.
+Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$.
+Similar to $r$, we found that setting $c=0.4$ performs well in our experiments (See Section C.8).
+This can be formulated as $n=\left\lfloor c\cdot N\cdot\nicefrac{{L}}{{(L+H)}}\right\rfloor$.
 Reconstruction and Forecasting
@@
 After de-normalizing and flattening, the forecasting window can be extracted.
-4 Experiments
-We use MAE (Base) with 112M parameters as our backbone.
-We conduct experiments on popular TSF benchmarks, including long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021) and Monash benchmark (Godahewa et al., 2021).
-We select representative baselines for comparison, including TS-based foundation models: Moirai (Woo et al., 2024) and TimesFM (Das et al., 2024); Text-based foundation models: TimeLLM (Jin et al., 2024), GPT4TS (Zhou et al., 2023), and LLMTime (Gruver et al., 2023); and other popular TSF baselines covering both Transformer-based, MLP-based and CNN-based architectures.
-Details are elaborated in Appendix A.
-4.1 Zero-Shot Time Series Forecasting
-Setups
-We first evaluate VisionTS’s zero-shot TSF performance without any fine-tuning on time-series modalities.
-To prevent data leakage and assess the out-of-distribution capabilities, we selected six widely-used datasets from the long-term TSF benchmark that are not included in Moirai’s pre-training set for evaluation.
-Since most baselines cannot perform zero-shot forecasting, we report their few-shot results by fine-tuning on the 10% of the individual target datasets.
-We also evaluate the Monash benchmark including 29 test datasets, which is more challenging for VisionTS since they were used in Moirai’s pre-training but not for VisionTS.
-We set the hyperparameters to $r=c=0.4$.
-Following common practice (Nie et al., 2022; Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section B.1.
 Table 1: Zero-shot or few-shot results on the long-term TSF benchmark.
-Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Section B.2.
+Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Table 9 (Section C.2).
 Bold: the best result.
-Comparisons with TimesFM and LLMTime are detailed in Section B.3.
 [Uncaptioned image] Zero-Shot [Uncaptioned image] Few-Shot (10% In-distribution Downstream Dataset)
-Pretrain [Uncaptioned image] Images [Uncaptioned image] Time-series [Uncaptioned image] Text [Uncaptioned image] No Pretrain
+Pretrain [Uncaptioned image] Images [Uncaptioned image] Time series [Uncaptioned image] Text [Uncaptioned image] No Pretrain
 Method VisionTS MoiraiS MoiraiB MoiraiL TimeLLM GPT4TS DLinear PatchTST TimesNet Autoformer Informer
@@
 $\bf 1^{st}$ count 7 0 3 1 2 1 0 0 0 0 0
+Discussion on Multivariate Forecasting
+In addition to the temporal interactions, multivariate time series data sometimes show interactions between variables.
+While pre-trained vision models effectively capture temporal interactions based on the intrinsic similarities between images and time series, they struggle to capture inter-variable interactions due to a limited number of image channels, especially without further training.
+Fortunately, recent work shows that channel independence — forecasting each variable separately — can be effective and is widely used in recent deep forecasting models (Nie et al., 2022; Han et al., 2024; Jin et al., 2024; Zhou et al., 2023; Lin et al., 2024).
+Following these works, we adopt channel independence in our paper while leaving the exploration of capturing inter-variable interactions to future work.
+4 Experiments
+Figure 4: Performance on the GIFT-Eval Leaderboard.
+Figure 5: Aggregated results on the Monash TSF Benchmark, with full results in Table 13 (Section C.5).
+We follow the standard evaluation protocol proposed by Woo et al.
+(2024) to test our VisionTS on 35 widely-used TSF benchmarks, and additionally evaluate it on the GIFT-Eval (Aksu et al., 2024) which is the largest TSF benchmark for zero-shot foundation models.
+We use MAE (Base) as our backbone by default.
+Baseline and benchmark details are elaborated in Section B.1.
+4.1 Zero-Shot Time Series Forecasting
+Setups
+We first evaluate VisionTS’s zero-shot TSF performance without fine-tuning on time-series modalities.
+To prevent data leakage, we selected six widely-used datasets from the long-term TSF benchmark that are not included in Moirai’s pre-training set for evaluation.
+Since most baselines cannot perform zero-shot forecasting, we report their few-shot results by fine-tuning on the 10% of the individual target datasets.
+We also evaluate the Monash benchmark and GIFT-Eval benchmark.
+Notably, the Monash benchmark is more challenging for VisionTS since they were used in Moirai’s pre-training but not for VisionTS.
+We set the hyperparameters to $r=c=0.4$.
+Following common practice (Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section C.1.
+Table 2: Average MSE of different MAE variants, with full results in Table 15 (Section C.6).
+Base Large Huge
+112M 330M 657M
+ETTh1 0.390 0.378 0.391
+ETTh2 0.333 0.340 0.339
+ETTm1 0.374 0.379 0.383
+ETTm2 0.282 0.286 0.284
+Electricity 0.207 0.209 0.202
+Weather 0.269 0.272 0.292
+Avg. 0.309 0.311 0.315
+Table 3: Computational cost in terms of seconds for forecasting a batch of 32 time series data.
+Context Length 1k 1k 2k 3k 4k
+Prediction Length 1k 2k 3k 4k 1k
+PatchTST 0.01 0.01 0.01 0.01 0.01 0.02 0.03 0.04
+DeepAR 0.26 0.32 0.37 0.43 0.26 4.06 6.10 8.17
+GPT4TS 0.01 0.01 0.01 0.02 0.01 0.03 0.04 0.06
+MoiraiBase 0.03 0.04 0.04 0.05 0.03 0.04 0.05 0.06
+TimesFM 0.08 0.14 0.20 0.27 0.07 0.13 0.20 0.25
+LLMTime (8B) > 200 > 200
+VisionTS ($c=0.4$) 0.04 0.03 0.03 0.03 0.04 0.04 0.05 0.05
 Results on Long-Term TSF Benchmark
@@
 It suggests that the transferability from images to time-series is stronger than from text to time-series, and even comparable to the in-domain transferability between time-series.
-We also include a comparison with traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section B.4, where VisionTS still outperforms all of these traditional methods.
-Results on Monash Benchmark
+We also include a comparison with two TSF foundation models, TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023), in Section C.3, as well as traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section C.4.
+Results show that VisionTS still outperforms all of these baselines.
+Results on GIFT-Eval and Monash Benchmarks
 Fig.
-4 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset (e.g., FFNN, WaveNet, and TBATS) and significantly outperforms the other cross-domain baseline (i.e., LLMTime).
+4 shows the comparison of VisionTS with six previously published TSF foundation models on the GIFT-Eval TSF Leaderboard [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval, where VisionTS surprisingly ranked first in terms of normalized MASE.
+Besides, Fig.
+5 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset and significantly outperforms the other cross-domain baseline (i.e., LLMTime).
 It achieves second place among all baselines, just behind Moirai that pre-trained on all the training datasets.
-This highlights VisionTS’s strong zero-shot forecasting ability and the effective cross-modality transferability between image and time series.
-Qualitative Analysis: When does VisionTS perform well, and when does it not?
-In Appendix D, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails).
-When the input exhibits strong regularity (Fig.
-8), VisionTS effectively forecasts both the periodicity (via segmentation) and trends (via MAE’s capabilities).
-In contrast, Moirai, akin to seasonal naïve methods, struggles to capture inter-period trends.
-For less-structured input (Figs.
-9, 10 and 11), Moirai adopts a conservative approach with lower volatility to minimize errors, while VisionTS takes a more aggressive stance.
-This strategy occasionally yields more accurate trend predictions (Figs.
-9 and 10) but may also result in greater MAE (Fig.
-11).
-Figure 4: Aggregated results on the Monash TSF Benchmark, with full results in Section B.5.
-Base Large Huge
-112M 330M 657M
-ETTh1 0.390 0.378 0.391
-ETTh2 0.333 0.340 0.339
-ETTm1 0.374 0.379 0.383
-ETTm2 0.282 0.286 0.284
-Avg. 0.345 0.346 0.349
-Table 2: MSE of different MAE variants, averaged on four prediction lengths.
-Analysis: Impact of MAE’s parameter size
-In Table 2 (full results in Section B.6), we observe that the overall performance of three MAE variants (112M, 330M, and 657M) outperforms MoiraiSmall and MoiraiLarge.
+This promising result highlights VisionTS’s strong zero-shot forecasting ability and effective cross-modality transferability.
+4.2 Further Analysis of VisionTS
+Backbone Analysis
+In Table 3 (full results in Section C.6), we observe that the overall performance of three MAE variants (112M, 330M, and 657M) outperforms MoiraiSmall and MoiraiLarge.
 Particularly, larger models show a slight decrease in performance.
-This may be due to larger models overfitting image-specific features, reducing their transferability.
-A similar phenomenon was reported in Moirai, where larger models were found to degrade downstream task performance.
+This may be due to larger visual models overfitting image-specific features, reducing their transferability.
+A similar phenomenon was reported in Moirai, where larger models were found to degrade performance.
 We leave the exploration of scaling laws in image-based TSF foundation models for the future.
-Analysis: Impact of interpolation and image orientations
-We conduct experiments to validate our choices in the Alignment step, detailed in Section B.7.
+Additionally, to explore the potential with other vision models, we also test LaMa (Suvorov et al., 2022), a visual inpainting model.
+Results in Section C.6 demonstrate that VisionTS with LaMa performs similarly to Moirai in the zero-shot setting.
+This suggests that the performance is driven by the inherent similarity between images and time series, not solely by the MAE model.
+Computational Cost
+We evaluate the computation cost of different baselines on an NVIDIA A800 GPU.
+Results are averaged on 90 runs.
+Table 3 shows the results between various TSF foundation models, showing that VisionTS are comparable to MoiraiBase and GPT4TS and faster than TimesFM, which is an auto-regressive model.
+While computation time increases with context length for all the other Transformer-based baselines, VisionTS remains nearly constant.
+This is because VisionTS encodes input sequences into an image with constant size, ensuring $O(1)$ efficiency.
+In contrast, Transformer-based methods operate at $O(L^{2})$ relative to context length $L$.
+(a) Weather
+(b) ETTm1
+(c) ETTm2
+(d) Electricity
+Figure 6: MSE (Y-axis) performance of different context lengths $L$ (X-axis), averaged on four prediction lengths.
+Figure 7: Modality visualization of the images (ImageNet) and time series (Monash, Weather, Electricity, and ETTm1) based on the MAE encoder.
+Table 4: Aggregated full-shot forecasting performance on eight long-term TSF benchmarks (ETTh1, ETTh2, ETTm1, ETTm2, Illness, Weather, Traffic, and Electricity).
+VisionTS is fine-tuned only a single epoch on each dataset except for Illness.
+Due to the space limit, we report the $1^{\text{st}}$ count for each baseline, with full results in Table 19 (Section D.2).
+Pretrain [Uncaptioned image] Images [Uncaptioned image] Text [Uncaptioned image] No Pretrain
+Method VisionTS Time-LLM GPT4TS Dlinear PatchTST TimesNet FEDformer Autoformer Stationary ETSformer Informer
+$\bf 1^{st}$ count 46 4 12 0 19 0 0 0 0 0 0
+Hyperparameter Analysis
+Section C.8 illustrates the impact of three hyperparameters.
+For context length $L$, as shown in Fig.
+7, performance typically improves with increasing $L$, particularly on high-frequency datasets like Weather (10-minute frequency) and ETTm1/ETTm2 (15-minute frequency).
+This aligns with other TSF foundation models like Moirai.
+As for the normalization constant $r$ and alignment constant $c$, when both of them are around 0.4, performance is generally well across most benchmarks.
+Modality Analysis: Where does the zero-shot forecastability come from?
+We further examine the gap between time series and images to explain the transferability of zero-shot forecasting.
+We sampled 1,000 images from ImageNet-1k and 300 samples from each time series dataset.
+We fed them into the MAE, maintaining a consistent image mask across all data.
+Fig.
+7 visualizes the MAE encoder outputs of these data, which are flattened and reduced to 2-dimension by t-SNE.
+Notably, some time series, such as ETTm1 and Electricity, fall within the ImageNet distribution.
+It suggests a relatively small gap between images and some time series (e.g., Electricity and ETTm1), which could explain the good transferability.
+Additionally, while ImageNet displays a concentrated distribution, time series are generally more scattered.
+For instance, ETTm1 clusters in the upper right, whereas Monash is found in the lower left, with a significant gap.
+This indicates strong heterogeneity within time series data and suggests that images may serve as a bridge to connect isolated time series modality.
+Ablation Study
+We conduct experiments to validate our choices in the Alignment step, detailed in Section C.7.
 First, we test three different interpolation strategies, which shows that Bilinear interpolation performs best.
@@
 Results show that these changes do not significantly affect performance, suggesting that image reconstruction is isotropic and not influenced by certain orientation.
-Computational Cost and Hyperparameter Analysis
-Section B.8 compares the computation cost between various baselines, showing that VisionTS are faster than GPT4TS and three Moirai variants.
-Additionally, Section B.9 illustrates the impact of three hyperparameters: normalization constant $r$, alignment constant $c$, and context length $L$.
-When $r$ and $c$ are around 0.4, performance is generally well across most benchmarks.
-For context length, performance generally improves with increasing $L$, but may also show a decrease in some datasets like ETTh2.
+Qualitative Analysis: When does VisionTS perform well, and when does it not?
+In Appendix E, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails).
+When the input exhibits strong regularity (Fig.
+11), VisionTS effectively forecasts both the periodicity (via segmentation) and trends (via MAE’s capabilities).
+In contrast, Moirai, akin to seasonal naïve methods, struggles to capture inter-period trends.
+For less-structured input (Figs.
+12, 13 and 14), Moirai adopts a conservative approach with lower volatility to minimize errors, while VisionTS takes a more aggressive stance.
+This strategy occasionally yields more accurate trend predictions (Figs.
+12 and 13) but may also result in greater MAE (Fig.
+14).
 4.3 Full-Shot Long-Term Time Series Forecasting
-Table 3: Full-short forecasting performance on the Informer long-term TSF benchmark.
+Setups
+We evaluate the full-shot capability of each baseline trained on individual long-term TSF benchmarks.
+In addition to the six datasets used for zero-shot forecasting, we also include the popular Traffic and Illness datasets.
+As self-attention and feed-forward layers contain rich knowledge that can be transferred to TSF, we choose to fine-tune only the layer normalization (LN) layers while freezing the other parameters, which is also adopted by Zhou et al.
+(2023).
+Training details are elaborated in Section D.1.
+Main Results
+Table 4 summarizes the full-shot results, with the full results and standard deviations detailed in Section D.2.
+It shows that VisionTS outperforms other baselines in most cases (46 out of 80), surpassing the non-pretrained PatchTST and the language-pretrained GPT4TS.
+Remarkably, except for Illness with the least data, VisionTS demands only a single epoch of fine-tuning.
+This suggests that even minimal fine-tuning enables VisionTS to adapt to time series effectively.
+Compared with Table 1, fine-tuning provides limited benefits for ETTh1 and ETTh2 but significantly improves other datasets.
+We attribute this to the smaller data scale of ETTh1 and ETTh2.
+Ablation Study
+Tan et al.
+(2024) proposed several ablation variants for text-based foundation models, including w/o LLM (removing the LLM), LLM2Attn/LLM2Trsf (replacing the LLM with a single self-attention/Transformer layer), and RandLLM (randomly initializing the LLM).
+They found no significant performance differences and concluded that textual knowledge is unnecessary for TSF.
+We conducted similar ablations to assess the role of the vision model (VM), including w/o VM, VM2Attn, VM2Trsf, and RandVM.
+Section D.3 shows that these variants lead to worse performance, indicating that visual knowledge is beneficial for TSF.
+Analysis: Fine-tuning strategies
+As stated before, we fine-tune only the layer normalization (LN).
+We also tested fine-tuning the bias, MLP, or attention layers, in addition to full fine-tuning and freezing.
+All hyperparameters were kept constant.
+Note that freezing differs from the previous zero-shot experiment, where a longer context length was used.
+Section D.3 show that fine-tuning LN is the best.
+Modifying MLP or attention layers results in significant performance drops, suggesting that valuable knowledge resides in these components.
+5 Related Work
+Depending on the pre-training data, TSF foundation models can be categorized into Text-based and TS-based models.
+We first review related works and then introduce recent research on image-based time series analysis.
+Text-based TSF Foundation Models
+Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks. We review this line of work in Appendix A.
+Time Series-Based TSF Foundation Models
+Self-supervised pre-training a TSF model on the same dataset used for downstream TSF tasks is a well-explored topic (Ma et al., 2023; Zhang et al., 2024), such as denoising autoencoders (Zerveas et al., 2021) or contrastive learning (Woo et al., 2022a; Yue et al., 2022).
+They follow a similar paradigm to the masked autoencoder (MAE) in computer vision.
+However, these methods rarely examine the cross-dataset generalization capabilities.
+Recently, research has shifted towards training universal foundation models, by collecting large-scale time series datasets from diverse domains (Goswami et al., 2024; Liu et al., 2024; Das et al., 2024; Dong et al., 2024; Feng et al., 2024) or generating numerous synthetic time series data (Fu et al., 2024; Yang et al., 2024).
+As a representative method, Woo et al.
+(2024) collected 27 billion observations across nine domains and trained TSF foundation models of various scales, achieving strong zero-shot performance.
+However, given the severe heterogeneity, constructing high-quality large datasets poses significant challenges for building these foundation models.
+Image-Based Time-Series Analysis
+Previous research has investigated encoding time series data into images and used convolutional neural networks (CNNs) trained from scratch for classification (Wang & Oates, 2015a, b; Hatami et al., 2018) or forecasting (Li et al., 2020; Sood et al., 2021; Semenoglou et al., 2023).
+Recent researchers explored using pre-trained models for these imaging time series.
+Li et al.
+(2024) used a pre-trained vision transformer (ViT) for classification.
+Wimmer & Rekabsaz (2023) and Zhang et al.
+(2023) employed vision-language multimodal pre-trained models to extract predictive features and generate text descriptions.
+Yang et al.
+(2024) generated synthetic time series data to pre-train a vision model for the TSF task.
+However, these studies did not deeply examine the transferability from natural images to TSF.
+Despite early efforts by Zhou et al.
+(2023) to fine-tune a BEiT (Bao et al., 2022) trained on images for time series forecasting, it still falls short of the leading text-based and TS-based TSF foundation models.
+To the best of our knowledge, we are the first to show that an image-based foundation model, without further time-series adaptation, can match or even surpass other types of TSF foundation models.
+6 Conclusion
+In this paper, we explore a novel approach to building a time series forecasting (TSF) foundation model using natural images, offering a new perspective distinct from the traditional text-based and TS-based methods.
+By leveraging the intrinsic similarities between images and time series, we introduced VisionTS, an MAE-based TSF foundation model that reformulates the TSF task as an image reconstruction problem.
+Our extensive evaluations demonstrate that VisionTS achieves outstanding forecasting performance in zero-shot and full-shot settings, being a free lunch for a TSF foundation model.
+We hope our findings could open new avenues for further cross-modality research.
+Limitations and Future Directions.
+(1) As a preliminary study, we employed MAE and LaMa.
+Utilizing more advanced models like diffusion models (Rombach et al., 2022; Peebles & Xie, 2023) presents a promising research direction.
+(2) Due to limitations in the visual model, VisionTS cannot capture multivariate interactions and perform distribution forecasting.
+Future modifications to the model structure may empower it with more time series capabilities.
+Impact Statement
+This paper presents work whose goal is to advance the field of time series forecasting.
+There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix
+Appendix A Extended Related Work
+Text-based TSF Foundation Models
+Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks.
+For example, Zhou et al.
+(2023) fine-tuned a pre-trained GPT (Radford et al., 2019) on each time-series downstream task, such as forecasting, classification, imputation, and anomaly detection.
+Based on Llama (Touvron et al., 2023), Jin et al.
+(2024) froze the pre-trained LLM and reprogrammed the time series to align with the language modality.
+Bian et al.
+(2024) adopted a two-stage approach by continually pre-training GPT (Radford et al., 2019) on the time-series domain.
+Nevertheless, the TSF performance of LLMs has recently been questioned by Tan et al.
+(2024), which designed several ablation studies to show that textual knowledge is unnecessary for forecasting.
+In this paper, we attribute it to the large modality gap.
+Some recent approaches focus on directly transforming the time series into natural texts for LLMs, allowing for zero-shot forecasting.
+For example, PromptCast (Xue & Salim, 2023) used pre-defined templates to describe numerical time series data, while LLMTime (Gruver et al., 2023) directly separated time steps using commas and separates digits using spaces to construct the text input.
+However, due to the efficiency issue of the autoregressive decoding strategy and the expensive inference cost of large language models, their practical use is limited.
+Appendix B Details of Experiments
+B.1 Benchmark and baselines
+Long-Term TSF Benchmark
+We evaluate our model on 8 widely used long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021), including ETTh1, ETTh2, ETTm1, ETTm2, Electricity, Traffic, Illness, and Weather.
+Performance is assessed using Mean Squared Error (MSE) and Mean Absolute Error (MAE), with lower values indicating better forecasting accuracy.
+Monash Benchmark
+Following Woo et al.
+(2024), we tested 29 Monash datasets (Godahewa et al., 2021) using GluonTS (Alexandrov et al., 2020), including M1 Monthly, M3 Monthly, M3 Other, M4 Monthly, M4 Weekly, M4 Daily, M4 Hourly, Tourism Quarterly, Tourism Monthly, CIF 2016, Australian Electricity Demand, Bitcoin, Pedestrian Counts, Vehicle Trips, KDD Cup, Weather, NN5 Daily, NN5 Weekly, Carparts, FRED-MD, Traffic Hourly, Traffic Weekly, Rideshare, Hospital, COVID Deaths, Temperature Rain, Sunspot, Saugeen River Flow, and US Births.
+Performance is assessed using MAE.
+GIFT-Eval Benchmark
+Aksu et al.
+(2024) introduces the General Time Series Forecasting Model Evaluation, GIFT-Eval, encompasses 23 datasets over 144,000 time series and 177 million data points, spanning seven domains, 10 frequencies, multivariate inputs, and prediction lengths ranging from short to long-term forecasts.
+We use a constant context length 2,000 for VisionTS and we report the point forecast performance using MAPE.
+Baselines
+We select representative baselines for comparison, including TS-based and Text-based foundation models, and other popular TSF baselines covering both Transformer-based, MLP-based and CNN-based architectures.
+The baseline models selected for comparison are briefly described below:
+  1. 1.
+    Moirai (Woo et al., 2024) is a TSF foundation model trained on the Large-scale Open Time Series Archive (LOTSA), with over 27B observations across nine domains.
+It has three variants: small, base, and large.
+  2. 2.
+    TimesFM (Das et al., 2024) is a decoder-style TSF foundation model, using a large time-series corpus comprising both real-world and synthetic datasets.
+  3. 3.
+    Time-LLM (Jin et al., 2024) is a text-based TSF foundation model built on Llama, which reprograms time series data to align with the language modality, keeping the LLM frozen.
+  4. 4.
+    GPT4TS (Zhou et al., 2023) (OneFitsAll) is another text-based model based on GPT, fine-tuned for forecasting tasks.
+  5. 5.
+    LLMTime (Gruver et al., 2023) encodes time series data to a text sequence, supporting zero-shot forecasting.
+  6. 6.
+    DLinear (Zeng et al., 2023) proposes a linear forecasting model, enhanced by seasonal-trend decomposition or normalization.
+  7. 7.
+    PatchTST (Nie et al., 2022) uses Transformer encoders with patching and channel independence techniques for improved predictions.
+  8. 8.
+    TimesNet (Wu et al., 2023) applies convolution kernels along the time dimension, using temporal decomposition and periodical segmentation to capture temporal patterns.
+  9. 9.
+    FEDformer (Zhou et al., 2022) employs a sparse frequency domain representation, using frequency-enhanced blocks for cross-time dependency.
+  10. 10.
+    Autoformer (Wu et al., 2021) uses series decomposition blocks and Auto-Correlation to capture cross-time dependency.
+  11. 11.
+    Stationary (Liu et al., 2022) introduces stationarization and de-stationary attention mechanisms.
+  12. 12.
+    ETSFormer (Woo et al., 2022b) leverages exponential smoothing principles, including exponential smoothing and frequency attention mechanisms.
+  13. 13.
+    Informer (Zhou et al., 2021) proposes ProbSparse self-attention and distillation operations.
+For the long-term TSF benchmark, we include TS-based foundation model results from their original papers, Text-based model results from Tan et al.
+(2024), and other baseline results from Zhou et al.
+(2023).
+For the Monash and PF benchmark, we include results from Woo et al.
+(2024).
+Table 5: Periodicity ($P$) search range for the sampling frequency.
+$x$ denotes the number of sampling frequencies.
+For example, for data with a sampling frequency of 2 minutes (2T), we have $x=2$, and the possible search range of $P$ is $\{\nicefrac{{1440}}{{x}},\nicefrac{{10080}}{{x}},1\}=\{720,5040,1\}$.
+Sampling Frequency Possible Seasonalities Possible P
+Second (S) 1 hour $\{\nicefrac{{3600}}{{x}},1\}$
+Minute (T) 1 day or 1 week $\{\nicefrac{{1440}}{{x}},\nicefrac{{10080}}{{x}},1\}$
+Hour (H) 1 day or 1 week $\{\nicefrac{{24}}{{x}},\nicefrac{{168}}{{x}},1\}$
+Day (D) 1 week, 1 month, or 1 year $\{\nicefrac{{7}}{{x}},\nicefrac{{30}}{{x}},\nicefrac{{365}}{{x}},1\}$
+Week (W) 1 year or 1 month $\{\nicefrac{{52}}{{x}},\nicefrac{{4}}{{x}},1\}$
+Month (M) 1 year, 6 months, or 3 months $\{\nicefrac{{12}}{{x}},\nicefrac{{6}}{{x}},\nicefrac{{3}}{{x}},1\}$
+Business Day (B) 1 week $\{\nicefrac{{5}}{{x}},1\}$
+Quarter (Q) 1 year or 6 months $\{\nicefrac{{4}}{{x}},\nicefrac{{2}}{{x}},1\}$
+Others - $\{1\}$
+Table 6: Final $P$ used for each dataset in our experiment.
+Frequency $P$ Datasets
+Long-Term TSF H 24 ETTh1 ETTh2 Electricity Traffic
+W 52 Illness
+15T 96 ETTm1 ETTm2
+10T 144 Weather
+Monash D 1 M4 Daily COVID Deaths
+W 1 NN5 Weekly
+M 1 FRED-MD
+Q 1 M3 Other
+M 3 M3 Monthly M4 Monthly CIF 2016 (6)
+W 4 M4 Weekly Traffic Weekly
+Q 4 Tourism Quarterly
+M 6 CIF 2016 (12) Car Parts
+D 7 Bitcoin Vehicle Trips Weather NN5 Daily
+D 7 US Births Saugeen Day Temperature Rain
+M 12 Tourism Monthly Hospital M1 Monthly
+H 24 M4 Hourly KDD cup Pedestrian Counts
+H 24 Traffic Hourly Rideshare
+D 30 Sunspot
+0.5H 336 Aus. Elec. Demand
+Table 7: Comparison of setting $P=1$ for VisionTS.
+VisionTS $P=1$
+MSE MAE MSE MAE
+ETTh1 0.390 0.414 0.840 0.628
+ETTh2 0.333 0.375 0.424 0.445
+ETTm1 0.374 0.372 0.660 0.533
+ETTm2 0.282 0.321 0.312 0.363
+Average 0.344 0.370 0.559 0.492
+Environment
+All experiments are conducted using Time-Series-Library (https://github.com/thuml/Time-Series-Library) and GluonTS library (Alexandrov et al., 2020) on an NVIDIA A800 GPU.
+B.2 Periodicity selection
+We first determine a range of period lengths based on the sampling frequency of the data, shown in Table 5.
+This frequency-based strategy is also employed by Alexandrov et al.
+(2020) while we extend the search range for tuning.
+We select the optimal $P$ from this range on the validation set.
+The final $P$ used in our experiments are summarized in Table 6.
+To demonstrate the influence of $P$ and the effectiveness of our periodicity selection strategy, we set $P=1$ and compare the results with the above strategy.
+Table 7 shows that such strategy (denoted as VisionTS) significantly outperforms the naive strategy that sets $P=1$.
+Appendix C Zero-Shot Forecasting
+C.1 Hyperparameters
+Table 8: Hyperparameters for VisionTS used in our zero-shot forecasting (Long-term TSF).
+ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
+Normalization constant $r$ 0.4 0.4 0.4 0.4 0.4 0.4
+Alignment constant $c$ 0.4 0.4 0.4 0.4 0.4 0.4
+Context length $L$ 2880 1728 2304 4032 4032 2880
+We conduct hyperparameter tuning on validation sets to determine the optimal context length $L$. Final used hyperparameters are summarized in Table 8.
+C.2 Full forecasting results of the long-term TSF benchmark
+Table 9: Full results of Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Bold: the best result.
+[Uncaptioned image] Zero-Shot [Uncaptioned image] Few-Shot (10% Downstream Dataset)
+Pretrain [Uncaptioned image] Images [Uncaptioned image] Time-series [Uncaptioned image] Text [Uncaptioned image] No Pretrain
+Method VisionTime MoiraiS MoiraiB MoiraiL TimeLLM GPT4TS DLinear PatchTST TimesNet Autoformer Informer
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+$ETTh1$ 96 0.353 0.383 0.375 0.402 0.384 0.402 0.380 0.398 0.448 0.460 0.458 0.456 0.492 0.495 0.516 0.485 0.861 0.628 0.613 0.552 1.179 0.792
+192 0.392 0.410 0.399 0.419 0.425 0.429 0.440 0.434 0.484 0.483 0.570 0.516 0.565 0.538 0.598 0.524 0.797 0.593 0.722 0.598 1.199 0.806
+336 0.407 0.423 0.412 0.429 0.456 0.450 0.514 0.474 0.589 0.540 0.608 0.535 0.721 0.622 0.657 0.550 0.941 0.648 0.750 0.619 1.202 0.811
+720 0.406 0.441 0.413 0.444 0.470 0.473 0.705 0.568 0.700 0.604 0.725 0.591 0.986 0.743 0.762 0.610 0.877 0.641 0.721 0.616 1.217 0.825
+avg 0.390 0.414 0.400 0.424 0.434 0.439 0.510 0.469 0.556 0.522 0.590 0.525 0.691 0.600 0.633 0.542 0.869 0.628 0.702 0.596 1.199 0.809
+$ETTh2$ 96 0.271 0.328 0.281 0.334 0.277 0.327 0.287 0.325 0.275 0.326 0.331 0.374 0.357 0.411 0.353 0.389 0.378 0.409 0.413 0.451 3.837 1.508
+192 0.328 0.367 0.340 0.373 0.340 0.374 0.347 0.367 0.374 0.373 0.402 0.411 0.569 0.519 0.403 0.414 0.490 0.467 0.474 0.477 3.856 1.513
+336 0.345 0.381 0.362 0.393 0.371 0.401 0.377 0.393 0.406 0.429 0.406 0.433 0.671 0.572 0.426 0.441 0.537 0.494 0.547 0.543 3.952 1.526
+720 0.388 0.422 0.380 0.416 0.394 0.426 0.404 0.421 0.427 0.449 0.449 0.464 0.824 0.648 0.477 0.480 0.510 0.491 0.516 0.523 3.842 1.503
+avg 0.333 0.375 0.341 0.379 0.346 0.382 0.354 0.377 0.370 0.394 0.397 0.421 0.605 0.538 0.415 0.431 0.479 0.465 0.488 0.499 3.872 1.513
+$ETTm1$ 96 0.341 0.347 0.404 0.383 0.335 0.360 0.353 0.363 0.346 0.388 0.390 0.404 0.352 0.392 0.410 0.419 0.583 0.501 0.774 0.614 1.162 0.785
+192 0.360 0.360 0.435 0.402 0.366 0.379 0.376 0.380 0.373 0.416 0.429 0.423 0.382 0.412 0.437 0.434 0.630 0.528 0.754 0.592 1.172 0.793
+336 0.377 0.374 0.462 0.416 0.391 0.394 0.399 0.395 0.413 0.426 0.469 0.439 0.419 0.434 0.476 0.454 0.725 0.568 0.869 0.677 1.227 0.908
+720 0.416 0.405 0.490 0.437 0.434 0.419 0.432 0.417 0.485 0.476 0.569 0.498 0.490 0.477 0.681 0.556 0.769 0.549 0.810 0.630 1.207 0.797
+avg 0.374 0.372 0.448 0.410 0.382 0.388 0.390 0.389 0.404 0.427 0.464 0.441 0.411 0.429 0.501 0.466 0.677 0.537 0.802 0.628 1.192 0.821
+$ETTm2$ 96 0.228 0.282 0.205 0.282 0.195 0.269 0.189 0.260 0.177 0.261 0.188 0.269 0.213 0.303 0.191 0.274 0.212 0.285 0.352 0.454 3.203 1.407
+192 0.262 0.305 0.261 0.318 0.247 0.303 0.247 0.300 0.241 0.314 0.251 0.309 0.278 0.345 0.252 0.317 0.270 0.323 0.694 0.691 3.112 1.387
+336 0.293 0.328 0.319 0.355 0.291 0.333 0.295 0.334 0.274 0.327 0.307 0.346 0.338 0.385 0.306 0.353 0.323 0.353 2.408 1.407 3.255 1.421
+720 0.343 0.370 0.415 0.410 0.355 0.377 0.372 0.386 0.417 0.390 0.426 0.417 0.436 0.440 0.433 0.427 0.474 0.449 1.913 1.166 3.909 1.543
+avg 0.282 0.321 0.300 0.341 0.272 0.321 0.276 0.320 0.277 0.323 0.293 0.335 0.316 0.368 0.296 0.343 0.320 0.353 1.342 0.930 3.370 1.440
+$Electricity$ 96 0.177 0.266 0.205 0.299 0.158 0.248 0.152 0.242 0.139 0.241 0.139 0.237 0.150 0.253 0.140 0.238 0.299 0.373 0.261 0.348 1.259 0.919
+192 0.188 0.277 0.220 0.310 0.174 0.263 0.171 0.259 0.151 0.248 0.156 0.252 0.164 0.264 0.160 0.255 0.305 0.379 0.338 0.406 1.160 0.873
+336 0.207 0.296 0.236 0.323 0.191 0.278 0.192 0.278 0.169 0.270 0.175 0.270 0.181 0.282 0.180 0.276 0.319 0.391 0.410 0.474 1.157 0.872
+720 0.256 0.337 0.270 0.347 0.229 0.307 0.236 0.313 0.240 0.322 0.233 0.317 0.223 0.321 0.241 0.323 0.369 0.426 0.715 0.685 1.203 0.898
+avg 0.207 0.294 0.233 0.320 0.188 0.274 0.188 0.273 0.175 0.270 0.176 0.269 0.180 0.280 0.180 0.273 0.323 0.392 0.431 0.478 1.195 0.891
+$Weather$ 96 0.220 0.257 0.173 0.212 0.167 0.203 0.177 0.208 0.161 0.210 0.163 0.215 0.171 0.224 0.165 0.215 0.184 0.230 0.221 0.297 0.374 0.401
+192 0.244 0.275 0.216 0.250 0.209 0.241 0.219 0.249 0.204 0.248 0.210 0.254 0.215 0.263 0.210 0.257 0.245 0.283 0.270 0.322 0.552 0.478
+336 0.280 0.299 0.260 0.282 0.256 0.276 0.277 0.292 0.261 0.302 0.256 0.292 0.258 0.299 0.259 0.297 0.305 0.321 0.320 0.351 0.724 0.541
+720 0.330 0.337 0.320 0.322 0.321 0.323 0.365 0.350 0.309 0.332 0.321 0.339 0.320 0.346 0.332 0.346 0.381 0.371 0.390 0.396 0.739 0.558
+avg 0.269 0.292 0.242 0.267 0.238 0.261 0.260 0.275 0.234 0.273 0.238 0.275 0.241 0.283 0.242 0.279 0.279 0.301 0.300 0.342 0.597 0.495
+Average 0.309 0.345 0.327 0.357 0.310 0.344 0.329 0.350 0.336 0.368 0.360 0.378 0.407 0.416 0.378 0.389 0.491 0.446 0.678 0.579 1.904 0.995
+$\bf 1^{st}$ count 32 0 10 8 10 6 0 0 0 0 0
+Table 9 shows the full results of zero-shot/few-shot long-term forecasting performance.
+VisionTS achieves the best results in most cases (32 out of 62), outperforming MoiraiBase (10 out of 62) and MoiraiLarge (8 out of 62).
+C.3 Comparison of TimesFM and LLMTime
+Table 10: MAE results of TimesFM and LLMTime for zero-shot forecasting, on the last test window of the original test split.
+Method VisionTS TimesFM LLMTime
+ETTh1 96 0.35 0.45 0.42
+192 0.45 0.53 0.50
+ETTh2 96 0.24 0.35 0.33
+192 0.60 0.62 0.70
+ETTm1 96 0.12 0.19 0.37
+192 0.23 0.26 0.71
+ETTm2 96 0.19 0.24 0.29
+192 0.24 0.27 0.31
+Average 0.30 0.36 0.45
+Table 11: Comparison of traditional forecasting baselines in the zero-shot setting.
+Method VisionTS ETS ARIMA Seasonal Naïve Seasonal Avg
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+$ETTh1$ 96 0.353 0.383 1.289 0.710 0.900 0.719 0.512 0.433 0.589 0.585
+192 0.392 0.410 1.319 0.730 0.906 0.724 0.581 0.469 0.598 0.590
+336 0.407 0.423 1.324 0.742 0.908 0.731 0.650 0.501 0.610 0.597
+720 0.406 0.441 1.329 0.751 0.932 0.753 0.655 0.514 0.656 0.624
+avg 0.390 0.414 1.315 0.733 0.912 0.732 0.600 0.479 0.613 0.599
+$ETTh2$ 96 0.271 0.328 0.399 0.408 0.488 0.508 0.391 0.380 0.457 0.494
+192 0.328 0.367 0.500 0.459 0.497 0.514 0.482 0.429 0.466 0.500
+336 0.345 0.381 0.562 0.498 0.507 0.522 0.532 0.466 0.476 0.509
+720 0.388 0.422 0.558 0.506 0.572 0.557 0.525 0.474 0.542 0.548
+avg 0.333 0.375 0.505 0.468 0.516 0.525 0.483 0.437 0.485 0.513
+$ETTm1$ 96 0.341 0.347 1.204 0.659 0.702 0.568 0.423 0.387 0.369 0.399
+192 0.360 0.360 1.251 0.685 0.704 0.570 0.463 0.406 0.374 0.402
+336 0.377 0.374 1.276 0.702 0.709 0.574 0.496 0.426 0.382 0.407
+720 0.416 0.405 1.311 0.724 0.713 0.580 0.574 0.464 0.394 0.416
+avg 0.374 0.372 1.261 0.693 0.707 0.573 0.489 0.421 0.380 0.406
+$ETTm2$ 96 0.228 0.282 0.257 0.324 0.397 0.434 0.263 0.301 0.365 0.411
+192 0.262 0.305 0.331 0.366 0.402 0.436 0.321 0.337 0.369 0.414
+336 0.293 0.328 0.402 0.406 0.407 0.439 0.376 0.370 0.375 0.418
+720 0.343 0.370 0.512 0.462 0.413 0.443 0.471 0.422 0.380 0.423
+avg 0.282 0.321 0.376 0.390 0.405 0.438 0.358 0.357 0.372 0.417
+Average 0.344 0.370 0.864 0.571 0.635 0.567 0.482 0.424 0.463 0.484
+$\bf 1^{st}$ count 41 0 0 0 1
+Table 12: Comparison of traditional zero-shot forecasting baselines.
+Method VisionTS ETS ARIMA Seasonal Naïve Seasonal Avg
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+$ETTh1$ 96 0.353 0.383 1.289 0.710 0.900 0.719 0.512 0.433 0.589 0.585
+192 0.392 0.410 1.319 0.730 0.906 0.724 0.581 0.469 0.598 0.590
+336 0.407 0.423 1.324 0.742 0.908 0.731 0.650 0.501 0.610 0.597
+720 0.406 0.441 1.329 0.751 0.932 0.753 0.655 0.514 0.656 0.624
+avg 0.390 0.414 1.315 0.733 0.912 0.732 0.600 0.479 0.613 0.599
+$ETTh2$ 96 0.271 0.328 0.399 0.408 0.488 0.508 0.391 0.380 0.457 0.494
+192 0.328 0.367 0.500 0.459 0.497 0.514 0.482 0.429 0.466 0.500
+336 0.345 0.381 0.562 0.498 0.507 0.522 0.532 0.466 0.476 0.509
+720 0.388 0.422 0.558 0.506 0.572 0.557 0.525 0.474 0.542 0.548
+avg 0.333 0.375 0.505 0.468 0.516 0.525 0.483 0.437 0.485 0.513
+$ETTm1$ 96 0.341 0.347 1.204 0.659 0.702 0.568 0.423 0.387 0.369 0.399
+192 0.360 0.360 1.251 0.685 0.704 0.570 0.463 0.406 0.374 0.402
+336 0.377 0.374 1.276 0.702 0.709 0.574 0.496 0.426 0.382 0.407
+720 0.416 0.405 1.311 0.724 0.713 0.580 0.574 0.464 0.394 0.416
+avg 0.374 0.372 1.261 0.693 0.707 0.573 0.489 0.421 0.380 0.406
+$ETTm2$ 96 0.228 0.282 0.257 0.324 0.397 0.434 0.263 0.301 0.365 0.411
+192 0.262 0.305 0.331 0.366 0.402 0.436 0.321 0.337 0.369 0.414
+336 0.293 0.328 0.402 0.406 0.407 0.439 0.376 0.370 0.375 0.418
+720 0.343 0.370 0.512 0.462 0.413 0.443 0.471 0.422 0.380 0.423
+avg 0.282 0.321 0.376 0.390 0.405 0.438 0.358 0.357 0.372 0.417
+Average 0.344 0.370 0.864 0.571 0.635 0.567 0.482 0.424 0.463 0.484
+$\bf 1^{st}$ count 41 0 0 0 1
+Due to the step-by-step output of the decoder architecture, the efficiency of TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023) are relatively slower.
+Thus, Das et al.
+(2024) only reported results for the last test window of the original split.
+We compared VisionTS with their results under the same setting, as shown in Table 11.
+VisionTS outperforms TimesFM and LLMTime in terms of MAE, indicating that image-based TSF models are on par with or even better than TS-based and text-based models.
+C.4 Comparison of traditional methods
+In addition to deep learning models, we also compare traditional methods, including ARIMA, ETS, and two methods that require periodicity as our VisionTS: Seasonal Naïve (repeating the last period) and Seasonal Avg (similar to Seasonal Naïve but repeating the average of all periods in the look-back window).
+Due to the high computational cost of ARIMA and ETS, we only compare them on the small-scale benchmarks, i.e., four ETT datasets.
+Table 12 shows that VisionTS also achieves the best performance.
+C.5 Full forecasting results of the Monash TSF benchmark
+Table 13: Full results of Fig. 5: Forecasting results (MAE) on the Monash TSF benchmark. We reported the reproduction results of LLMTime based on the GPT3.5 API from Woo et al. (2024).
+VisionTS LLMTime MoiraiSmall Naive SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer
+M1 Monthly 1987.69 2562.84 2082.26 2707.75 2259.04 2166.18 2237.5 1905.28 2080.13 2088.25 2052.32 2162.58 1860.81 1820.37 2184.42 2723.88
+M3 Monthly 737.93 877.97 713.41 837.14 743.41 623.71 630.59 626.46 654.8 692.97 732 692.48 728.81 648.6 699.3 798.38
+M3 Other 315.85 300.3 263.54 278.43 277.83 215.35 189.42 194.98 193.02 234.43 318.13 240.17 247.56 221.85 245.29 239.24
+M4 Monthly 666.54 728.27 597.6 671.27 625.24 563.58 589.52 582.6 575.36 596.19 611.69 612.52 615.22 578.48 655.51 780.47
+M4 Weekly 404.23 518.44 339.76 347.99 336.82 333.32 296.15 335.66 321.61 293.21 364.65 338.37 351.78 277.73 359.46 378.89
+M4 Daily 215.63 266.52 189.1 180.83 178.27 178.86 176.6 193.26 179.67 181.92 231.36 177.91 299.79 190.44 189.47 201.08
+M4 Hourly 288.37 576.06 268.04 1218.06 1218.06 1220.97 386.27 3358.1 1310.85 257.39 285.35 385.49 886.02 425.75 393.63 320.54
+Tourism Quarterly 12931.88 16918.86 18352.44 15845.1 15014.19 7656.49 9972.42 8925.52 10475.47 9092.58 10267.97 8981.04 9511.37 8640.56 9137.12 9521.67
+Tourism Monthly 2560.19 5608.61 3569.85 5636.83 5302.1 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98
+CIF 2016 570907.24 599313.8 655888.58 578596.5 581875.97 714818.6 855578.4 642421.4 469059 563205.57 603551.3 1495923 3200418 679034.8 5998225 4057973
+Aus. Elec. Demand 237.44 760.81 266.57 659.6 659.6 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.5 231.45
+Bitcoin 2.33E+18 1.74E+18 1.76E+18 7.78E+17 5.33E+18 5.33E+18 9.9E+17 1.1E+18 3.62E+18 6.66E+17 1.93E+18 1.45E+18 1.95E+18 1.06E+18 2.46E+18 2.61E+18
+Pedestrian Counts 52.01 97.77 54.88 170.88 170.87 170.94 222.38 216.5 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29
+Vehicle Trips 22.08 31.48 24.46 31.42 29.98 30.76 21.21 30.95 30.07 27.24 22.61 22.93 22 28.16 24.15 28.01
+KDD cup 38.16 42.72 39.81 42.13 42.04 42.06 39.2 44.88 52.2 36.85 34.82 37.16 48.98 49.1 37.08 44.46
+Weather 2.06 2.17 1.96 2.36 2.24 2.51 2.3 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03
+NN5 Daily 3.51 7.1 5.37 8.26 6.63 3.8 3.7 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16
+NN5 Weekly 14.67 15.76 15.07 16.71 15.66 15.3 14.98 15.7 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34
+Carparts 0.58 0.44 0.53 0.65 0.55 0.53 0.58 0.56 0.56 0.41 0.53 0.39 0.39 0.98 0.4 0.39
+FRED-MD 1893.67 2804.64 2568.48 2825.67 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.8 2508.4 4666.04
+Traffic Hourly 0.01 0.03 0.02 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01
+Traffic Weekly 1.14 1.15 1.17 1.19 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.2 1.42
+Rideshare 5.92 6.28 1.35 6.29 6.29 7.62 6.45 6.29 3.37 6.3 6.07 6.59 6.28 5.55 2.75 6.29
+Hospital 19.36 25.68 23 24.07 21.76 18.54 17.43 17.97 19.6 19.24 19.17 22.86 18.25 20.18 19.35 36.19
+COVID Deaths 137.51 653.31 124.32 353.71 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66
+Temperature Rain 6.37 6.37 5.3 9.39 8.18 8.22 7.14 8.21 7.19 6.13 6.76 5.56 5.37 7.28 5.81 5.24
+Sunspot 2.81 5.07 0.11 3.93 4.93 4.93 2.57 4.93 2.57 3.83 2.27 7.97 0.77 14.47 0.17 0.13
+Saugeen River Flow 30.22 34.84 24.07 21.5 21.5 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06
+US Births 519.94 1374.99 872.51 1152.67 1192.2 586.93 399 419.73 526.33 574.93 441.7 557.87 424.93 422 504.4 452.87
+Normalized MAE 0.729 1.041 0.657 1.000 1.028 0.927 0.758 0.872 0.898 0.785 0.760 0.741 0.759 0.783 0.749 0.770
+Rank 2 16 1 14 15 13 5 11 12 10 7 3 6 9 4 8
+Setup
+Table 6 lists the sampling frequency and the selected period $P$ for each dataset.
+Datasets with $P=1$ indicate no significant periodicity, where we use a context length of $L=300$.
+For other datasets with $P>1$, we select a longer context length of $L=1000$.
+All datasets were tested with the hyperparameters $r=c=0.4$ as we had done for the long-term TSF benchmark.
+Results
+Table 13 presents VisionTS ’s MAE test results, with the normalized MAE calculated by dividing each dataset’s MAE by the naive forecast’s MAE and aggregated using the geometric mean across datasets.
+We include the result of each baseline from Woo et al.
+(2024).
+Particularly, we find that VisionTS outperforms Moirai on some datasets with $P=1$ (e.g., FRED-MD and NN5 Weekly), showing that VisionTS can still work effectively without significant periodicity.
+Table 14: Comparison of LaMa as the backbone. Results are averaged on four prediction lengths.
+MAE LaMa MoiraiSmall MoiraiLarge
+MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1 0.390 0.414 0.425 0.433 0.400 0.424 0.510 0.469
+ETTh2 0.333 0.375 0.376 0.408 0.341 0.379 0.354 0.377
+ETTm1 0.374 0.372 0.400 0.391 0.448 0.410 0.390 0.389
+ETTm2 0.282 0.321 0.294 0.337 0.300 0.341 0.276 0.320
+Average 0.344 0.370 0.374 0.392 0.372 0.388 0.382 0.388
+C.6 Impact of backbones
+Table 15: Full results of Table 3: zero-shot forecasting results of different MAE variants. Bold: best results among three variants. We also include the results from Moirai for reference.
+Method MAE (Base) MAE (Large) MAE (Huge) Moirai (Small) Moirai (Base) Moirai (Huge)
+112M 330M 657M 14M 91M 311M
+Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+$ETTh1$ 96 0.353 0.383 0.346 0.382 0.362 0.384 0.375 0.402 0.384 0.402 0.380 0.398
+192 0.392 0.410 0.379 0.406 0.407 0.414 0.399 0.419 0.425 0.429 0.440 0.434
+336 0.407 0.423 0.391 0.416 0.399 0.419 0.412 0.429 0.456 0.450 0.514 0.474
+720 0.406 0.441 0.397 0.433 0.395 0.433 0.413 0.444 0.470 0.473 0.705 0.568
+avg 0.390 0.414 0.378 0.409 0.391 0.412 0.400 0.424 0.434 0.439 0.510 0.469
+$ETTh2$ 96 0.271 0.328 0.286 0.334 0.285 0.333 0.281 0.334 0.277 0.327 0.287 0.325
+192 0.328 0.367 0.346 0.375 0.337 0.369 0.340 0.373 0.340 0.374 0.347 0.367
+336 0.345 0.381 0.356 0.387 0.357 0.388 0.362 0.393 0.371 0.401 0.377 0.393
+720 0.388 0.422 0.371 0.409 0.379 0.412 0.380 0.416 0.394 0.426 0.404 0.421
+avg 0.333 0.375 0.340 0.377 0.339 0.375 0.341 0.379 0.346 0.382 0.354 0.377
+$ETTm1$ 96 0.341 0.347 0.344 0.349 0.352 0.351 0.404 0.383 0.335 0.360 0.353 0.363
+192 0.360 0.360 0.365 0.363 0.360 0.367 0.435 0.402 0.366 0.379 0.376 0.380
+336 0.377 0.374 0.381 0.376 0.381 0.383 0.462 0.416 0.391 0.394 0.399 0.395
+720 0.416 0.405 0.429 0.411 0.440 0.412 0.490 0.437 0.434 0.419 0.432 0.417
+avg 0.374 0.372 0.379 0.375 0.383 0.378 0.448 0.410 0.382 0.388 0.390 0.389
+$ETTm2$ 96 0.228 0.282 0.225 0.282 0.229 0.282 0.205 0.282 0.195 0.269 0.189 0.260
+192 0.262 0.305 0.262 0.305 0.265 0.306 0.261 0.318 0.247 0.303 0.247 0.300
+336 0.293 0.328 0.299 0.331 0.286 0.324 0.319 0.355 0.291 0.333 0.295 0.334
+720 0.343 0.370 0.358 0.377 0.355 0.374 0.415 0.410 0.355 0.377 0.372 0.386
+avg 0.282 0.321 0.286 0.324 0.284 0.322 0.300 0.341 0.272 0.321 0.276 0.320
+$Electricity$ 96 0.177 0.266 0.177 0.268 0.170 0.259 0.205 0.299 0.158 0.248 0.152 0.242
+192 0.188 0.277 0.192 0.283 0.182 0.273 0.220 0.310 0.174 0.263 0.171 0.259
+336 0.207 0.296 0.213 0.303 0.207 0.295 0.236 0.323 0.191 0.278 0.192 0.278
+720 0.256 0.337 0.256 0.337 0.250 0.333 0.270 0.347 0.229 0.307 0.236 0.313
+avg 0.207 0.294 0.209 0.298 0.202 0.290 0.233 0.320 0.188 0.274 0.188 0.273
+$Weather$ 96 0.220 0.257 0.222 0.257 0.235 0.265 0.173 0.212 0.167 0.203 0.177 0.208
+192 0.244 0.275 0.246 0.275 0.276 0.288 0.216 0.250 0.209 0.241 0.219 0.249
+336 0.280 0.299 0.283 0.301 0.304 0.309 0.260 0.282 0.256 0.276 0.277 0.292
+720 0.330 0.337 0.338 0.343 0.351 0.350 0.320 0.322 0.321 0.323 0.365 0.350
+avg 0.269 0.292 0.272 0.294 0.292 0.303 0.242 0.267 0.238 0.261 0.260 0.275
+Average 0.309 0.345 0.311 0.346 0.315 0.347 0.327 0.357 0.310 0.344 0.329 0.350
+$\bf 1^{st}$ count 38 17 17 - - -
+Table 15 compares zero-shot forecasting performance of three MAE variants (112M, 330M, and 657M), showing that the three variants are similar, but larger models show a slight decrease.
+Particularly, the smallest model excels in ETTh2, ETTm1, ETTm2, and Weather, while the largest model excels in Electricity.
+Additionally, Table 14 compares VisionTS with another visual backbone, LaMa.
+C.7 Impact of the different image encoding strategies
+Table 16: Impact of resampling filters and image orientations.
+Interpolation strategies in resampling Image orientation
+Method Bilinear Bicubic Nearest Neighbor Method - Horizontal flip Vertical flip
+Metric MSE MAE MSE MAE MSE MAE Metric MSE MAE MSE MAE MSE MAE
+$ETTh1$ 96 0.353 0.383 0.351 0.383 0.426 0.424 $ETTh1$ 96 0.353 0.383 0.348 0.379 0.355 0.385
+192 0.392 0.410 0.392 0.409 0.450 0.443 192 0.392 0.410 0.386 0.404 0.394 0.411
+336 0.407 0.423 0.407 0.422 0.451 0.450 336 0.407 0.423 0.401 0.416 0.408 0.423
+720 0.406 0.441 0.405 0.440 0.454 0.470 720 0.406 0.441 0.399 0.430 0.406 0.442
+avg 0.390 0.414 0.389 0.414 0.445 0.446 avg 0.390 0.414 0.384 0.407 0.391 0.415
+$ETTh2$ 96 0.271 0.328 0.274 0.329 0.298 0.349 $ETTh2$ 96 0.271 0.328 0.274 0.329 0.274 0.330
+192 0.328 0.367 0.330 0.367 0.343 0.380 192 0.328 0.367 0.331 0.370 0.330 0.367
+336 0.345 0.381 0.345 0.380 0.373 0.401 336 0.345 0.381 0.347 0.386 0.345 0.381
+720 0.388 0.422 0.386 0.419 0.404 0.431 720 0.388 0.422 0.376 0.416 0.388 0.422
+avg 0.333 0.375 0.334 0.374 0.354 0.390 avg 0.333 0.375 0.332 0.375 0.334 0.375
+$ETTm1$ 96 0.341 0.347 0.366 0.354 0.399 0.374 $ETTm1$ 96 0.341 0.347 0.345 0.348 0.342 0.347
+192 0.360 0.360 0.383 0.367 0.397 0.376 192 0.360 0.360 0.364 0.362 0.360 0.360
+336 0.377 0.374 0.396 0.381 0.386 0.380 336 0.377 0.374 0.378 0.375 0.377 0.374
+720 0.416 0.405 0.429 0.409 0.417 0.409 720 0.416 0.405 0.419 0.408 0.417 0.405
+avg 0.374 0.372 0.393 0.378 0.400 0.384 avg 0.374 0.372 0.376 0.373 0.374 0.372
+$ETTm2$ 96 0.228 0.282 0.246 0.296 0.264 0.326 $ETTm2$ 96 0.228 0.282 0.230 0.286 0.228 0.283
+192 0.262 0.305 0.273 0.313 0.273 0.328 192 0.262 0.305 0.264 0.308 0.262 0.305
+336 0.293 0.328 0.303 0.334 0.297 0.343 336 0.293 0.328 0.298 0.332 0.293 0.328
+720 0.343 0.370 0.343 0.370 0.334 0.369 720 0.343 0.370 0.350 0.373 0.343 0.369
+avg 0.282 0.321 0.291 0.328 0.292 0.341 avg 0.282 0.321 0.285 0.325 0.282 0.321
+Average 0.344 0.370 0.352 0.373 0.373 0.391 Average 0.344 0.370 0.344 0.370 0.345 0.371
+$\bf 1^{st}$ count 30 18 2 $\bf 1^{st}$ count 28 16 21
+Table 16 summarizes the impact of interpolation strategies and image orientations in the Alignment step.
+It shows that the smoother Bilinear and Bicubic interpolation perform similarly, both significantly better than the rougher Nearest Neighbor.
+This suggests that smooth resizing effectively handles time series interpolation.
+Moreover, image orientation has little impact on performance.
+C.8 Hyperparameter analysis
+Figs. 8, 9 and 10 show the influence of three hyperparameters, $r$, $c$, and $L$. We report the MSE averaged on four prediction lengths {96, 192, 336, 720}.
+(a) ETTh1
+(b) ETTh2
+(c) ETTm1
+(d) ETTm2
+(e) Electricity
+(f) Weather
+Figure 8: MSE (Y-axis) performance of different normalization constants $r$ (X-axis).
+(a) ETTh1
+(b) ETTh2
+(c) ETTm1
+(d) ETTm2
+(e) Electricity
+(f) Weather
+Figure 9: MSE (Y-axis) performance of different alignment constants $c$ (X-axis).
+(a) ETTh1
+(b) ETTh2
+(c) ETTm1
+(d) ETTm2
+(e) Electricity
+(f) Weather
+Figure 10: MSE (Y-axis) performance of different context lengths $L$ (X-axis).
+Appendix D Full-Shot Forecasting
+D.1 Training details
+Table 17: Final hyperparameters for VisionTS used in our full-shot forecasting.
+ETTh1 ETTh2 ETTm1 ETTm2 Illness Weather Traffic Electricity
+Normalization constant $r$ 0.4 0.4 0.4 0.4 1.0 1.0 0.4 0.4
+Alignment constant $c$ 0.4 0.4 0.4 0.4 0.4 0.7 0.4 0.4
+Context length $L$ 1152 1152 2304 1152 104 576 1152 1152
+Based on the principle of channel independence (Nie et al., 2022; Han et al., 2024), we treat the variables of each time series as individual data samples.
+We use an Adam optimizer with a learning rate 0.0001 and a batch size 256 to fine-tune MAE.
+All experiments are repeated three times.
+The training epoch is one for all the datasets except Illness, for which we train MAE for 100 epochs with an early stop due to the limited training dataset scale.
+We conduct tuning on validation sets for the three hyperparameters, $r$, $c$, and $L$.
+The final hyperparameters used are summarized in Table 17.
+D.2 Full results and standard deviations
+Table 18: Standard deviations of full-shot experiments.
+Method VisionTS Time-LLM GPT4TS
+Metric MSE MAE MSE MAE MSE MAE
+ETTh1 96 0.347 $\pm$ 0.002 0.376 $\pm$ 0.000 0.376 $\pm$ 0.003 0.402 $\pm$ 0.002 0.370 $\pm$ 0.003 0.389 $\pm$ 0.001
+192 0.385 $\pm$ 0.001 0.400 $\pm$ 0.000 0.407 $\pm$ 0.003 0.421 $\pm$ 0.002 0.412 $\pm$ 0.003 0.413 $\pm$ 0.001
+336 0.407 $\pm$ 0.001 0.415 $\pm$ 0.001 0.430 $\pm$ 0.004 0.438 $\pm$ 0.001 0.448 $\pm$ 0.003 0.431 $\pm$ 0.001
+720 0.439 $\pm$ 0.001 0.443 $\pm$ 0.000 0.457 $\pm$ 0.003 0.468 $\pm$ 0.001 0.441 $\pm$ 0.003 0.449 $\pm$ 0.001
+ETTh2 96 0.269 $\pm$ 0.003 0.328 $\pm$ 0.002 0.286 $\pm$ 0.003 0.346 $\pm$ 0.002 0.280 $\pm$ 0.001 0.335 $\pm$ 0.001
+192 0.332 $\pm$ 0.001 0.374 $\pm$ 0.001 0.361 $\pm$ 0.003 0.391 $\pm$ 0.002 0.348 $\pm$ 0.002 0.380 $\pm$ 0.001
+336 0.351 $\pm$ 0.002 0.395 $\pm$ 0.002 0.390 $\pm$ 0.003 0.414 $\pm$ 0.002 0.380 $\pm$ 0.002 0.405 $\pm$ 0.001
+720 0.390 $\pm$ 0.003 0.430 $\pm$ 0.002 0.405 $\pm$ 0.003 0.434 $\pm$ 0.002 0.406 $\pm$ 0.002 0.436 $\pm$ 0.001
+ETTm1 96 0.281 $\pm$ 0.001 0.322 $\pm$ 0.001 0.291 $\pm$ 0.001 0.341 $\pm$ 0.001 0.300 $\pm$ 0.001 0.340 $\pm$ 0.000
+192 0.322 $\pm$ 0.006 0.353 $\pm$ 0.002 0.341 $\pm$ 0.001 0.369 $\pm$ 0.001 0.343 $\pm$ 0.001 0.368 $\pm$ 0.000
+336 0.356 $\pm$ 0.003 0.379 $\pm$ 0.002 0.359 $\pm$ 0.002 0.379 $\pm$ 0.001 0.376 $\pm$ 0.001 0.386 $\pm$ 0.000
+720 0.391 $\pm$ 0.001 0.413 $\pm$ 0.001 0.433 $\pm$ 0.001 0.419 $\pm$ 0.001 0.431 $\pm$ 0.001 0.416 $\pm$ 0.000
+ETTm2 96 0.169 $\pm$ 0.003 0.256 $\pm$ 0.002 0.162 $\pm$ 0.001 0.248 $\pm$ 0.001 0.163 $\pm$ 0.001 0.249 $\pm$ 0.001
+192 0.225 $\pm$ 0.003 0.294 $\pm$ 0.003 0.235 $\pm$ 0.002 0.304 $\pm$ 0.001 0.222 $\pm$ 0.001 0.291 $\pm$ 0.000
+336 0.278 $\pm$ 0.002 0.334 $\pm$ 0.001 0.280 $\pm$ 0.002 0.329 $\pm$ 0.001 0.273 $\pm$ 0.001 0.327 $\pm$ 0.001
+720 0.372 $\pm$ 0.002 0.392 $\pm$ 0.002 0.366 $\pm$ 0.002 0.382 $\pm$ 0.001 0.357 $\pm$ 0.001 0.376 $\pm$ 0.001
+Weather 96 0.142 $\pm$ 0.000 0.192 $\pm$ 0.001 0.155 $\pm$ 0.001 0.199 $\pm$ 0.001 0.148 $\pm$ 0.001 0.188 $\pm$ 0.000
+192 0.191 $\pm$ 0.000 0.238 $\pm$ 0.000 0.223 $\pm$ 0.001 0.261 $\pm$ 0.001 0.192 $\pm$ 0.001 0.230 $\pm$ 0.000
+336 0.246 $\pm$ 0.003 0.282 $\pm$ 0.001 0.251 $\pm$ 0.001 0.279 $\pm$ 0.001 0.246 $\pm$ 0.001 0.273 $\pm$ 0.000
+720 0.328 $\pm$ 0.004 0.337 $\pm$ 0.001 0.345 $\pm$ 0.001 0.342 $\pm$ 0.001 0.320 $\pm$ 0.001 0.328 $\pm$ 0.000
+Traffic 96 0.344 $\pm$ 0.001 0.236 $\pm$ 0.000 0.392 $\pm$ 0.001 0.267 $\pm$ 0.000 0.396 $\pm$ 0.001 0.264 $\pm$ 0.000
+192 0.372 $\pm$ 0.001 0.249 $\pm$ 0.001 0.409 $\pm$ 0.001 0.271 $\pm$ 0.000 0.412 $\pm$ 0.001 0.268 $\pm$ 0.000
+336 0.383 $\pm$ 0.001 0.257 $\pm$ 0.001 0.434 $\pm$ 0.001 0.296 $\pm$ 0.000 0.421 $\pm$ 0.001 0.273 $\pm$ 0.000
+720 0.422 $\pm$ 0.001 0.280 $\pm$ 0.000 0.451 $\pm$ 0.001 0.291 $\pm$ 0.000 0.455 $\pm$ 0.001 0.291 $\pm$ 0.000
+Electricity 96 0.126 $\pm$ 0.000 0.218 $\pm$ 0.000 0.137 $\pm$ 0.000 0.233 $\pm$ 0.000 0.141 $\pm$ 0.000 0.239 $\pm$ 0.000
+192 0.146 $\pm$ 0.001 0.239 $\pm$ 0.001 0.152 $\pm$ 0.000 0.247 $\pm$ 0.000 0.158 $\pm$ 0.000 0.253 $\pm$ 0.000
+336 0.161 $\pm$ 0.001 0.255 $\pm$ 0.001 0.169 $\pm$ 0.000 0.267 $\pm$ 0.000 0.172 $\pm$ 0.000 0.266 $\pm$ 0.000
+720 0.193 $\pm$ 0.000 0.286 $\pm$ 0.000 0.200 $\pm$ 0.000 0.290 $\pm$ 0.000 0.207 $\pm$ 0.000 0.293 $\pm$ 0.000
+$\bf 1^{st}$ count 42 2 12
+Table 19: Full results of Table 4: Full-shot forecasting performance on the long-term TSF benchmark. VisionTS is fine-tuned only a single epoch on each dataset except for Illness.
 Pretrain [Uncaptioned image] Images [Uncaptioned image] Text [Uncaptioned image] No Pretrain
@@
 $\bf 1^{st}$ count 46 4 12 0 19 0 0 0 0 0 0
-Setups
-We evaluate the full-shot capability of each baseline trained on individual long-term TSF benchmarks.
-In addition to the six datasets used for zero-shot forecasting, we also include the popular Traffic and Illness datasets.
-As self-attention and feed-forward layers contain rich knowledge that can be transferred to TSF, we choose to fine-tune only the layer normalization (LN) layers while freezing the other parameters, which is also adopted by Zhou et al.
-(2023).
-Training details are elaborated in Section C.1.
-Main Results
-Table 3 summarizes the full-shot results, with standard deviations detailed in Section C.2.
-It shows that VisionTS outperforms other baselines in most cases (46 out of 80), surpassing the non-pretrained PatchTST and the language-pretrained GPT4TS.
-Remarkably, except for Illness with the least data, VisionTS demands only a single epoch of fine-tuning.
-This suggests that even minimal fine-tuning enables VisionTS to effectively adapt to time series.
-Compared with Table 1, fine-tuning provides limited benefits for ETTh1 and ETTh2 but leads to significant improvements for other datasets.
-We attribute this to the smaller data scale of ETTh1 and ETTh2.
-- w/o VM VM2Attn VM2Trsf Rand-VM
-ETTh1 0.395 0.785 0.448 0.459 0.534
-ETTh2 0.336 0.420 0.418 0.448 0.411
-ETTm1 0.338 0.676 0.397 0.398 0.433
-ETTm2 0.261 0.379 0.274 0.292 0.288
-Avg. 0.333 0.565 0.384 0.399 0.417
-Table 4: MSE results for ablation studies, averaged on four prediction lengths.
-All LN Bias MLP Attn Freeze
-ETTh1 0.534 0.395 0.401 0.534 0.554 0.419
-ETTh2 0.411 0.336 0.347 0.401 0.392 0.340
-ETTm1 0.433 0.338 0.343 0.441 0.444 0.374
-ETTm2 0.288 0.261 0.256 0.292 0.289 0.305
-Avg. 0.417 0.333 0.337 0.417 0.420 0.360
-Table 5: MSE for different fine-tuning strategies, averaged on four prediction lengths.
-Ablation Study
-Tan et al.
-(2024) proposed several ablation variants for text-based foundation models, including w/o LLM (removing the LLM), LLM2Attn/LLM2Trsf (replacing the LLM with a single self-attention/Transformer layer), and RandLLM (randomly initializing the LLM).
-They found no significant performance differences and concluded that textual knowledge is not necessary for TSF.
-We conducted similar ablations to assess the role of the vision model (VM), including w/o VM, VM2Attn, VM2Trsf, and RandVM.
-Table 5 with full results in Section C.3 shows that these variants lead to worse performance, indicating that visual knowledge is beneficial for TSF.
-Analysis: Fine-tuning strategies
-As stated before, we fine-tune only the layer normalization (LN).
-We also tested fine-tuning the bias, MLP, or attention layers, in addition to full fine-tuning and freezing.
-All hyperparameters were kept constant.
-Note that freezing differs from the previous zero-shot experiment, where a longer context length was used.
-Table 5 with full results in Section C.3 show that fine-tuning LN is the best.
-Modifying MLP or attention layers results in significant performance drops, suggesting that valuable knowledge resides in these components.
-5 Related Work
-Depending on the pre-training data, TSF foundation models can be categorized into Text-based and TS-based.
-We firstly review these related work, then introduce recent research for image-based time series analysis.
-Text-based TSF Foundation Models
-Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks.
-For example, Zhou et al.
-(2023) fine-tuned a pre-trained GPT (Radford et al., 2019) on each time-series downstream task, such as forecasting, classification, imputation, and anomaly detection.
-Based on Llama (Touvron et al., 2023), Jin et al.
-(2024) froze the pre-trained LLM and reprogrammed the time series to align with the language modality.
-Bian et al.
-(2024) adopted a two-stage approach by continually pre-training GPT (Radford et al., 2019) on the time-series domain.
-Nevertheless, the TSF performance of LLMs has recently been questioned by Tan et al.
-(2024), which designed several ablation studies to show that textual knowledge is unnecessary for forecasting.
-In this paper, we attribute it to the large modality gap.
-Some recent approaches focus on directly transforming the time series into natural texts for LLMs, allowing for zero-shot forecasting.
-For example, PromptCast (Xue & Salim, 2023) uses pre-defined templates to describe numerical time series data, while LLMTime (Gruver et al., 2023) directly separates time steps using commas and separates digits using spaces to construct the text input.
-However, due to the efficiency issue caused by the autoregressive decoding strategy, their practical use is limited.
-Time Series-Based TSF Foundation Models
-Self-supervised pre-training a TSF model on the same dataset used for downstream TSF tasks is a well-explored topic (Ma et al., 2023; Zhang et al., 2024), such as denoising autoencoders (Zerveas et al., 2021) or contrastive learning (Woo et al., 2022a; Yue et al., 2022).
-They follow a similar paradigm to the masked autoencoder (MAE) in computer vision.
-However, these methods rarely examine the cross-dataset generalization capabilities.
-Recently, research has shifted towards training universal foundation models, by collecting large-scale time series datasets from diverse domains (Goswami et al., 2024; Liu et al., 2024; Das et al., 2024; Dong et al., 2024; Feng et al., 2024) or generating numerous synthetic time series data (Yang et al., 2024).
-As a representative method, Woo et al.
-(2024) collected 27 billion observations across nine domains and trained TSF foundation models of various scales, achieving strong zero-shot performance.
-However, given the severe heterogeneity, constructing high-quality large datasets poses significant challenges for building these foundation models.
-Image-Based Time-Series Analysis
-Previous research has investigated encoding time series data into images and used convolutional neural networks (CNNs) trained from scratch for classification (Wang & Oates, 2015a; b; Hatami et al., 2018) or forecasting (Li et al., 2020; Sood et al., 2021; Semenoglou et al., 2023).
-Recent researchers explored using pre-trained models for these imaging time series.
-Li et al.
-(2024) used a pre-trained vision transformer (ViT) for classification.
-Wimmer & Rekabsaz (2023) and Zhang et al.
-(2023) employed vision-language multimodal pre-trained models to extract predictive features and generate text descriptions.
-However, these studies did not deeply examine the transferability from natural images to TSF.
-Despite early efforts by Zhou et al.
-(2023) to fine-tune a BEiT (Bao et al., 2022) trained on images for time series forecasting, it still falls short of the leading text-based and TS-based TSF foundation models.
-To the best of our knowledge, we are the first to show that an image-based foundation model, without further time-series adaptation, can match or even surpass other types of TSF foundation models.
-6 Conclusion
-In this paper, we explore a novel approach to building a time series forecasting (TSF) foundation model using natural images, offering a new perspective distinct from the traditional text-based and TS-based methods.
-By leveraging the intrinsic similarities between images and time series, we introduced VisionTS, an MAE-based TSF foundation model that reformulates the TSF task as an image reconstruction problem.
-Our extensive evaluations demonstrate that VisionTS achieves outstanding forecasting performance in zero-shot and full-shot settings, being a free lunch for a TSF foundation model.
-We hope our findings could open new avenues for further cross-domain research.
-Limitations and Future Directions.
-As a preliminary study, we employed a basic MAE model and reformulated TSF as a patch-based image reconstruction task.
-Reformulating TSF as a more refined image inpainting task and utilizing more advanced models like diffusion models (Rombach et al., 2022; Peebles & Xie, 2023) presents a promising research direction.
-Appendix A Details of Experiments
-Table 6: The statistics of 8 long-term TSF datasets.
-Dataset ETTh1 ETTh2 ETTm1 ETTm2 Electricity Traffic Weather Illness
-Variates 7 7 7 7 321 862 21 7
-Timesteps 17,420 17,420 69,680 69,680 26,304 17,544 52,695 966
-Frequency 1hour 1hour 15min 15min 1hour 1hour 10min 1week
-Periodicity 24 24 96 96 24 24 144 52
-Long-term TSF Benchmark
-We evaluate our model on 8 widely used long-term TSF datasets (Zhou et al., 2021; Wu et al., 2021), summarized in Table 6. We also report the periodicity of each dataset for our segmentation step.
-  1. 1.
-    ETTh1/ETTh2/ETTm1/ETTm2 record 7 indicators from electricity transformers between July 2016 and July 2018. Data is collected hourly for ETTh1/ETTh2 and every 15 minutes for ETTm1/ETTm2.
-  2. 2.
-    Electricity contains hourly electricity consumption (in KWh) data from 321 customers between 2012 and 2014.
-  3. 3.
-    Traffic provides hourly road occupancy rates measured by various sensors on San Francisco Bay area freeways over 2 years, sourced from the California Department of Transportation.
-  4. 4.
-    Illness tracks 7 weekly indicators related to influenza-like illness from the CDC in the United States, spanning 2002 to 2021.
-  5. 5.
-    Weather includes 21 meteorological indicators, such as temperature and humidity, recorded every 10 minutes throughout 2020.
-Monash Benchmark
-Following Woo et al.
-(2024), we tested 29 Monash datasets (Godahewa et al., 2021) using GluonTS (Alexandrov et al., 2020), including M1 Monthly, M3 Monthly, M3 Other, M4 Monthly, M4 Weekly, M4 Daily, M4 Hourly, Tourism Quarterly, Tourism Monthly, CIF 2016, Australian Electricity Demand, Bitcoin, Pedestrian Counts, Vehicle Trips, KDD Cup, Weather, NN5 Daily, NN5 Weekly, Carparts, FRED-MD, Traffic Hourly, Traffic Weekly, Rideshare, Hospital, COVID Deaths, Temperature Rain, Sunspot, Saugeen River Flow, and US Births.
-Baselines
-The baseline models selected for comparison are briefly described below:
-  1. 1.
-    Moirai (Woo et al., 2024) is a TSF foundation model trained on the Large-scale Open Time Series Archive (LOTSA), with over 27B observations across nine domains.
-It has three variants: small, base, and large.
-  2. 2.
-    TimesFM (Das et al., 2024) is a decoder-style TSF foundation model, using a large time-series corpus comprising both real-world and synthetic datasets.
-  3. 3.
-    Time-LLM (Jin et al., 2024) is a text-based TSF foundation model built on Llama, which reprograms time series data to align with the language modality, keeping the LLM frozen.
-  4. 4.
-    GPT4TS (Zhou et al., 2023) (OneFitsAll) is another text-based model based on GPT, fine-tuned for forecasting tasks.
-  5. 5.
-    LLMTime (Gruver et al., 2023) encodes time series data to a text sequence, supporting zero-shot forecasting.
-  6. 6.
-    DLinear (Zeng et al., 2023) proposes a linear forecasting model, enhanced by seasonal-trend decomposition or normalization.
-  7. 7.
-    PatchTST (Nie et al., 2022) uses Transformer encoders with patching and channel independence techniques for improved predictions.
-  8. 8.
-    TimesNet (Wu et al., 2023) applies convolution kernels along the time dimension, using temporal decomposition and periodical segmentation to capture temporal patterns.
-  9. 9.
-    FEDformer (Zhou et al., 2022) employs a sparse frequency domain representation, using frequency-enhanced blocks for cross-time dependency.
-  10. 10.
-    Autoformer (Wu et al., 2021) uses series decomposition blocks and Auto-Correlation to capture cross-time dependency.
-  11. 11.
-    Stationary (Liu et al., 2022) introduces stationarization and de-stationary attention mechanisms.
-  12. 12.
-    ETSFormer (Woo et al., 2022b) leverages exponential smoothing principles, including exponential smoothing and frequency attention mechanisms.
-  13. 13.
-    Informer (Zhou et al., 2021) proposes ProbSparse self-attention and distillation operations.
-For the long-term TSF benchmark, we include TS-based foundation model results from their original papers, Text-based model results from Tan et al.
-(2024), and other baseline results from Zhou et al.
-(2023).
-For the Monash benchmark, we include results from Woo et al.
-(2024).
-Metrics and Environment
-Performance is assessed using Mean Squared Error (MSE) and Mean Absolute Error (MAE), with lower values indicating better forecasting accuracy.
-All experiments are conducted using Time-Series-Library (https://github.com/thuml/Time-Series-Library) on an NVIDIA A800 GPU.
-Appendix B Zero-Shot Forecasting
-B.1 Hyperparameters
-Table 7: Hyperparameters for VisionTS used in our zero-shot forecasting (Informer benchmark).
-ETTh1 ETTh2 ETTm1 ETTm2 Weather Electricity
-Normalization constant $r$ 0.4 0.4 0.4 0.4 0.4 0.4
-Alignment constant $c$ 0.4 0.4 0.4 0.4 0.4 0.4
-Context length $L$ 2880 1728 2304 4032 4032 2880
-We conduct hyperparameter tuning on validation sets to determine the optimal context length $L$. Final used hyperparameters are summarized in Table 7.
-B.2 Full forecasting results of the long-term TSF benchmark
-Table 8: Full results of Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Bold: the best result.
-[Uncaptioned image] Zero-Shot [Uncaptioned image] Few-Shot (10% Downstream Dataset)
-Pretrain [Uncaptioned image] Images [Uncaptioned image] Time-series [Uncaptioned image] Text [Uncaptioned image] No Pretrain
-Method VisionTime MoiraiS MoiraiB MoiraiL TimeLLM GPT4TS DLinear PatchTST TimesNet Autoformer Informer
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-$ETTh1$ 96 0.353 0.383 0.375 0.402 0.384 0.402 0.380 0.398 0.448 0.460 0.458 0.456 0.492 0.495 0.516 0.485 0.861 0.628 0.613 0.552 1.179 0.792
-192 0.392 0.410 0.399 0.419 0.425 0.429 0.440 0.434 0.484 0.483 0.570 0.516 0.565 0.538 0.598 0.524 0.797 0.593 0.722 0.598 1.199 0.806
-336 0.407 0.423 0.412 0.429 0.456 0.450 0.514 0.474 0.589 0.540 0.608 0.535 0.721 0.622 0.657 0.550 0.941 0.648 0.750 0.619 1.202 0.811
-720 0.406 0.441 0.413 0.444 0.470 0.473 0.705 0.568 0.700 0.604 0.725 0.591 0.986 0.743 0.762 0.610 0.877 0.641 0.721 0.616 1.217 0.825
-avg 0.390 0.414 0.400 0.424 0.434 0.439 0.510 0.469 0.556 0.522 0.590 0.525 0.691 0.600 0.633 0.542 0.869 0.628 0.702 0.596 1.199 0.809
-$ETTh2$ 96 0.271 0.328 0.281 0.334 0.277 0.327 0.287 0.325 0.275 0.326 0.331 0.374 0.357 0.411 0.353 0.389 0.378 0.409 0.413 0.451 3.837 1.508
-192 0.328 0.367 0.340 0.373 0.340 0.374 0.347 0.367 0.374 0.373 0.402 0.411 0.569 0.519 0.403 0.414 0.490 0.467 0.474 0.477 3.856 1.513
-336 0.345 0.381 0.362 0.393 0.371 0.401 0.377 0.393 0.406 0.429 0.406 0.433 0.671 0.572 0.426 0.441 0.537 0.494 0.547 0.543 3.952 1.526
-720 0.388 0.422 0.380 0.416 0.394 0.426 0.404 0.421 0.427 0.449 0.449 0.464 0.824 0.648 0.477 0.480 0.510 0.491 0.516 0.523 3.842 1.503
-avg 0.333 0.375 0.341 0.379 0.346 0.382 0.354 0.377 0.370 0.394 0.397 0.421 0.605 0.538 0.415 0.431 0.479 0.465 0.488 0.499 3.872 1.513
-$ETTm1$ 96 0.341 0.347 0.404 0.383 0.335 0.360 0.353 0.363 0.346 0.388 0.390 0.404 0.352 0.392 0.410 0.419 0.583 0.501 0.774 0.614 1.162 0.785
-192 0.360 0.360 0.435 0.402 0.366 0.379 0.376 0.380 0.373 0.416 0.429 0.423 0.382 0.412 0.437 0.434 0.630 0.528 0.754 0.592 1.172 0.793
-336 0.377 0.374 0.462 0.416 0.391 0.394 0.399 0.395 0.413 0.426 0.469 0.439 0.419 0.434 0.476 0.454 0.725 0.568 0.869 0.677 1.227 0.908
-720 0.416 0.405 0.490 0.437 0.434 0.419 0.432 0.417 0.485 0.476 0.569 0.498 0.490 0.477 0.681 0.556 0.769 0.549 0.810 0.630 1.207 0.797
-avg 0.374 0.372 0.448 0.410 0.382 0.388 0.390 0.389 0.404 0.427 0.464 0.441 0.411 0.429 0.501 0.466 0.677 0.537 0.802 0.628 1.192 0.821
-$ETTm2$ 96 0.228 0.282 0.205 0.282 0.195 0.269 0.189 0.260 0.177 0.261 0.188 0.269 0.213 0.303 0.191 0.274 0.212 0.285 0.352 0.454 3.203 1.407
-192 0.262 0.305 0.261 0.318 0.247 0.303 0.247 0.300 0.241 0.314 0.251 0.309 0.278 0.345 0.252 0.317 0.270 0.323 0.694 0.691 3.112 1.387
-336 0.293 0.328 0.319 0.355 0.291 0.333 0.295 0.334 0.274 0.327 0.307 0.346 0.338 0.385 0.306 0.353 0.323 0.353 2.408 1.407 3.255 1.421
-720 0.343 0.370 0.415 0.410 0.355 0.377 0.372 0.386 0.417 0.390 0.426 0.417 0.436 0.440 0.433 0.427 0.474 0.449 1.913 1.166 3.909 1.543
-avg 0.282 0.321 0.300 0.341 0.272 0.321 0.276 0.320 0.277 0.323 0.293 0.335 0.316 0.368 0.296 0.343 0.320 0.353 1.342 0.930 3.370 1.440
-$Electricity$ 96 0.177 0.266 0.205 0.299 0.158 0.248 0.152 0.242 0.139 0.241 0.139 0.237 0.150 0.253 0.140 0.238 0.299 0.373 0.261 0.348 1.259 0.919
-192 0.188 0.277 0.220 0.310 0.174 0.263 0.171 0.259 0.151 0.248 0.156 0.252 0.164 0.264 0.160 0.255 0.305 0.379 0.338 0.406 1.160 0.873
-336 0.207 0.296 0.236 0.323 0.191 0.278 0.192 0.278 0.169 0.270 0.175 0.270 0.181 0.282 0.180 0.276 0.319 0.391 0.410 0.474 1.157 0.872
-720 0.256 0.337 0.270 0.347 0.229 0.307 0.236 0.313 0.240 0.322 0.233 0.317 0.223 0.321 0.241 0.323 0.369 0.426 0.715 0.685 1.203 0.898
-avg 0.207 0.294 0.233 0.320 0.188 0.274 0.188 0.273 0.175 0.270 0.176 0.269 0.180 0.280 0.180 0.273 0.323 0.392 0.431 0.478 1.195 0.891
-$Weather$ 96 0.220 0.257 0.173 0.212 0.167 0.203 0.177 0.208 0.161 0.210 0.163 0.215 0.171 0.224 0.165 0.215 0.184 0.230 0.221 0.297 0.374 0.401
-192 0.244 0.275 0.216 0.250 0.209 0.241 0.219 0.249 0.204 0.248 0.210 0.254 0.215 0.263 0.210 0.257 0.245 0.283 0.270 0.322 0.552 0.478
-336 0.280 0.299 0.260 0.282 0.256 0.276 0.277 0.292 0.261 0.302 0.256 0.292 0.258 0.299 0.259 0.297 0.305 0.321 0.320 0.351 0.724 0.541
-720 0.330 0.337 0.320 0.322 0.321 0.323 0.365 0.350 0.309 0.332 0.321 0.339 0.320 0.346 0.332 0.346 0.381 0.371 0.390 0.396 0.739 0.558
-avg 0.269 0.292 0.242 0.267 0.238 0.261 0.260 0.275 0.234 0.273 0.238 0.275 0.241 0.283 0.242 0.279 0.279 0.301 0.300 0.342 0.597 0.495
-Average 0.309 0.345 0.327 0.357 0.310 0.344 0.329 0.350 0.336 0.368 0.360 0.378 0.407 0.416 0.378 0.389 0.491 0.446 0.678 0.579 1.904 0.995
-$\bf 1^{st}$ count 32 0 10 8 10 6 0 0 0 0 0
-Table 8 shows the full results of zero-shot/few-shot long-term forecasting performance.
-VisionTS achieves the best results in most cases (32 out of 62), outperforming MoiraiBase (10 out of 62) and MoiraiLarge (8 out of 62).
-B.3 Comparison of TimesFM and LLMTime
-Table 9: Comparison of TimesFM and LLMTime for zero-shot forecasting in terms of MAE.
-Owing to expensive evaluations for LLMTime, the results are reported on the last test window of the original test split, as done in Das et al.
-(2024); Gruver et al.
-(2023).
-Method VisionTS TimesFM LLMTime
-ETTh1 96 0.35 0.45 0.42
-192 0.45 0.53 0.50
-ETTh2 96 0.24 0.35 0.33
-192 0.60 0.62 0.70
-ETTm1 96 0.12 0.19 0.37
-192 0.23 0.26 0.71
-ETTm2 96 0.19 0.24 0.29
-192 0.24 0.27 0.31
-Average 0.30 0.36 0.45
-Table 10: Comparison of traditional zero-shot forecasting baselines.
-Method VisionTS ETS ARIMA Seasonal Naïve Seasonal Avg
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-$ETTh1$ 96 0.353 0.383 1.289 0.710 0.900 0.719 0.512 0.433 0.589 0.585
-192 0.392 0.410 1.319 0.730 0.906 0.724 0.581 0.469 0.598 0.590
-336 0.407 0.423 1.324 0.742 0.908 0.731 0.650 0.501 0.610 0.597
-720 0.406 0.441 1.329 0.751 0.932 0.753 0.655 0.514 0.656 0.624
-avg 0.390 0.414 1.315 0.733 0.912 0.732 0.600 0.479 0.613 0.599
-$ETTh2$ 96 0.271 0.328 0.399 0.408 0.488 0.508 0.391 0.380 0.457 0.494
-192 0.328 0.367 0.500 0.459 0.497 0.514 0.482 0.429 0.466 0.500
-336 0.345 0.381 0.562 0.498 0.507 0.522 0.532 0.466 0.476 0.509
-720 0.388 0.422 0.558 0.506 0.572 0.557 0.525 0.474 0.542 0.548
-avg 0.333 0.375 0.505 0.468 0.516 0.525 0.483 0.437 0.485 0.513
-$ETTm1$ 96 0.341 0.347 1.204 0.659 0.702 0.568 0.423 0.387 0.369 0.399
-192 0.360 0.360 1.251 0.685 0.704 0.570 0.463 0.406 0.374 0.402
-336 0.377 0.374 1.276 0.702 0.709 0.574 0.496 0.426 0.382 0.407
-720 0.416 0.405 1.311 0.724 0.713 0.580 0.574 0.464 0.394 0.416
-avg 0.374 0.372 1.261 0.693 0.707 0.573 0.489 0.421 0.380 0.406
-$ETTm2$ 96 0.228 0.282 0.257 0.324 0.397 0.434 0.263 0.301 0.365 0.411
-192 0.262 0.305 0.331 0.366 0.402 0.436 0.321 0.337 0.369 0.414
-336 0.293 0.328 0.402 0.406 0.407 0.439 0.376 0.370 0.375 0.418
-720 0.343 0.370 0.512 0.462 0.413 0.443 0.471 0.422 0.380 0.423
-avg 0.282 0.321 0.376 0.390 0.405 0.438 0.358 0.357 0.372 0.417
-Average 0.344 0.370 0.864 0.571 0.635 0.567 0.482 0.424 0.463 0.484
-$\bf 1^{st}$ count 41 0 0 0 1
-Due to the step-by-step output of the decoder architecture, the efficiency of TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023) are relatively slower.
-Thus, Das et al.
-(2024) only reported results for the last test window of the original split.
-We compared VisionTS with their results under the same setting, as shown in Table 9.
-VisionTS outperforms TimesFM and LLMTime in terms of MAE, indicating that image-based TSF models are on par with or even better than TS-based and text-based models.
-B.4 Comparison of traditional methods
-In addition to deep learning models, we also compare traditional methods, including ARIMA, ETS, and two methods that require periodicity as our VisionTS: Seasonal Naïve (repeating the last period) and Seasonal Avg (similar to Seasonal Naïve but repeating the average of all periods in the look-back window).
-Due to the high computational cost of ARIMA and ETS, we only compare them on the small-scale benchmarks, i.e., four ETT datasets.
-Table 10 shows that VisionTS also achieves the best performance.
-B.5 Full forecasting results of the Monash TSF benchmark
-Table 11: The statistics of 29 Monash TSF datasets. $P=1$ indicates no significant periodicity.
-Datasets Frequency Periodiciy ($P$)
-Rideshare H 1
-M4 Daily, Bitcoin, COVID Deaths D 1
-M4 Weekly, NN5 Weekly, Traffic Weekly W 1
-FRED-MD M 1
-M3 Other Q 2
-CIF 2016 (horizon=6) M 3
-Tourism Quarterly Q 4
-M3 Monthly, M4 Monthly, CIF 2016 (horizon=12), Car Parts M 6
-Vehicle Trips, NN5 Daily D 7
-M1 Monthly, Tourism Monthly, Hospital M 12
-M4 Hourly, Traffic Hourly H 24
-Sunspot, Saugeen River Flow D 30
-Pedestrian Counts H 48
-KDD cup H 96
-Aus. Elec. Demand 30T 336
-Weather, Temperature Rain, US Births D 365
-Table 12: Full results of Fig. 4: Forecasting results (MAE) on the Monash TSF benchmark. We reported the reproduction results of LLMTime based on the GPT3.5 API from Woo et al. (2024).
-VisionTS LLMTime MoiraiSmall Naive SES Theta TBATS ETS (DHR-)ARIMA PR CatBoost FFNN DeepAR N-BEATS WaveNet Transformer
-M1 Monthly 1987.69 2562.84 2082.26 2707.75 2259.04 2166.18 2237.5 1905.28 2080.13 2088.25 2052.32 2162.58 1860.81 1820.37 2184.42 2723.88
-M3 Monthly 741 877.97 713.41 837.14 743.41 623.71 630.59 626.46 654.8 692.97 732 692.48 728.81 648.6 699.3 798.38
-M3 Other 283.76 300.3 263.54 278.43 277.83 215.35 189.42 194.98 193.02 234.43 318.13 240.17 247.56 221.85 245.29 239.24
-M4 Monthly 654.27 728.27 597.6 671.27 625.24 563.58 589.52 582.6 575.36 596.19 611.69 612.52 615.22 578.48 655.51 780.47
-M4 Weekly 314.83 518.44 339.76 347.99 336.82 333.32 296.15 335.66 321.61 293.21 364.65 338.37 351.78 277.73 359.46 378.89
-M4 Daily 215.63 266.52 189.1 180.83 178.27 178.86 176.6 193.26 179.67 181.92 231.36 177.91 299.79 190.44 189.47 201.08
-M4 Hourly 288.37 576.06 268.04 1218.06 1218.06 1220.97 386.27 3358.1 1310.85 257.39 285.35 385.49 886.02 425.75 393.63 320.54
-Tourism Quarterly 12931.88 16918.9 18352.44 15845.1 15014.19 7656.49 9972.42 8925.52 10475.5 9092.58 10268 8981.04 9511.37 8640.56 9137.12 9521.67
-Tourism Monthly 2560.19 5608.61 3569.85 5636.83 5302.1 2069.96 2940.08 2004.51 2536.77 2187.28 2537.04 2022.21 1871.69 2003.02 2095.13 2146.98
-CIF 2016 570907.24 599314 655888.58 578597 581875.97 714819 855578 642421 469059 563205.6 603551 1495923 3200418 679035 5998225 4057973
-Aus. Elec. Demand 237.44 760.81 266.57 659.6 659.6 665.04 370.74 1282.99 1045.92 247.18 241.77 258.76 302.41 213.83 227.5 231.45
-Bitcoin 2.34E+18 1.7E+18 1.76E+18 7.8E+17 5.33E+18 5.3E+18 9.9E+17 1.1E+18 3.6E+18 6.66E+17 1.9E+18 1.5E+18 2E+18 1.1E+18 2.5E+18 2.6E+18
-Pedestrian Counts 46.37 97.77 54.88 170.88 170.87 170.94 222.38 216.5 635.16 44.18 43.41 46.41 44.78 66.84 46.46 47.29
-Vehicle Trips 22.08 31.48 24.46 31.42 29.98 30.76 21.21 30.95 30.07 27.24 22.61 22.93 22 28.16 24.15 28.01
-KDD cup 35.13 42.72 39.81 42.13 42.04 42.06 39.2 44.88 52.2 36.85 34.82 37.16 48.98 49.1 37.08 44.46
-Weather 2.07 2.17 1.96 2.36 2.24 2.51 2.3 2.35 2.45 8.17 2.51 2.09 2.02 2.34 2.29 2.03
-NN5 Daily 3.52 7.1 5.37 8.26 6.63 3.8 3.7 3.72 4.41 5.47 4.22 4.06 3.94 4.92 3.97 4.16
-NN5 Weekly 14.67 15.76 15.07 16.71 15.66 15.3 14.98 15.7 15.38 14.94 15.29 15.02 14.69 14.19 19.34 20.34
-Carparts 0.58 0.44 0.53 0.65 0.55 0.53 0.58 0.56 0.56 0.41 0.53 0.39 0.39 0.98 0.4 0.39
-FRED-MD 1893.67 2804.64 2568.48 2825.67 2798.22 3492.84 1989.97 2041.42 2957.11 8921.94 2475.68 2339.57 4264.36 2557.8 2508.4 4666.04
-Traffic Hourly 0.01 0.03 0.02 0.03 0.03 0.03 0.04 0.03 0.04 0.02 0.02 0.01 0.01 0.02 0.02 0.01
-Traffic Weekly 1.2 1.15 1.17 1.19 1.12 1.13 1.17 1.14 1.22 1.13 1.17 1.15 1.18 1.11 1.2 1.42
-Rideshare 4.6 6.28 1.35 6.29 6.29 7.62 6.45 6.29 3.37 6.3 6.07 6.59 6.28 5.55 2.75 6.29
-Hospital 19.36 25.68 23 24.07 21.76 18.54 17.43 17.97 19.6 19.24 19.17 22.86 18.25 20.18 19.35 36.19
-COVID Deaths 137.51 653.31 124.32 353.71 353.71 321.32 96.29 85.59 85.77 347.98 475.15 144.14 201.98 158.81 1049.48 408.66
-Temperature Rain 5.88 6.37 5.3 9.39 8.18 8.22 7.14 8.21 7.19 6.13 6.76 5.56 5.37 7.28 5.81 5.24
-Sunspot 2.81 5.07 0.11 3.93 4.93 4.93 2.57 4.93 2.57 3.83 2.27 7.97 0.77 14.47 0.17 0.13
-Saugeen River Flow 24.47 34.84 24.07 21.5 21.5 21.49 22.26 30.69 22.38 25.24 21.28 22.98 23.51 27.92 22.17 28.06
-US Births 820.21 1374.99 872.51 1152.67 1192.2 586.93 399 419.73 526.33 574.93 441.7 557.87 424.93 422 504.4 452.87
-Normalized MAE 0.714 1.041 0.657 1.000 1.028 0.927 0.758 0.872 0.898 0.785 0.760 0.741 0.759 0.783 0.749 0.770
-Rank 2 16 1 14 15 13 5 11 12 10 7 3 6 9 4 8
-Setup
-Table 11 lists the sampling frequency and the selected period $P$ for each dataset.
-Datasets with $P=1$ indicate no significant periodicity, where we use a context length of $L=300$.
-For other datasets with $P>1$, we select a longer context length of $L=1000$.
-All datasets were tested with the hyperparameters $r=c=0.4$ as we had done for the long-term TSF benchmark.
-Results
-Table 12 presents VisionTS ’s MAE test results, with the normalized MAE calculated by dividing each dataset’s MAE by the naive forecast’s MAE and aggregated using the geometric mean across datasets.
-We include the result of each baseline from Woo et al.
-(2024).
-Particularly, we find that VisionTS outperforms Moirai on some datasets with $P=1$ (e.g., FRED-MD and M4 Weekly), showing that VisionTS can still work effectively without significant periodicity.
-B.6 Impact of MAE’s parameter size
-Table 13: Full results of Table 2: zero-shot forecasting results of different MAE variants. Bold: best results among three variants. We also include the results from Moirai for reference.
-Method MAE (Base) MAE (Large) MAE (Huge) Moirai (Small) Moirai (Base) Moirai (Huge)
-112M 330M 657M 14M 91M 311M
-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-$ETTh1$ 96 0.353 0.383 0.346 0.382 0.362 0.384 0.375 0.402 0.384 0.402 0.380 0.398
-192 0.392 0.410 0.379 0.406 0.407 0.414 0.399 0.419 0.425 0.429 0.440 0.434
-336 0.407 0.423 0.391 0.416 0.399 0.419 0.412 0.429 0.456 0.450 0.514 0.474
-720 0.406 0.441 0.397 0.433 0.395 0.433 0.413 0.444 0.470 0.473 0.705 0.568
-avg 0.390 0.414 0.378 0.409 0.391 0.412 0.400 0.424 0.434 0.439 0.510 0.469
-$ETTh2$ 96 0.271 0.328 0.286 0.334 0.285 0.333 0.281 0.334 0.277 0.327 0.287 0.325
-192 0.328 0.367 0.346 0.375 0.337 0.369 0.340 0.373 0.340 0.374 0.347 0.367
-336 0.345 0.381 0.356 0.387 0.357 0.388 0.362 0.393 0.371 0.401 0.377 0.393
-720 0.388 0.422 0.371 0.409 0.379 0.412 0.380 0.416 0.394 0.426 0.404 0.421
-avg 0.333 0.375 0.340 0.377 0.339 0.375 0.341 0.379 0.346 0.382 0.354 0.377
-$ETTm1$ 96 0.341 0.347 0.344 0.349 0.352 0.351 0.404 0.383 0.335 0.360 0.353 0.363
-192 0.360 0.360 0.365 0.363 0.360 0.367 0.435 0.402 0.366 0.379 0.376 0.380
-336 0.377 0.374 0.381 0.376 0.381 0.383 0.462 0.416 0.391 0.394 0.399 0.395
-720 0.416 0.405 0.429 0.411 0.440 0.412 0.490 0.437 0.434 0.419 0.432 0.417
-avg 0.374 0.372 0.379 0.375 0.383 0.378 0.448 0.410 0.382 0.388 0.390 0.389
-$ETTm2$ 96 0.228 0.282 0.225 0.282 0.229 0.282 0.205 0.282 0.195 0.269 0.189 0.260
-192 0.262 0.305 0.262 0.305 0.265 0.306 0.261 0.318 0.247 0.303 0.247 0.300
-336 0.293 0.328 0.299 0.331 0.286 0.324 0.319 0.355 0.291 0.333 0.295 0.334
-720 0.343 0.370 0.358 0.377 0.355 0.374 0.415 0.410 0.355 0.377 0.372 0.386
-avg 0.282 0.321 0.286 0.324 0.284 0.322 0.300 0.341 0.272 0.321 0.276 0.320
-$Electricity$ 96 0.177 0.266 0.177 0.268 0.170 0.259 0.205 0.299 0.158 0.248 0.152 0.242
-192 0.188 0.277 0.192 0.283 0.182 0.273 0.220 0.310 0.174 0.263 0.171 0.259
-336 0.207 0.296 0.213 0.303 0.207 0.295 0.236 0.323 0.191 0.278 0.192 0.278
-720 0.256 0.337 0.256 0.337 0.250 0.333 0.270 0.347 0.229 0.307 0.236 0.313
-avg 0.207 0.294 0.209 0.298 0.202 0.290 0.233 0.320 0.188 0.274 0.188 0.273
-$Weather$ 96 0.220 0.257 0.222 0.257 0.235 0.265 0.173 0.212 0.167 0.203 0.177 0.208
-192 0.244 0.275 0.246 0.275 0.276 0.288 0.216 0.250 0.209 0.241 0.219 0.249
-336 0.280 0.299 0.283 0.301 0.304 0.309 0.260 0.282 0.256 0.276 0.277 0.292
-720 0.330 0.337 0.338 0.343 0.351 0.350 0.320 0.322 0.321 0.323 0.365 0.350
-avg 0.269 0.292 0.272 0.294 0.292 0.303 0.242 0.267 0.238 0.261 0.260 0.275
-Average 0.309 0.345 0.311 0.346 0.315 0.347 0.327 0.357 0.310 0.344 0.329 0.350
-$\bf 1^{st}$ count 38 17 17 - - -
-Table 13 compares zero-shot forecasting performance of three MAE variants (112M, 330M, and 657M), showing that the three variants are similar, but larger models show a slight decrease.
-Particularly, the smallest model excels in ETTh2, ETTm1, ETTm2, and Weather, while the largest model excels in Electricity.
-B.7 Impact of the different image encoding strategies
-Table 14: Impact of resampling filters and image orientations.
-Interpolation strategies in resampling Image orientation
-Method Bilinear Bicubic Nearest Neighbor Method - Horizontal flip Vertical flip
-Metric MSE MAE MSE MAE MSE MAE Metric MSE MAE MSE MAE MSE MAE
-$ETTh1$ 96 0.353 0.383 0.351 0.383 0.426 0.424 $ETTh1$ 96 0.353 0.383 0.348 0.379 0.355 0.385
-192 0.392 0.410 0.392 0.409 0.450 0.443 192 0.392 0.410 0.386 0.404 0.394 0.411
-336 0.407 0.423 0.407 0.422 0.451 0.450 336 0.407 0.423 0.401 0.416 0.408 0.423
-720 0.406 0.441 0.405 0.440 0.454 0.470 720 0.406 0.441 0.399 0.430 0.406 0.442
-avg 0.390 0.414 0.389 0.414 0.445 0.446 avg 0.390 0.414 0.384 0.407 0.391 0.415
-$ETTh2$ 96 0.271 0.328 0.274 0.329 0.298 0.349 $ETTh2$ 96 0.271 0.328 0.274 0.329 0.274 0.330
-192 0.328 0.367 0.330 0.367 0.343 0.380 192 0.328 0.367 0.331 0.370 0.330 0.367
-336 0.345 0.381 0.345 0.380 0.373 0.401 336 0.345 0.381 0.347 0.386 0.345 0.381
-720 0.388 0.422 0.386 0.419 0.404 0.431 720 0.388 0.422 0.376 0.416 0.388 0.422
-avg 0.333 0.375 0.334 0.374 0.354 0.390 avg 0.333 0.375 0.332 0.375 0.334 0.375
-$ETTm1$ 96 0.341 0.347 0.366 0.354 0.399 0.374 $ETTm1$ 96 0.341 0.347 0.345 0.348 0.342 0.347
-192 0.360 0.360 0.383 0.367 0.397 0.376 192 0.360 0.360 0.364 0.362 0.360 0.360
-336 0.377 0.374 0.396 0.381 0.386 0.380 336 0.377 0.374 0.378 0.375 0.377 0.374
-720 0.416 0.405 0.429 0.409 0.417 0.409 720 0.416 0.405 0.419 0.408 0.417 0.405
-avg 0.374 0.372 0.393 0.378 0.400 0.384 avg 0.374 0.372 0.376 0.373 0.374 0.372
-$ETTm2$ 96 0.228 0.282 0.246 0.296 0.264 0.326 $ETTm2$ 96 0.228 0.282 0.230 0.286 0.228 0.283
-192 0.262 0.305 0.273 0.313 0.273 0.328 192 0.262 0.305 0.264 0.308 0.262 0.305
-336 0.293 0.328 0.303 0.334 0.297 0.343 336 0.293 0.328 0.298 0.332 0.293 0.328
-720 0.343 0.370 0.343 0.370 0.334 0.369 720 0.343 0.370 0.350 0.373 0.343 0.369
-avg 0.282 0.321 0.291 0.328 0.292 0.341 avg 0.282 0.321 0.285 0.325 0.282 0.321
-Average 0.344 0.370 0.352 0.373 0.373 0.391 Average 0.344 0.370 0.344 0.370 0.345 0.371
-$\bf 1^{st}$ count 30 18 2 $\bf 1^{st}$ count 28 16 21
-Table 15: Computational cost in terms of seconds for forecasting a batch of 32 time series data.
-Context Length 1k 2k 3k 4k 1k
-Prediction Length 1k 1k 2k 3k 4k
-PatchTST 0.01 0.01 0.01 0.02 0.01 0.01 0.01 0.02
-GPT4TS 0.03 0.05 0.09 0.12 0.03 0.03 0.03 0.03
-MoiraiSmall 0.04 0.05 0.06 0.07 0.04 0.05 0.06 0.07
-MoiraiBase 0.06 0.08 0.11 0.13 0.06 0.08 0.11 0.13
-MoiraiLarge 0.14 0.19 0.25 0.30 0.14 0.19 0.25 0.30
-VisionTS ($c=1.0$) 0.03 0.03 0.03 0.03 0.02 0.02 0.02 0.02
-VisionTS ($c=0.4$) 0.02 0.02 0.02 0.02 0.02 0.02 0.02 0.02
-Table 14 summarizes the impact of interporation strategies and image orientations in the Alignment step.
-It shows that the smoother Bilinear and Bicubic interporation perform similarly, both significantly better than the rougher Nearest Neighbor.
-This suggests that smooth resizing effectively handles time series interpolation.
-Moreover, image orientation has little impact on performance.
-B.8 Computation cost
-We evaluate the computation cost of different baselines on an NVIDIA A800 GPU.
-For VisionTS, we test two variants, $c=0.4$ and $c=1.0$, since the alignment constant $c$ influences the number of visible patches and determines the computation cost of MAE encoder.
-Results are averaged on 20 runs.
-Table 15 compares the computation cost between various baselines, showing that VisionTS are faster than GPT4TS and three Moirai variants.
-While computation time increases with context length for all the other Transformer-based baselines, VisionTS remains nearly constant.
-This is because VisionTS encodes input sequences into an image with constant size, ensuring $O(1)$ efficiency.
-In contrast, Transformer-based methods operate at $O(L^{2})$ relative to context length $L$.
-B.9 Hyperparameter analysis
-Figs. 5, 6 and 7 show the influence of three hyperparameters, $r$, $c$, and $L$. We report the MSE averaged on four prediction lengths {96, 192, 336, 720}.
-(a) ETTh1
-(b) ETTh2
-(c) ETTm1
-(d) ETTm2
-(e) Electricity
-(f) Weather
-Figure 5: MSE (Y-axis) performance of different normalization constants $r$ (X-axis).
-(a) ETTh1
-(b) ETTh2
-(c) ETTm1
-(d) ETTm2
-(e) Electricity
-(f) Weather
-Figure 6: MSE (Y-axis) performance of different alignment constants $c$ (X-axis).
-(a) ETTh1
-(b) ETTh2
-(c) ETTm1
-(d) ETTm2
-(e) Electricity
-(f) Weather
-Figure 7: MSE (Y-axis) performance of different context lengths $L$ (X-axis).
-Appendix C Full-Shot Forecasting
-C.1 Training details
-Table 16: Final hyperparameters for VisionTS used in our full-shot forecasting.
-ETTh1 ETTh2 ETTm1 ETTm2 Illness Weather Traffic Electricity
-Normalization constant $r$ 0.4 0.4 0.4 0.4 1.0 1.0 0.4 0.4
-Alignment constant $c$ 0.4 0.4 0.4 0.4 0.4 0.7 0.4 0.4
-Context length $L$ 1152 1152 2304 1152 104 576 1152 1152
-Based on the principle of channel independence (Nie et al., 2022; Han et al., 2024), we treat the variables of each time series as individual data samples.
-We use an Adam optimizer with a learning rate 0.0001 and a batch size 256 to fine-tune MAE.
-All experiments are repeated three times.
-The training epoch is one for all the datasets except Illness, for which we train MAE for 100 epochs with an early stop due to the limited training dataset scale.
-We conduct tuning on validation sets for the three hyperparameters, $r$, $c$, and $L$.
-The final hyperparameters used are summarized in Table 16.
-C.2 Standard deviations
-Table 17: Standard deviations of full-shot experiments.
-Method VisionTS Time-LLM GPT4TS
-Metric MSE MAE MSE MAE MSE MAE
-ETTh1 96 0.347 $\pm$ 0.002 0.376 $\pm$ 0.000 0.376 $\pm$ 0.003 0.402 $\pm$ 0.002 0.370 $\pm$ 0.003 0.389 $\pm$ 0.001
-192 0.385 $\pm$ 0.001 0.400 $\pm$ 0.000 0.407 $\pm$ 0.003 0.421 $\pm$ 0.002 0.412 $\pm$ 0.003 0.413 $\pm$ 0.001
-336 0.407 $\pm$ 0.001 0.415 $\pm$ 0.001 0.430 $\pm$ 0.004 0.438 $\pm$ 0.001 0.448 $\pm$ 0.003 0.431 $\pm$ 0.001
-720 0.439 $\pm$ 0.001 0.443 $\pm$ 0.000 0.457 $\pm$ 0.003 0.468 $\pm$ 0.001 0.441 $\pm$ 0.003 0.449 $\pm$ 0.001
-ETTh2 96 0.269 $\pm$ 0.003 0.328 $\pm$ 0.002 0.286 $\pm$ 0.003 0.346 $\pm$ 0.002 0.280 $\pm$ 0.001 0.335 $\pm$ 0.001
-192 0.332 $\pm$ 0.001 0.374 $\pm$ 0.001 0.361 $\pm$ 0.003 0.391 $\pm$ 0.002 0.348 $\pm$ 0.002 0.380 $\pm$ 0.001
-336 0.351 $\pm$ 0.002 0.395 $\pm$ 0.002 0.390 $\pm$ 0.003 0.414 $\pm$ 0.002 0.380 $\pm$ 0.002 0.405 $\pm$ 0.001
-720 0.390 $\pm$ 0.003 0.430 $\pm$ 0.002 0.405 $\pm$ 0.003 0.434 $\pm$ 0.002 0.406 $\pm$ 0.002 0.436 $\pm$ 0.001
-ETTm1 96 0.281 $\pm$ 0.001 0.322 $\pm$ 0.001 0.291 $\pm$ 0.001 0.341 $\pm$ 0.001 0.300 $\pm$ 0.001 0.340 $\pm$ 0.000
-192 0.322 $\pm$ 0.006 0.353 $\pm$ 0.002 0.341 $\pm$ 0.001 0.369 $\pm$ 0.001 0.343 $\pm$ 0.001 0.368 $\pm$ 0.000
-336 0.356 $\pm$ 0.003 0.379 $\pm$ 0.002 0.359 $\pm$ 0.002 0.379 $\pm$ 0.001 0.376 $\pm$ 0.001 0.386 $\pm$ 0.000
-720 0.391 $\pm$ 0.001 0.413 $\pm$ 0.001 0.433 $\pm$ 0.001 0.419 $\pm$ 0.001 0.431 $\pm$ 0.001 0.416 $\pm$ 0.000
-ETTm2 96 0.169 $\pm$ 0.003 0.256 $\pm$ 0.002 0.162 $\pm$ 0.001 0.248 $\pm$ 0.001 0.163 $\pm$ 0.001 0.249 $\pm$ 0.001
-192 0.225 $\pm$ 0.003 0.294 $\pm$ 0.003 0.235 $\pm$ 0.002 0.304 $\pm$ 0.001 0.222 $\pm$ 0.001 0.291 $\pm$ 0.000
-336 0.278 $\pm$ 0.002 0.334 $\pm$ 0.001 0.280 $\pm$ 0.002 0.329 $\pm$ 0.001 0.273 $\pm$ 0.001 0.327 $\pm$ 0.001
-720 0.372 $\pm$ 0.002 0.392 $\pm$ 0.002 0.366 $\pm$ 0.002 0.382 $\pm$ 0.001 0.357 $\pm$ 0.001 0.376 $\pm$ 0.001
-Weather 96 0.142 $\pm$ 0.000 0.192 $\pm$ 0.001 0.155 $\pm$ 0.001 0.199 $\pm$ 0.001 0.148 $\pm$ 0.001 0.188 $\pm$ 0.000
-192 0.191 $\pm$ 0.000 0.238 $\pm$ 0.000 0.223 $\pm$ 0.001 0.261 $\pm$ 0.001 0.192 $\pm$ 0.001 0.230 $\pm$ 0.000
-336 0.246 $\pm$ 0.003 0.282 $\pm$ 0.001 0.251 $\pm$ 0.001 0.279 $\pm$ 0.001 0.246 $\pm$ 0.001 0.273 $\pm$ 0.000
-720 0.328 $\pm$ 0.004 0.337 $\pm$ 0.001 0.345 $\pm$ 0.001 0.342 $\pm$ 0.001 0.320 $\pm$ 0.001 0.328 $\pm$ 0.000
-Traffic 96 0.344 $\pm$ 0.001 0.236 $\pm$ 0.000 0.392 $\pm$ 0.001 0.267 $\pm$ 0.000 0.396 $\pm$ 0.001 0.264 $\pm$ 0.000
-192 0.372 $\pm$ 0.001 0.249 $\pm$ 0.001 0.409 $\pm$ 0.001 0.271 $\pm$ 0.000 0.412 $\pm$ 0.001 0.268 $\pm$ 0.000
-336 0.383 $\pm$ 0.001 0.257 $\pm$ 0.001 0.434 $\pm$ 0.001 0.296 $\pm$ 0.000 0.421 $\pm$ 0.001 0.273 $\pm$ 0.000
-720 0.422 $\pm$ 0.001 0.280 $\pm$ 0.000 0.451 $\pm$ 0.001 0.291 $\pm$ 0.000 0.455 $\pm$ 0.001 0.291 $\pm$ 0.000
-Electricity 96 0.126 $\pm$ 0.000 0.218 $\pm$ 0.000 0.137 $\pm$ 0.000 0.233 $\pm$ 0.000 0.141 $\pm$ 0.000 0.239 $\pm$ 0.000
-192 0.146 $\pm$ 0.001 0.239 $\pm$ 0.001 0.152 $\pm$ 0.000 0.247 $\pm$ 0.000 0.158 $\pm$ 0.000 0.253 $\pm$ 0.000
-336 0.161 $\pm$ 0.001 0.255 $\pm$ 0.001 0.169 $\pm$ 0.000 0.267 $\pm$ 0.000 0.172 $\pm$ 0.000 0.266 $\pm$ 0.000
-720 0.193 $\pm$ 0.000 0.286 $\pm$ 0.000 0.200 $\pm$ 0.000 0.290 $\pm$ 0.000 0.207 $\pm$ 0.000 0.293 $\pm$ 0.000
-$\bf 1^{st}$ count 42 2 12
-We report the standard deviations of our full-shot experiments computed on three runs in Table 17, including the results of Time-LLM and GPT4TS from Tan et al. (2024) for reference.
-C.3 Ablation study and fine-tuning strategy comparison
-Table 18: Full results of Tables 5 and 5: Ablation studies (left) and fine-tuning strategies (right). Results are averaged on four prediction lengths: {96, 192, 336, 720}.
+Table 19 shows the full results of the full-shot experiments.
+We also report the standard deviations of our full-shot experiments computed on three runs in Table 18, including the results of Time-LLM and GPT4TS from Tan et al.
+(2024) for reference.
+D.3 Ablation study and fine-tuning strategy comparison
+Table 20: Ablation studies (left) and fine-tuning strategies (right). Results are averaged on four prediction lengths: {96, 192, 336, 720}.
 Ablation on Visual MAE (VM) Ablation on trained parameters
@@
 The results are shown in Table 20, suggesting that visual knowledge is crucial for VisionTS and fine-tuning the layer normalization is the best.
-Appendix D Visualization
+Appendix E Visualization
 We visualized the predictions of VisionTS in the zero-shot setting, including its input and reconstructed images.
-We also visualized the predictions of MoiraiLarge and Seasonal Naïve, wich their MAE metrics for comparison.
+We also visualized the predictions of MoiraiLarge and Seasonal Naïve, with their MAE metrics for comparison.
 Figs.
-8, 9 and 10 show examples where VisionTS performed well, with Fig.
+11, 12 and 13 show examples where VisionTS performed well, with Fig.
 11 depicting a more regular pattern, while Figs.
-9 and 10 display less obvious patterns.
+12 and 13 display less obvious patterns.
 Fig.

### v4

@@
 Extensive experiments reveal intrinsic similarities between images and real-world time series, suggesting that visual models may offer a “free lunch” for TSF and highlight the potential for future cross-modality research.
-Our code is available in the Supplementary Material.
+Our code is publicly available at https://github.com/Keytoyze/VisionTS.
 Time Series Forecasting, Foundation Models, Transfer Learning
-Figure 1: Long-term forecasting performance comparison. Our VisionTS, without any training on time series data, outperforms the pure time series foundation model MoiraiLarge in the zero-shot setting.
+Figure 1: Long-term forecasting (left) and GIFT-Eval (right) performance comparison.
+Our VisionTS, without any training on time series data, outperforms the pure time series foundation models in the zero-shot setting.
 1 Introduction
@@
 We propose to segment it into $\lfloor\nicefrac{{L}}{{P}}\rfloor$ subsequences of length $P$, where $P$ is the periodicity.
-Notably, when the time series lacks clear periodicity, we can set $P=1$ directly, which is also effective in our experiments (Section C.5).
+Notably, when the time series lacks clear periodicity, we can set $P=1$ directly, which is also effective in our experiments (Section B.6).
 In practice, $P$ can be determined using statistical methods like Fast Fourier Transform (Wu et al., 2023; Chen et al., 2024) or domain knowledge like sampling frequency (Godahewa et al., 2021; Alexandrov et al., 2020).
-In this paper, we select $P$ based on the sampling frequency, elaborated in Section B.2.
+In this paper, we select $P$ based on the sampling frequency, elaborated in Section A.2.
 After that, these subsequences are then stacked into a 2D matrix, denoted by ${\bm{I}}_{\text{raw}}\in\mathbb{R}^{P\times\lfloor\nicefrac{{L}}{{P}}\rfloor}$.
@@
 However, an excessively low $r$ can result in values that are difficult to distinguish.
-We found that a moderate value (0.4) of $r$ performs well across most scenarios (See Section C.8 for more details).
+We found that a moderate value (0.4) of $r$ performs well across most scenarios (See Section B.9 for more details).
 Let ${\bm{I}}_{\text{norm}}$ denote the normalized matrix, which is computed as follows:
@@
 Therefore, we propose multiplying $n$ by a hyperparameter $c\in[0,1]$.
-Similar to $r$, we found that setting $c=0.4$ performs well in our experiments (See Section C.8).
+Similar to $r$, we found that setting $c=0.4$ performs well in our experiments (See Section B.9).
 This can be formulated as $n=\left\lfloor c\cdot N\cdot\nicefrac{{L}}{{(L+H)}}\right\rfloor$.
@@
 Table 1: Zero-shot or few-shot results on the long-term TSF benchmark.
-Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Table 9 (Section C.2).
+Results are averaged across prediction lengths {96, 192, 336, 720}, with full results in Table 9 (Section B.2).
 Bold: the best result.
@@
 4 Experiments
-Figure 4: Performance on the GIFT-Eval Leaderboard.
-Figure 5: Aggregated results on the Monash TSF Benchmark, with full results in Table 13 (Section C.5).
+Figure 4: Performance on the GIFT-Eval Leaderboard (cut-off at VisionTS’s release).
+Figure 5: Aggregated results on the Monash TSF Benchmark, with full results in Table 15 (Section B.6).
 We follow the standard evaluation protocol proposed by Woo et al.
@@
 We use MAE (Base) as our backbone by default.
-Baseline and benchmark details are elaborated in Section B.1.
+Baseline and benchmark details are elaborated in Section A.1.
 4.1 Zero-Shot Time Series Forecasting
@@
 We set the hyperparameters to $r=c=0.4$.
-Following common practice (Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section C.1.
-Table 2: Average MSE of different MAE variants, with full results in Table 15 (Section C.6).
+Following common practice (Zhou et al., 2023; Woo et al., 2024), we conduct hyperparameter tuning on validation sets to determine the optimal context length $L$, detailed in Section B.1.
+Table 2: Average MSE of different MAE variants, with full results in Table 17 (Section B.7).
 Base Large Huge
@@
 It suggests that the transferability from images to time-series is stronger than from text to time-series, and even comparable to the in-domain transferability between time-series.
-We also include a comparison with two TSF foundation models, TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023), in Section C.3, as well as traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section C.4.
+We also include a comparison with two TSF foundation models, TimesFM (Das et al., 2024) and LLMTime (Gruver et al., 2023), in Section B.3, as well as traditional algorithms (ETS, ARIMA, and Seasonal Naïve) in Section B.4.
 Results show that VisionTS still outperforms all of these baselines.
-Results on GIFT-Eval and Monash Benchmarks
+Results on GIFT-Eval Benchmarks
 Fig.
 4 shows the comparison of VisionTS with six previously published TSF foundation models on the GIFT-Eval TSF Leaderboard [^2] https://huggingface.co/spaces/Salesforce/GIFT-Eval, where VisionTS surprisingly ranked first in terms of normalized MASE.
-Besides, Fig.
+It should be noted that although some concurrent works (after the release of VisionTS) in the current leaderboard outperform VisionTS, there may be data leakage issues for these works.
+In contrast, visual MAE was trained on ImageNet, long before the release of GIFT-Eval leaderboard, which can ensure no leakage.
+Results on Monash Benchmark
+Fig.
 5 shows the results aggregated from 29 Monash datasets, showing that VisionTS in the zero-shot setting surpasses all models individually trained on each dataset and significantly outperforms the other cross-domain baseline (i.e., LLMTime).
@@
 Backbone Analysis
-In Table 3 (full results in Section C.6), we observe that the overall performance of three MAE variants (112M, 330M, and 657M) outperforms MoiraiSmall and MoiraiLarge.
+In Table 3 (full results in Section B.7), we observe that the overall performance of three MAE variants (112M, 330M, and 657M) outperforms MoiraiSmall and MoiraiLarge.
 Particularly, larger models show a slight decrease in performance.
@@
 Additionally, to explore the potential with other vision models, we also test LaMa (Suvorov et al., 2022), a visual inpainting model.
-Results in Section C.6 demonstrate that VisionTS with LaMa performs similarly to Moirai in the zero-shot setting.
+Results in Section B.7 demonstrate that VisionTS with LaMa performs similarly to Moirai in the zero-shot setting.
 This suggests that the performance is driven by the inherent similarity between images and time series, not solely by the MAE model.
@@
 VisionTS is fine-tuned only a single epoch on each dataset except for Illness.
-Due to the space limit, we report the $1^{\text{st}}$ count for each baseline, with full results in Table 19 (Section D.2).
+Due to the space limit, we report the $1^{\text{st}}$ count for each baseline, with full results in Table 21 (Section C.2).
 Pretrain [Uncaptioned image] Images [Uncaptioned image] Text [Uncaptioned image] No Pretrain
@@
 Hyperparameter Analysis
-Section C.8 illustrates the impact of three hyperparameters.
+Section B.9 illustrates the impact of three hyperparameters.
 For context length $L$, as shown in Fig.
@@
 Ablation Study
-We conduct experiments to validate our choices in the Alignment step, detailed in Section C.7.
+We conduct experiments to validate our choices in the Alignment step, detailed in Section B.8.
 First, we test three different interpolation strategies, which shows that Bilinear interpolation performs best.
@@
 Qualitative Analysis: When does VisionTS perform well, and when does it not?
-In Appendix E, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails).
+In Appendix D, we visualize the zero-shot forecasting of VisionTS alongside the input and reconstruction images, highlighting both successful cases (where VisionTS outperforms Moirai) and failures (where Moirai prevails).
 When the input exhibits strong regularity (Fig.
@@
 (2023).
-Training details are elaborated in Section D.1.
+Training details are elaborated in Section C.1.
 Main Results
-Table 4 summarizes the full-shot results, with the full results and standard deviations detailed in Section D.2.
+Table 4 summarizes the full-shot results, with the full results and standard deviations detailed in Section C.2.
 It shows that VisionTS outperforms other baselines in most cases (46 out of 80), surpassing the non-pretrained PatchTST and the language-pretrained GPT4TS.
@@
 We conducted similar ablations to assess the role of the vision model (VM), including w/o VM, VM2Attn, VM2Trsf, and RandVM.
-Section D.3 shows that these variants lead to worse performance, indicating that visual knowledge is beneficial for TSF.
+Section C.3 shows that these variants lead to worse performance, indicating that visual knowledge is beneficial for TSF.
 Analysis: Fine-tuning strategies
@@
 Note that freezing differs from the previous zero-shot experiment, where a longer context length was used.
-Section D.3 show that fine-tuning LN is the best.
+Section C.3 show that fine-tuning LN is the best.
 Modifying MLP or attention layers results in significant performance drops, suggesting that valuable knowledge resides in these components.
@@
 Text-based TSF Foundation Models
-Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks. We review this line of work in Appendix A.
+Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks.
+For example, Zhou et al.
+(2023) fine-tuned a pre-trained GPT (Radford et al., 2019) on each time-series downstream task, such as forecasting, classification, imputation, and anomaly detection.
+Based on Llama (Touvron et al., 2023), Jin et al.
+(2024) froze the pre-trained LLM and reprogrammed the time series to align with the language modality.
+Bian et al.
+(2024) adopted a two-stage approach by continually pre-training GPT (Radford et al., 2019) on the time-series domain.
+Nevertheless, the TSF performance of LLMs has recently been questioned by Tan et al.
+(2024), which designed several ablation studies to show that textual knowledge is unnecessary for forecasting.
+In this paper, we attribute it to the large modality gap.
+Some recent approaches focus on directly transforming the time series into natural texts for LLMs, allowing for zero-shot forecasting.
+For example, PromptCast (Xue & Salim, 2023) used pre-defined templates to describe numerical time series data, while LLMTime (Gruver et al., 2023) directly separated time steps using commas and separates digits using spaces to construct the text input.
+However, due to the efficiency issue of the autoregressive decoding strategy and the expensive inference cost of large language models, their practical use is limited.
 Time Series-Based TSF Foundation Models
 Self-supervised pre-training a TSF model on the same dataset used for downstream TSF tasks is a well-explored topic (Ma et al., 2023; Zhang et al., 2024), such as denoising autoencoders (Zerveas et al., 2021) or contrastive learning (Woo et al., 2022a; Yue et al., 2022).
-They follow a similar paradigm to the masked autoencoder (MAE) in computer vision.
+They follow a similar paradigm to the masked autoencoder (MAE) in computer vision, which is a well-studied topic in other machine learning fields, such as BERT (Devlin et al., 2019), CBraMod (Wang et al., 2025), and HuBERT (Hsu et al., 2021).
 However, these methods rarely examine the cross-dataset generalization capabilities.
-Recently, research has shifted towards training universal foundation models, by collecting large-scale time series datasets from diverse domains (Goswami et al., 2024; Liu et al., 2024; Das et al., 2024; Dong et al., 2024; Feng et al., 2024) or generating numerous synthetic time series data (Fu et al., 2024; Yang et al., 2024).
+Recently, research has shifted towards training universal foundation models, by collecting large-scale time series datasets from diverse domains (Ansari et al., 2024; Goswami et al., 2024; Liu et al., 2024; Das et al., 2024; Dong et al., 2024; Feng et al., 2024) or generating numerous synthetic time series data (Fu et al., 2024; Yang et al., 2024).
 As a representative method, Woo et al.
@@
 Future modifications to the model structure may empower it with more time series capabilities.
+Acknowledgments
+Research work mentioned in this paper is supported by State Street Zhejiang University Technology Center. We would also like to thank reviewers for their valuable comments.
 Impact Statement
 This paper presents work whose goal is to advance the field of time series forecasting.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix
-Appendix A Extended Related Work
-Text-based TSF Foundation Models
-Large Language Models (LLMs) pre-trained on large amounts of text data are being applied to TSF tasks.
-For example, Zhou et al.
-(2023) fine-tuned a pre-trained GPT (Radford et al., 2019) on each time-series downstream task, such as forecasting, classification, imputation, and anomaly detection.
-Based on Llama (Touvron et al., 2023), Jin et al.
-(2024) froze the pre-trained LLM and reprogrammed the time series to align with the language modality.
-Bian et al.
-(2024) adopted a two-stage approach by continually pre-training GPT (Radford et al., 2019) on the time-series domain.
-Nevertheless, the TSF performance of LLMs has recently been questioned by Tan et al.
-(2024), which designed several ablation studies to show that textual knowledge is unnecessary for forecasting.
-In this paper, we attribute it to the large modality gap.
-Some recent approaches focus on directly transforming the time series into natural texts for LLMs, allowing for zero-shot forecasting.
-For example, PromptCast (Xue & Salim, 2023) used pre-defined templates to describe numerical time series data, while LLMTime (Gruver et al., 2023) directly separated time steps using commas and separates digits using spaces to construct the text input.
-However, due to the efficiency issue of the autoregressive decoding strategy and the expensive inference cost of large language models, their practical use is limited.
-Appendix B Details of Experiments
-B.1 Benchmark and baselines
+Appendix A Details of Experiments
+A.1 Benchmark and baselines
 Long-Term TSF Benchmark
@@
 All experiments are conducted using Time-Series-Library (https://github.com/thuml/Time-Series-Library) and GluonTS library (Alexandrov et al., 2020) on an NVIDIA A800 GPU.
-B.2 Periodicity selection
+A.2 Periodicity selection
 We first determine a range of period lengths based on the sampling frequency of the data, shown in Table 5.
@@
 Table 7 shows that such strategy (denoted as VisionTS) significantly outperforms the naive strategy that sets $P=1$.
-Appendix C Zero-Shot Forecasting
-C.1 Hyperparameters
+Appendix B Zero-Shot Forecasting
+B.1 Hyperparameters
 Table 8: Hyperparameters for VisionTS used in our zero-shot forecasting (Long-term TSF).
@@
 We conduct hyperparameter tuning on validation sets to determine the optimal context length $L$. Final used hyperparameters are summarized in Table 8.
-C.2 Full forecasting results of the long-term TSF benchmark
+B.2 Full forecasting results of the long-term TSF benchmark
 Table 9: Full results of Table 1: Zero-shot or few-shot results on the long-term TSF benchmark. Bold: the best result.
@@
 VisionTS achieves the best results in most cases (32 out of 62), outperforming MoiraiBase (10 out of 62) and MoiraiLarge (8 out of 62).
-C.3 Comparison of TimesFM and LLMTime
+B.3 Comparison of TimesFM and LLMTime
 Table 10: MAE results of TimesFM and LLMTime for zero-shot forecasting, on the last test window of the original test split.
@@
 VisionTS outperforms TimesFM and LLMTime in terms of MAE, indicating that image-based TSF models are on par with or even better than TS-based and text-based models.
-C.4 Comparison of traditional methods
+B.4 Comparison of traditional methods
 In addition to deep learning models, we also compare traditional methods, including ARIMA, ETS, and two methods that require periodicity as our VisionTS: Seasonal Naïve (repeating the last period) and Seasonal Avg (similar to Seasonal Naïve but repeating the average of all periods in the look-back window).
@@
 Table 12 shows that VisionTS also achieves the best performance.
-C.5 Full forecasting results of the Monash TSF benchmark
+B.5 Comparison of concurrent works
+Table 13: Comparison of Time-MoE and TTM in the zero-shot setting.
+We report the base and large model results for Time-MoE, as the ultra model weights are not yet released.
+For TTM, we used the official HuggingFace model for replication.
+The following table summarizes the performance of various zero-shot foundation models.
+VisionTS Time-MoE (base) Time-MoE (large) TTM (v1)
+ETTh1 MSE 0.390 0.400 0.394 0.398
+MAE 0.414 0.424 0.419 0.421
+ETTh2 MSE 0.333 0.366 0.405 0.348
+MAE 0.375 0.404 0.415 0.393
+ETTm1 MSE 0.374 0.394 0.376 0.520
+MAE 0.372 0.415 0.405 0.479
+ETTm2 MSE 0.282 0.317 0.316 0.312
+MAE 0.321 0.365 0.361 0.348
+Electricity MSE 0.207 (data leakage) (data leakage) 0.201
+MAE 0.294 (data leakage) (data leakage) 0.293
+Weather MSE 0.269 0.265 0.270 0.234
+MAE 0.292 0.297 0.300 0.266
+Average MSE 0.309 - - 0.335
+MAE 0.345 - - 0.367
+[…]
+Table 14: Comparison of CALF in both zero-shot and full-shot settings.
+VisionTS (zero-shot) VisionTS (full-shot) CALF
+ETTh1 MSE 0.390 0.395 0.432
+MAE 0.414 0.409 0.428
+ETTh2 MSE 0.333 0.336 0.349
+MAE 0.375 0.382 0.382
+ETTm1 MSE 0.374 0.338 0.395
+MAE 0.372 0.367 0.390
+ETTm2 MSE 0.282 0.261 0.281
+MAE 0.321 0.319 0.321
+Electricity MSE 0.207 0.156 0.175
+MAE 0.294 0.249 0.265
+Weather MSE 0.269 0.227 0.250
+MAE 0.292 0.262 0.274
+Average MSE 0.309 0.286 0.314
+MAE 0.345 0.331 0.343
+We compare our work with other concurrent TSF methods.
+Table 13 presents the comparison from Time-MoE (Shi et al., 2024) and TTM (Ekambaram et al., 2024), and Table 14 shows the comparison with CALF (Liu et al., 2025), which is the existing SOTA LLMs-based time series forecasting work.
+These findings highlight the promising potential of vision models in TSF scenarios.
+B.6 Full forecasting results of the Monash TSF benchmark
 Table 15: Full results of Fig. 5: Forecasting results (MAE) on the Monash TSF benchmark. We reported the reproduction results of LLMTime based on the GPT3.5 API from Woo et al. (2024).
@@
 Average 0.344 0.370 0.374 0.392 0.372 0.388 0.382 0.388
-C.6 Impact of backbones
+B.7 Impact of backbones
 Table 17: Full results of Table 3: zero-shot forecasting results of different MAE variants. Bold: best results among three variants. We also include the results from Moirai for reference.
@@
 Additionally, Table 16 compares VisionTS with another visual backbone, LaMa.
-C.7 Impact of the different image encoding strategies
+B.8 Impact of the different image encoding strategies
 Table 18: Impact of resampling filters and image orientations.
@@
 Moreover, image orientation has little impact on performance.
-C.8 Hyperparameter analysis
+B.9 Hyperparameter analysis
 Figs. 8, 9 and 10 show the influence of three hyperparameters, $r$, $c$, and $L$. We report the MSE averaged on four prediction lengths {96, 192, 336, 720}.
@@
 Figure 10: MSE (Y-axis) performance of different context lengths $L$ (X-axis).
-Appendix D Full-Shot Forecasting
-D.1 Training details
+Appendix C Full-Shot Forecasting
+C.1 Training details
 Table 19: Final hyperparameters for VisionTS used in our full-shot forecasting.
@@
 The final hyperparameters used are summarized in Table 19.
-D.2 Full results and standard deviations
+C.2 Full results and standard deviations
 Table 20: Standard deviations of full-shot experiments.
@@
 (2024) for reference.
-D.3 Ablation study and fine-tuning strategy comparison
+C.3 Ablation study and fine-tuning strategy comparison
 Table 22: Ablation studies (left) and fine-tuning strategies (right). Results are averaged on four prediction lengths: {96, 192, 336, 720}.
@@
 The results are shown in Table 22, suggesting that visual knowledge is crucial for VisionTS and fine-tuning the layer normalization is the best.
-Appendix E Visualization
+Appendix D Visualization
 We visualized the predictions of VisionTS in the zero-shot setting, including its input and reconstructed images.